### v4

# Approximations on Spirallike domains of $\mathbb{C}^{n}$

Abstract.
In this paper, we first show that any domain $\Omega$ in $\mathbb{C}^{n}(n\geq 2)$, which is spirallike with respect to a complete, holomorphic, globally asymptotic stable vector field $V$, is a Runge domain. Next, we prove an Andersén-Lempert type approximation theorem: If a domain $\Omega\subseteq\mathbb{C}^{n}$ is spirallike with respect to a complete, globally exponentially stable, vector field $V$, with a certain rate of the convergence of the flow of the vector field $V$ in $\Omega$ then any biholomorphism $\Phi\colon\Omega\to\Phi(\Omega)$, with $\Phi(\Omega)$ is Runge, can be approximated by automorphisms of $\mathbb{C}^{n}$ uniformly on compacts. This generalizes all earlier known theorems in this direction substancially, even when the vector field is linear. We further show that, if $J(\Phi)\equiv 1$ and $div(V)$ is constant then the biholomorphism $\Phi\colon\Omega\to\Phi(\Omega)$ can be approximated by volume-preserving automorphism of $\mathbb{C}^{n}$. We also provide an analogous result when the given biholomorphism preserves the symplectic form. As an application of our approximation results, we show that any Loewner PDE in a complete hyperbolic domain $\Omega$ which is spirallike with respect to a globally asymptotic stable vector field $V$, having some uniform exponential convergence rate on the domain $\Omega$, admits an essentially unique univalent solution with values in $\mathbb{C}^{n}$. We also provide an example of a Hartogs domain in $\mathbb{C}^{2}$ which spirallike with respect to a complete holomorphic vector field $F(z_{1},z_{2})=(-2z_{1},-3z_{2}+z_{1}z_{2})$, but the domain is not spirallike with respect to any linear vector field. Some more examples are provided at the end of this paper.
Key words and phrases:
Loewner theory, Runge domain, Spirallike domain, Automorphisms of $\mathbb{C}^{n}$
2020 Mathematics Subject Classification:
Primary: 32M17, 32E30, 30C45
Sanjoy Chatterjee is supported by CSIR fellowship (File No-09/921(0283)/2019-EMR-I). Sushil Gorai is partially supported by a Matrics Research Grant (MTR/2017/000974) and a Core Research Grant (CRG/2022/003560) of SERB
1. Introduction
The question that we study in this paper is: For what classes of domains any biholomorphic maps into $\mathbb{C}^{n}$ can be approximated by automorphisms of $\mathbb{C}^{n}$? Thanks to the Andersén-Lempert theory, a positive answer of this question will lead to approximating any biholomorphic map by simpler classes of maps, namely the finite composition of shears and overshears. In this paper, we give a large class of domains for which such approximation holds. Before going into the discussion of this approximation question we mention a result about Runge domains which is very crucial in the sequel. Recall that a domain $\Omega\subseteq\mathbb{C}^{n}$ is said to be a Runge domain if every holomorphic function $f:\Omega\to\mathbb{C}$ can be approximated by holomorphic polynomials uniformly on every compact subset of $\Omega$. Runge domains are fundamental in the study of function theory in several complex variables. In one variable, they are the simply connected domains in $\mathbb{C}$, hence, a topological property classifies them. In higher dimensions, no such characterization exists. Characterizing the Runge domain in higher dimensions is a difficult problem. In this paper, we give a class of domains that are Runge. More precisely, we prove that
Theorem 1.1.
Let $F\in\mathfrak{X}_{\mathcal{O}}{(\mathbb{C}^{n})}$ be a complete globally asymptotically stable vector field such that $F(0)=0$. If $\Omega$ is a spirallike domain with respect to $F$ containing the origin then $\Omega$ is a Runge domain.
A holomorphic vector field $V$ on a domain $\Omega\subset\mathbb{C}^{n}$ is a real vector field on $\Omega$ such that
$V(z)=\sum_{i=1}^{n}a_{i}(z)\frac{\partial}{\partial x_{i}}+b_{i}(z)\frac{\partial}{\partial y_{i}}$
such that $(a_{j}+ib_{j})$ is holomorphic function on $\Omega$ for all $j\in\{1,2,\cdots,n\}$, $z_{j}=x_{j}+iy_{j}$ are coordinate of $\mathbb{C}^{n}$. We denote the set of all holomorphic vector fields on $\mathbb{C}^{n}$ by $\mathfrak{X}_{\mathcal{O}}(\mathbb{C}^{n})$. A complete vector field $V$ is said to be globally asymptotic stable if the vector field admits an equilibrium point such that the integral curve of the vector field, initiated at any given point in $\mathbb{C}^{n}$, converges to the equilibrium point as $t\to\infty$ (see Section 2 for detail). It is known that any star shaped domain in $\mathbb{C}^{n}$ is Runge (see [Kasimi]). Any star shaped domain with respect to the origin is spirallike with respect to the vector field $-I$. In [Hamada15, Theorem-3.1], Hamada proved that spirallike domain with respect to certain matrix is Runge. For detailed discussion about the spirallike domain see Section 2.
We now begin our discussion about the approximation of biholomorphic maps on spirallike domains by automorphisms of $\mathbb{C}^{n}$. For this purpose, a brief description of the automorphism group of $\mathbb{C}^{n}$ will be useful. Let Aut$(\mathbb{C}^{n})$ be set of all holomorphic mappings $f:\mathbb{C}^{n}\to\mathbb{C}^{n}$ such that $f^{-1}$ exist and also holomorphic. For $n=1$, we know that $(\mathbb{C})$ consists only of affine mappings $z\mapsto az+b,(a,b\in\mathbb{C},\;a\neq 0)$. However, $Aut(\mathbb{C}^{n})$ is huge for $n\geq 2$. Writing the standard coordinates as $(z_{1},z_{2},\cdots z_{n})=(z^{\prime},z_{n})$ on $\mathbb{C}^{n}$ we see that $Aut(\mathbb{C}^{n})$ contains mappings of the form
$(z^{\prime},z_{n})\mapsto(z^{\prime},z_{n}+f(z^{\prime}))$ (1.1)
$(z^{\prime},z_{n})\mapsto(z^{\prime},e^{f(z^{\prime})}z_{n}),$ (1.2)
where $f:\mathbb{C}^{n-1}\to\mathbb{C}$ is any holomorphic mapping. The inverse of the above mapping is obtained by replacing $f$ with $-f$ in (1.1) and (1.2). The mappings of forms (1.1) and their $Sl(n,\mathbb{C})$ conjugates are called shears or additive shears, and the mappings (1.2) and their $Gl(n,\mathbb{C})$ conjugates are called overshears or multiplicative shears. The subgroup generated by elements of overshears and shears is denoted by $S(n)$, $S_{1}(n)$, respectively. The set of all volume preserving automorphisms of $\mathbb{C}^{n}$ is denoted by $Aut_{1}(\mathbb{C}^{n})$ and is defined by $Aut_{1}(\mathbb{C}^{n}):=\{F\in\text{Aut}(\mathbb{C}^{n}):\text{det}(DF(z))\equiv 1\}$. Our next result is about approximating biholomorphic maps on some spirallike domains by compositions of shears and overshears, special members of the $\text{Aut}(\mathbb{C}^{n})$. Before starting our main result let us survey a few things from the existing literature. The study of $Aut(\mathbb{C}^{n})$ was initiated by Rosay and Rudin [RosayRudin]. Questions about the density of the compositions of shears in the class of all volume preserving automorphisms of $\mathbb{C}^{n}$ and the density of compositions of overshears in $Aut(\mathbb{C}^{n})$ were asked in [RosayRudin], which were answered positively by Andersén [Andersen] and by Andersén and Lempert [AndersenLemp] respectively, giving birth of the now-famous Andersén-Lempert theory [Andersen, AndersenLemp] (see [Forstbook] for more details). It is a natural question to ask for characterization of the domains $\Omega\subset\mathbb{C}^{n}$ for which any biholomorphism from $\Omega$ into $\mathbb{C}^{n}$ can be approximated uniformly on compacts of $\Omega$ by automorphisms of $\mathbb{C}^{n}$, i.e. in view of Andersén-Lempert theory by composition of Overshears. Andersén and Lempert [AndersenLemp] considered this question and proved:
Result 1.2.
[AndersenLemp, Theorem 2.1] Let $\Omega$ be a starshaped domain $\mathbb{C}^{n}$ and $\Phi:\Omega\to\mathbb{C}^{n}$ be a biholomorphic map such that $\Phi(\Omega)$ is Runge. Then $\Phi$ can be approximated uniformly on compact subsets by compositions of overshears. If $\det D\Phi\equiv 1$, then $\Phi$ can be approximated uniformly on compact subsets by the composition of shears.
Note that, if a biholomorphism $\Phi:\Omega\to\Phi(\Omega)$, with $\Phi(\Omega)$ Runge, can be approximated uniformly on compacts by elements of Aut$({\mathbb{C}^{n}})$, then in view of the density theorem due to Andersén and Lempert [AndersenLemp, Theorem 1.3], we get that the biholomorphism $\Phi$ can be approximated by the composition of overshears. In a seminal paper Forstnerič and Rosay [FR1993, Theorem-1.1], [Forst1994, Theorem-1.1] gave a sufficient condition for Runge domain under which any biholomorphic map can be approximated by automorphisms of $\mathbb{C}^{n}$.
Result 1.3 (Forstnerič-Rosay).
Let $\Omega\subset\mathbb{C}^{n}$ be a Runge domain. Assume that $H:[0,1]\times\Omega\to\mathbb{C}^{n}$ is a mapping of class $\mathcal{C}^{p}(0\leq p<\infty)$ such that for each $t\in[0,1]$, $H_{t}=H(t,\cdot)\colon\Omega\to\mathbb{C}^{n}$ is a biholomorphic mapping onto a Runge domain $\Omega_{t}:=H_{t}(\Omega)\subset\mathbb{C}^{n}$ and the map $H_{0}$ can be approximated uniformly on every compact subset of $\Omega$ by automorphisms of $\mathbb{C}^{n}$. Then
  • (i)
[…]
  • (ii)
    if we have $H_{0}\in Aut(\mathbb{C}^{n})$, then we can choose $\Psi$ such that $\Psi_{0}=H_{0}$,
  • (iii)
[…]
This result by Forstnerič and Rosay is the most general result in this direction. The smoothness assumption of the family is lowered in [Forst1994] and [FKu]. But, for a given domain, constructing such an isotopy $\{\Phi_{t}\}$ in Result 1.3 is often difficult. Therefore, it is interesting to know the classes of domains to which the above result can be applied. Hamada [Hamada15] gave a class of domains in the following result.
Result 1.4.
[Hamada15, Theorem 4.2] Let $\Omega$ be a domain containing the origin that satisfies the following conditions: $e^{-tA}w\in\Omega$ for all $w\in\Omega$, where $A\in M_{n}(\mathbb{C})$ such that $\inf_{||z||=1}{\sf Re}\langle Az,z\rangle>0$. If $\max_{\lambda\in\sigma(A)}{\sf Re}(\lambda)<2\inf_{||z||=1}{\sf Re}\langle Az,z\rangle$, where $\sigma(A)$ is the spectrum of $A$, then any biholomorphism $\phi:\Omega\to\mathbb{C}^{n}$ with $\phi(\Omega)$ Runge, can be approximated uniformly on compacts of $\Omega$ by automorphisms of $\mathbb{C}^{n}$.
In this paper, we consider the spirallike domain with respect to an exponentially stable, non-linear vector field (see Section 2 for the precise definition) and show that under certain conditions the result holds. To state it precisely we need the following:
Let $\sigma(A)$ denote the spectrum of the matrix $A$. Let $k_{+}(A):=\max\{{\sf Re}\lambda:\lambda\in\sigma(A)\}$ and $k_{-}(A):=\min\{{\sf Re}\lambda:\lambda\in\sigma(A)\}$. We now present the main result of this article:
Theorem 1.5.
Let $V\in\mathfrak{X}_{\mathcal{O}}{(\mathbb{C}^{n})}$ $(n\geq 2)$ be a complete globally asymptotic stable vector field with $V(0)=0$ and $DV(0)=A$. Assume that $\Omega\subseteq\mathbb{C}^{n}$ contains the origin and spirallike domain with respect to the vector field $V$. Suppose that there exists $\gamma,\alpha>0$ such that $\|X(t,z)\|\leq\gamma e^{-\alpha t}\|z\|$ for all $t\geq 0$ and for all $z\in\Omega$, where $X(t,z)$ is the flow of the vector field $V$. Assume that
$2\alpha+k_{-}(A)$ $>0.$ (1.3)
Let $\Phi:\Omega\to\Phi(\Omega)$ be a biholomorphism with $\Phi(\Omega)$ is Runge. Then the map $\Phi$ can be approximated by automorphisms of $\mathbb{C}^{n}$ uniformly on every compact subset of $\Omega$.
We obtain the following corollary of the Theorem 1.5.
Corollary 1.6.
Let $\Omega_{1}\subset\mathbb{C}^{n}$, $\Omega_{2}\subset\mathbb{C}^{m}$ be two domains containing the origin and spirallike with respect to $V_{1}\in\mathfrak{X}_{\mathcal{O}}(\mathbb{C}^{n})$ and $V_{2}\in\mathfrak{X}_{\mathcal{O}}(\mathbb{C}^{m})$ respectively. Suppose that $X_{1}(t,z)$, $X_{2}(t,w)$ denotes the flow of the vector field $V_{1}$, $V_{2}$. Assume that there exist $\gamma_{1},\gamma_{2},\alpha_{1},\alpha_{2}>0$ such that $\|X_{j}(t,z_{j})\|\leq\gamma_{j}e^{-\alpha_{j}t}\|z_{j}\|$, for all $t\geq 0$, for all $z_{j}\in\Omega_{j}$ for $j\in\{1,2\}$. Suppose that
$\min\{\alpha_{1},\alpha_{2}\}>\max\{\frac{-k_{-}(A)}{2},\frac{-k_{-}(B)}{2}\}.$
[…]
A holomorphic symplectic form $\omega$ on $\mathbb{C}^{2n}$ is closed holomorphic $2$-form on $\mathbb{C}^{2n}$ whose highest exterior power $\omega^{n}$ is nowhere vanishing. The standard holomorphic symplectic form on $\mathbb{C}^{2n}$ with coordinates $z_{1},z_{2},\cdots z_{n},~{}w_{1},w_{2},\cdots,w_{n}$ is $\omega=\sum_{j=1}^{n}\,dz_{i}\wedge\,dw_{i}$ Let $\Omega\subset\mathbb{C}^{n}$ be a domain. A holomorphic map $f:\Omega\to\mathbb{C}^{n}$ is called symplectic if $f^{*}(\omega)=\omega$. A holomorphic vector field $V$ on $\Omega$ is said to be symplectic if $\,d\iota_{V}(\omega)=0$. An automorphism $F:\mathbb{C}^{n}\to\mathbb{C}^{n}$ is symplectic if $F^{*}(\omega)=\omega$. We denote the set of all symplectic holomorphism by $Aut_{\omega}(\mathbb{C}^{n})$.
We now present an approximation result where a volume preserving biholomorphisms and symplectic biholWe now present an approximation result where a volume preserving biholomorphisms and symplectic biholomorphism that are defined on a spirallike domain, by volume preserving automorphism and symplectic automorphisms of $\mathbb{C}^{n}$ respectively. omorphism that are defined on a spirallike domain, by volume preserving automorphism and symplectic automorphisms of $\mathbb{C}^{n}$ respectively.
Theorem 1.7.
Let $V\in\mathfrak{X}_{\mathcal{O}}{(\mathbb{C}^{n})}$ $(n\geq 2)$ be a complete globally asymptotic stable vector field with $V(0)=0$ and $DV(0)=A$. Assume that $\Omega\subseteq\mathbb{C}^{n}$ is a pseudoconvex domain containing the origin and spirallike domain with respect to $V$. Suppose that there exists $\gamma,\alpha>0$ such that $\|X(t,z)\|\leq\gamma e^{-\alpha t}\|z\|$ for all $t\geq 0$ and for all $z\in\Omega$, where $X(t,z)$ is the flow of the vector field $V$. Assume that
$2\alpha+k_{-}(A)$ $>0.$ (1.4)
Let $\Phi:\Omega\to\Phi(\Omega)$ be a biholomorphism with $\Phi(\Omega)$ is Runge. Then the following holds:
  • (i)
[…]
  • (ii)
    Let $n=2m$ and $(z_{1},z_{2},\ldots,z_{m},w_{1},w_{2},\ldots,w_{m})$ be a coordinate of $\mathbb{C}^{n}$ and $\omega=\sum_{j=1}^{m}\,dz_{i}\wedge\,dw_{i}$ is a symplectic form on $\mathbb{C}^{n}$. If $V$ is symplectic vector field and $\Phi^{*}(\omega)=\omega$. Then $\Phi$ can be approximated by elements of $Aut_{\omega}(\mathbb{C}^{n})$ locally uniformly on $\Omega$.
Remark 1.8.
The condition on the flow given in (1.3), implies the uniform exponential stability of the vector field $V$ on the domain $\Omega$, where the flow of the vector field $V$ has a rate of convergence bigger than $-\frac{k_{-}(A)}{2}$. The question for spirallike domains with respect to a globally asymptotic stable vector field remains open.
Hence, we obtain the following corollary from Theorem 1.5, restricting the vector field to be linear which already gives a generalization to Result 1.4.
Corollary 1.9.
Let $A\in GL(n,\mathbb{C})$ with $2k_{+}(A)<k_{-}(A)$ and $\Omega\subset\mathbb{C}^{n}$ be a domain containing the origin and spirallike with respect to $P^{-1}AP$ for some $P\in GL(n,\mathbb{C})$. Assume that $\Phi\colon\Omega\to\Phi(\Omega)$ is biholomorphism with $\Phi(\Omega)$ Runge. Then:
  • (i)
    The map $\Phi$ can be approximated by elements of $Aut(\mathbb{C}^{n})$ uniformly on every compact subset of $\Omega$.
  • (ii)
    If $\Omega$ is pseudoconvex and $J\Phi\equiv 1$ then $\Phi$ can be approximated by elements of $Aut_{1}(\mathbb{C}^{n})$ uniformly on every compact subsets of $\Omega$.
  • (iii)
    Let $n=2m$ and $(z_{1},z_{2},\cdots,z_{m},w_{1},w_{2},\cdots,w_{m})$ be a coordinate of $\mathbb{C}^{n}$. Assume that $\Omega\subseteq\mathbb{C}^{n}$ is pseudoconvex domain and $\sum_{j=1}^{m}\,dz_{i}\wedge\,dw_{i}$ is a symplectic form on $\mathbb{C}^{n}$. If $A\in Sp(m,\mathbb{C}):=\{A\in Gl(n,\mathbb{C}):A^{t}JA=A\}$, where $J=\begin{bmatrix}0&I_{m}\\ -I_{m}&0\end{bmatrix}$ and $\Phi^{*}(\omega)=\omega$ then $\Phi$ can be approximated by $Aut_{sp}(\mathbb{C}^{n}):=\{\Psi\in Aut(\mathbb{C}^{n}):\Psi^{*}(\omega)=\omega\}$, where $\omega$ is the standard symplectic form on $\mathbb{C}^{n}$.
Remark 1.10.
All the Theorem 1.5, Theorem 1.1, Theorem 1.12 is also true if the vector field $F$ has non zero globally asymptotic stable equilibrium point and $\Omega$ contains the equilibrium point.
Remark 1.11.
The novelty of taking the approach through the vector fields is that the notion of spirallikeness can be defined in case of complex manifolds with density property. A result analogous to Theorem 1.5 can be proved in that case, which we will deal in a different paper.
Before moving to certain applications of the above results in Loewner PDE, we give a few comments about the proofs of the above approximation results. The main idea is to construct an isotopy in Result 1.3. In our proof of Theorem 1.5, we will also construct an isotopy by using the flows. Our next result is an application of Theorem 1.5 to the theory of Loewner PDE. The study was motivated by Arosio-Bracci-Wold [ABFW13] and Hamada [Hamada15]. We present a brief introduction to Loewner PDE before presenting our result. The Loewner ordinary differential equation in the unit disc was introduced by Loewner[Loew1923] in 1923 to study Bieberbach’s conjecture about the coefficients of a univalent map on the unit disc. The theory in one variable was developed later by Kufarev [Kuf] and Pommerenke [Pomm1965] (see [MBCD2010] for a brief survey of Loewner differential equations). The Loewner theory has been used to prove many deep results in geometric function theory [GKbook]. It was one of the main ingredients in de Brange’s proof of Bieberbach conjecture [dB1985]. Then there were several generalizations; Some of them are: chordal Loewner theory [KSS1968] and Loewner theory in several complex variables [Pfal1974, Pfal1975]. Pfaltzgraff [Pfal1974, Pfal1975] extended the theory of the Loewner equation in several complex variables, and later many researchers studied the higher dimensional Loewner ODE and PDE [Poreda1991, GHK2002, GHKK2008, HK2003, GKP2005]. A map $H:\Omega\times[0,\infty)\to\mathbb{C}^{n}$ is called a Herglotz vector field of order $d$ on $\Omega$ if
  • (i)
    For each $z\in\Omega$, $H(z,\cdot):[0,\infty)\to\Omega$ is measurable.
  • (ii)
    For each $s\in[0,\infty)$, $H(.,s)$ is holomorphic vector field and $H(\cdot,s)$ is a semicomplete holomorphic vector field for almost all $s\in[0,\infty)$, i.e. for each $z_{0}\in\Omega$, the ODE defined by
    $\dfrac{dw}{dt}=H(w(t),s),\;\;\;w(0)=z_{0}$
    has a solution for all $t\in[0,\infty)$ for almost all $s\in[0,\infty)$.
  • (iii)
    For each $t_{0}\in[0,\infty)$, and for each compact $K\subset\Omega$, there exists a function $c_{t_{0}}^{K}\in L^{d}([0,t_{0}],[0,\infty))$ such that $||H(z,t)||\leq c_{t_{0}}^{K}$ for all $z\in K$ almost all $t\in[0,\infty]$.
For a Herglotz vector field $H$ of order $d\in[1,\infty]$ on a complete hyperbolic domain $\Omega$ the Loewner PDE is the following equation:
$\dfrac{\partial f_{t}(z)}{\partial t}$ $=-df_{t}(z)H(z,t).$ (1.5)
A solution of the Loewner PDE is a family of holomorphic maps $\{f_{t}\}_{t\in[0,\infty)}$ from $\Omega$ to a complex manifold $M$ which is locally absolutely continuous on $[0,\infty)$ satisfies (1.5) and $t\mapsto f_{t}$ is continuous in the topology of holomorphic maps from $\Omega$ to $M$. Such a family of maps is closely related to evolution family $\{\varphi_{s,t}\}$ that satisfies Loewner ODE (1.6)
$\dfrac{\partial\varphi_{s,t}(z)}{\partial t}$ $=H(\varphi_{s,t}(z),t),~{}~{}a.e~{}~{}t\in[s,\infty)$ (1.6)
in the following sense:
[…]
The main open problem in the field is to solve whether, on a complex complete hyperbolic domain, there is a solution to Loewner PDE with values in $\mathbb{C}^{n}$. Solution of such Loewner PDE with values in certain abstract complex manifold has already been found in [BCD2009]. Recently the connection between the approximation of biholomorphic maps and the solution of Loewner PDE has been found[ABFW13, Hamada15]. As an application of Theorem 1.5 we will show that any Loewner PDE in a complete hyperbolic spirallike domain $\Omega$ with respect to an exponentially stable vector field has a solution with values in $\mathbb{C}^{n}$. We will also provide an application of Theorem 1.9. Both of them generalize the result [Hamada15, Theorem 5.3]. Our generalizations provide a new class of domains where Loewner PDE has a univalent solution with values in $\mathbb{C}^{n}$.
Theorem 1.12.
Let $n\geq 2$. Let $\Omega\subset\mathbb{C}^{n}$ be a complete hyperbolic domain containing the origin and spirallike with respect to complete globally asymptotic stable holomorphic vector field $V\in\mathfrak{X}_{\mathcal{O}}{(\mathbb{C}^{n})}$ with $V(0)=0$ and $DV(0)=A$. Suppose that there exist $\gamma,\alpha>0$ such that $\|X(t,z)\|\leq\gamma e^{-\alpha t}\|z\|$, for all $t\geq 0$ and for all $z\in\Omega$, where $X(t,z)$ is the flow of the vector field $V$. Assume that
$2\alpha+k_{-}(A)$ $>0.$ (1.8)
Let $G:\Omega\times\mathbb{R}^{+}\to\mathbb{C}^{n}$ be a Harglotz vector field of order $d\in[1,\infty]$. Then there exist a Loewner chain $f_{t}:\Omega\to\mathbb{C}^{n}$ of order $d$ which solves the Loewner PDE
$\frac{\partial f_{t}}{\partial t}(z)$ $=-df_{t}(G(z,t))~{}~{}a.e.~{}~{}t\geq 0~{}~{}\forall z\in\Omega.$ (1.9)
Moreover, $R=\cup_{t\geq 0}f_{t}(\Omega)$ is a Runge and Stein domain in $\mathbb{C}^{n}$ and any other solution to (1.9) with values in $\mathbb{C}^{n}$ is of the form $(\Phi\circ f_{t})$ for a suitable map $\varphi:R\to\mathbb{C}^{n}$.
Corollary 1.13.
Let $A\in GL(n,\mathbb{C})$ with $2k_{+}(A)<k_{-}(A)$ and $\Omega\subset\mathbb{C}^{n}$ be a domain containing the origin and spirallike with respect to $P^{-1}AP$ for some $P\in GL(n,\mathbb{C})$. Let $G:\Omega\times\mathbb{R}^{+}\to\mathbb{C}^{n}$ be a Harglotz vector field of order $d\in[1,\infty]$. Then there exist a Loewner chain $f_{t}:\Omega\to\mathbb{C}^{n}$ of order $d$ which solves the Loewner PDE
$\frac{\partial f_{t}}{\partial t}(z)$ $=-df_{t}(G(z,t))~{}~{}a.e.~{}~{}t\geq 0~{}~{}\forall z\in\Omega.$
Moreover, $R=\cup_{t\geq 0}f_{t}(\Omega)$ is a Runge and Stein domain in $\mathbb{C}^{n}$ and any other solution to (1.9) with values in $\mathbb{C}^{n}$ is of the form $(\Phi\circ f_{t})$ for a suitable map $\varphi:R\to\mathbb{C}^{n}$.
Theorem 1.12 and Corollary 1.13 give a partial answer to the open question whether any Loewner PDE in a complete hyperbolic domain in $\mathbb{C}^{n}$ admits a univalent solution in $\mathbb{C}^{n}$. Other partial positive answers are [ABFW13, Hamada15, Duren]. Theorem 1.13 generalizes that of [ABFW13, Hamada15].
In the final section of the paper, we provide several examples of spirallike domains that satisfy the conditions of our theorems. Thus those domains allow to approximate any biholomorphic maps onto a Runge domain by automorphisms of $\mathbb{C}^{n}$. we also provide an example of a complete hyperbolic Hartogs domain which is spirallike with respect to a nonlinear holomorphic vector field but not spirallike with respect to any linear vector field.
About the layout of the paper: In Section 2, we provide basic definitions and collect some results from the literature that we will use in our proofs. In Section 3, a proof of Theorem 1.1 is provided. Section 4 is the heart of this paper. The proof of Theorem 1.5, Theorem 1.7 and Corollary 1.6 are presented there. The proof of Theorem 1.12 is presented in Section 5. Section 6 we provide an example of a spirallike domain that is not spirallike with respect to any linear vector field. Here we present some other examples supporting our theorems.
2. Technical Preliminaries
In this section, we first discuss about the holomorphic vector fields and their flows. Then we make a comparison of spirallike domains that were considered by Hamada in [Hamada15] with ours when the vector field is linear. Let $\Omega\subset\mathbb{C}^{n}$ and $V$ be a holomorphic vector field on $\Omega$. From the local existence and uniqueness theory for systems of ordinary differential equations (see [perko]), it follows that, for every $p\in\Omega$, there exists an open set $U$ in $\Omega$ containing the point $p$ and $\epsilon>0$ such that the equation
$\dot{x}(t)$ $=V(x(t))$
$x(0)$ $=x_{0},$ (2.1)
has unique solution denoted by $X(t,x_{0})$ for every point $x_{0}\in U$ and $t\in(-\epsilon,\epsilon)$. Let $X(t,x_{0})=X_{t}(x_{0})$ be the solution of (2) at time $t\in(-\epsilon,\epsilon)$. In this case, the map $X(,)$ on $(-\varepsilon,\varepsilon)\times U$ defined by $(t,x_{0})\mapsto X(t,x_{0})$ is called the (local) flow of the vector field $V$. We have $X_{t}\circ X_{s}=X_{s+t}$ when both sides are well defined. If $V$ is holomorphic, we consider the holomorphic differential equation associated to (2) defined by
$\dot{x}(z)$ $=\widetilde{V}(x(z))$
$x(0)$ $=z_{0},$ (2.2)
[…]
$\frac{d\phi}{dt}=V(\phi(t));~{}~{}\frac{d\psi}{ds}=i\cdot V(\psi(s)),$
[…]
From [Forst1996, Corollary 2.2] we get that every $\mathbb{R}$-complete holomorphic vector field is $\mathbb{C}$-complete. We will use this crucially in our proofs. We always mean $\mathbb{C}$-complete vector field whenever we say a vector field is complete. If a holomorphic vector field $V\in\mathfrak{X}_{\mathcal{O}}(\mathbb{C}^{n})$ is $\mathbb{R}$-complete (or $\mathbb{C}$-complete), then its associated flow, denoted as $X\colon\mathbb{R}\times\mathbb{C}^{n}\to\mathbb{C}^{n}$ (or $X\colon\mathbb{C}\times\mathbb{C}^{n}\to\mathbb{C}^{n}$), defines a smooth action of the additive group $(\mathbb{R},+)$ (or holomorphic addition $(\mathbb{C},+)$ on $\mathbb{C}^{n}$ through holomorphic automorphisms of $\mathbb{C}^{n}$.
Next, we discuss the the notion of spirallike domain. The notion of a $\Phi$-like domain was introduced by Gurganus in [Gur]. Here we provide the definition of the spirallike domain, a modified definition of a $\Phi$-like domain.
Definition 2.1.
Let $V\in\mathfrak{X}_{\mathcal{O}}(\mathbb{C}^{n})$ be a complete vector field. A domain $\Omega\subset\mathbb{C}^{n}$ is said to be spirallike domain with respect to the vector field $V$ if $X(t,z)\in\Omega$ for all $t\in\mathbb{R}_{\geq 0}$, for all $z\in\Omega$. Here $X(,)$ denotes the flow of the vector field $V$.
Remark 2.2.
The Definition 2.1, is not new. In [Gur], Gurganus introduced the notion of $\Phi$-like domains where $\Phi$ is a holomorphic map. It turns out that a domain $\Omega\subseteq\mathbb{C}^{n}$ containing the origin is spirallike with respect to the globally asymptotic stable vector field $\Phi$ defined on $\Omega$ with $\Phi(0)=0$ and $D\Phi(0)+D\Phi^{*}(0)$ is a negative definite matrix if and only if $\Omega$ is $-\Phi$-like. The viewpoint to look at this notion in the context of a holomorphic vector field is new.
In this paper we will consider the holomorphic vector fields with the following properties:
  • i)
    The vector field is defined on the whole of $\mathbb{C}^{n}$.
  • ii)
    The vector field is $\mathbb{R}$-complete.
  • iv)
    The vector field is globally asymptotically stable or globally exponentially stable.
Since the above properties play a very important role in the statements of the results, we just mention those here briefly before going further in our exposition. We consider the following system of differential equations: Let $E\subset\mathbb{R}^{n}$ be an open set containing the origin and $f:E\to\mathbb{R}^{n}$ is continuously differentiable mappings such that $f(0)=0$
$\frac{dX(t)}{dt}$ $=f(X(t)),~{}~{}X(0)=x_{0}.$ (2.3)
Assume that the solution of the system (2.3) exists for every $t\geq 0$, $\forall x_{0}\in E$.
  • Then origin is said to be a stable equilibrium point of the system (2.3) if for every $\epsilon>0$ there exist $\delta>0$ such that $X(t,x_{0})\in B(0,\epsilon)$ for every $x_{0}\in B(0,\delta)$ for every $t\geq 0$.
  • The origin is said to be globally asymptotically stable equilibrium point if $E=\mathbb{R}^{n}$ , the origin is stable equilibrium point such that $\lim_{t\to\infty}X(t,x_{0})=0$ for all $x_{0}\in\mathbb{R}^{n}$. A vector field is said to be globally asymptotically stable if the origin is a globally asymptotically stable equilibrium point.
  • The origin is said to be the globally exponentially stable equilibrium point of the system (2.3) if $E=\mathbb{R}^{n}$ and there exist constants $c,\gamma>0$ such that $\|X(t,x_{0})\|<c.e^{-\gamma t}\|x_{0}\|$ for every $x_{0}\in\mathbb{R}^{n}$, for all $t\geq 0$. The constant $\gamma$ is called an estimate of the rate of convergence of the flow. A vector field is said to be a globally exponentially stable vector field if the origin is a globally exponentially stable equilibrium point for its flow.
Clearly, every globally exponentially stable equilibrium point is a globally asymptotically stable equilibrium point. A linear vector field is asymptotic stable if and only if it is exponentially stable. Any holomorphic vector field $F$ on $\mathbb{C}^{n}$ can also be seen as a holomorphic map from $F:\mathbb{C}^{n}\to\mathbb{C}^{n}$. For such a vector field $F$, we denote $DF(a)$ by the derivative matrix viewing it as a map.
Remark 2.3.
Let for $A\in M_{n}(\mathbb{C})$, $m(A):=\inf_{\|z\|=1}{\sf Re}\langle Az,z\rangle$. The condition in Result 1.4, becomes $k_{+}(A)<2m(A)$. As we had discussed earlier the differential equation that we consider in this case is:
$\frac{dX}{dt}$ $=AX(t)$ (2.4)
$X(0)$ $=z.$
The differential equation that Hamada [Hamada15] considered is the equation introduced by Gurganus [Gur] :
$\frac{dX}{dt}$ $=-AX(t)$ (2.5)
$X(0)$ $=z.$
If a domain $\Omega\subset\mathbb{C}^{n}$ is spirallike with respect to $A$ (as in the Definition 2.1) if and only if $\Omega$ is spirallike with respect to $-A$ in the sense of Hamada (considering the differential equation as (2.5)). The Runge property does not get affected by this change of definition. But, the approximation of the biholomorphic map of $\Omega$ by automorphisms of $\mathbb{C}^{n}$ get affected very much by this change. Hamada’s condition $k_{+}(A)<2m(A)$ (the condition in Result 1.4) does not remain invariant under the above change of definition. In general $m(A)$ is not a similarity invariant. Hence, the Corollary 1.9 really enlarges the class of domains for which every biholomorphism with Runge image can be approximated by elements of $\text{Aut}(\mathbb{C}^{n})$ locally uniformly on $\Omega$, even if the vector field become linear.
Next, we state a result due to Arosio-Bracci-Hamada-Kohr [ABHK2013] which gives the existence of a solution of Loewner PDE with values in a complex manifold on a complete hyperbolic domain. A manifold is said to be completely hyperbolic if it forms a complete metric space with respect to the Kobayashi distance (see [Kobaya] for details.)
Result 2.4.
[ABHK2013, Theorem 5.2] Let $M$ be a complete hyperbolic complex manifold of dimension $n$. Let $G:M\times\mathbb{R}^{+}\to TM$ be a Herglotz vector field of order $d$. Then there exists a Loewner chain $f_{t}:M\to N$ of order $d$ which solves the Loewner PDE
$\frac{\partial f_{t}}{\partial t}(z)$ $=-df_{t}(G(z,t))~{}~{}a.e.~{}t\geq 0~{}~{}\forall z\in M,$ (2.6)
where $N=\cup_{t\geq 0}f_{t}(M)$ is a complex manifold of dimension $n$ and any other solution to (2.6) with values in a complex manifold $Q$ is of the form $(\Lambda\circ f_{t})$ where $\Lambda:N\to Q$ is holomorphic.
We now state a result by Thai-Duc [Dudoc] which gives a characterization of the complete hyperbolic Hartogs domain. This plays a vital role in our proof of Theorem 6.4.
Result 2.5.
[Dudoc] Let $X$ be a complex space and $\phi$ a plurisubharmonic function on $X$. Assume that for every $x\in X$, there exists an open neighbourhood $U$ of $x$ in $X$ and a sequence $\{h_{j}\}$ of holomorphic functions on $U$ and a sequence $\{c_{j}\}$ of real numbers in the interval $(0,1)$ such that the sequence $\{c_{j}\log|h_{j}|\}$ converges uniformly on compact subsets of $U$ to the function $\phi$. Then the Hartogs domain $\Omega_{\phi}(X)=\{(x,z)\in X\times\mathbb{C}:|z|<e^{-\phi(x)}\}$ is complete hyperbolic if and only if $X$ is complete hyperbolic and $\phi$ is continuous on $X$.
Next, we state a result due to Hamada [Hamada15, Theorem 5.2] which describes a condition under which a pair of complex manifolds is a Runge pair.
Result 2.6.
[Hamada15] Let $M$ be a complete hyperbolic Stein Manifold and $Q$ be a complex manifold of the same dimension. Let $(f_{t}:M\to Q)$ be a Loewner chain of order $d\in[1,\infty]$. Then $(f_{s_{1}}(M),f_{s_{2}}(M))$ is a Runge pair for $0\leq s_{1}<s_{2}$.
The next result due to Lloyd [Lloyd, Theorem 3] will be used to prove Lemma 4.1.
Result 2.7.
[Lloyd, Theorem 3] Let $D$ be a bounded domain in $\mathbb{C}^{n}$. Suppose $f$ and $g$ are two continuous maps from $\overline{D}$ to $\mathbb{C}^{n}$ which are holomorphic on $D$ such that
$\|g(z)\|<\|f(z)\|~{}\forall z\in\partial D.$
[…]
We need the following lemma for the approximation result for the linear case.
Result 2.8.
[SimDav, Page-311] For every $\epsilon>0$ there exist $\rho(\epsilon),m(\epsilon)>0$, such that
$\|e^{-tA}z\|$ $\leq\rho(\epsilon)e^{-(k_{-}(A)-\epsilon)t}\|z\|$
$\|e^{tA}z\|$ $\leq m(\epsilon)e^{(k_{+}(A)+\epsilon)t}\|z\|$
for all $t\geq 0$, and for all $z\in\mathbb{C}^{n}$.
We now state and prove a lemma that will be useful for our proofs.
Lemma 2.9.
Let $\Omega$ be a domain in $\mathbb{C}^{n}$ and $\varphi:\Omega\to\mathbb{C}^{n}$ be a biholomorphic map with $\varphi(\Omega)$ Runge. Assume that $\mu:\Omega\to\mathbb{C}^{n}$ is a map such that $\mu(\Omega)$ is Runge and $\mu(\Omega)\subset\Omega$. Then $\lambda\circ\varphi\circ\mu(\Omega)$ is a Runge domain for every $\lambda\in Aut(\mathbb{C}^{n})$.
Proof.
It is enough to prove that $\varphi\circ\mu(\Omega)$ is Runge. Let $f\in\mathcal{O}(\varphi\circ\mu(\Omega))$ and $K$ be a compact subset of $\varphi\circ\mu(\Omega)$. Hence, $\varphi^{-1}(K)$ is a compact subset of $\mu(\Omega)$. Clearly, $f\circ\varphi\in\mathcal{O}(\mu(\Omega))$. Since $\mu(\Omega)$ is Runge, there exists a sequence of polynomials $\{p_{n}\}$ such that $p_{n}\to f\circ\varphi$ uniformly on $\varphi^{-1}(K)$. Hence, $p_{n}\circ\varphi^{-1}\to f$ uniformly on K. Since $\varphi^{-1}:\varphi(\Omega)\to\Omega$ is a holomorphic and $\varphi(\Omega)$ is Runge, there exists a sequence $\{q_{j}\}$ of polynomial maps from $\mathbb{C}^{n}\to\mathbb{C}^{n}$ such that $q_{j}\to\varphi^{-1}$ uniformly on $K$. By a diagonal-type process, we see that $f$ can be approximated by polynomials uniformly on $K$. Hence, $\varphi\circ\mu(\Omega)$ is Runge. ∎
We now state a result from [Forstbook, Page-110] that will be useful in the context of approximation by volume preserving automorphisms. Recall that, for holomorphic vector field $F=\sum_{j=1}^{n}F_{j}\frac{\partial}{\partial z_{j}}\in\mathfrak{X}_{\mathcal{O}}(\mathbb{C}^{n})$ and standard holomorphic $(n,0)$ form $\omega=\,dz_{1}\wedge\,dz_{2}\wedge\cdots\wedge\,dz_{n}$,
$\text{div}_{\omega}F=\frac{\partial F_{1}}{\partial z_{1}}+\frac{\partial F_{2}}{\partial z_{2}}+\cdots+\frac{\partial F_{n}}{\partial z_{n}}.$
Lemma 2.10.
Let $\Omega\subset\mathbb{C}^{n}$ be an open set and $F:\Omega\to\mathbb{C}^{n}$ be a complete $\mathcal{C}^{1}$- vector field with flow $X(t,z)$. Then
[…]
3. Runge property of spirallike domains
Before proving the main result of this section we state and prove a proposition that will be useful in our proof of Theorem 1.1.
Proposition 3.1.
Let $F\in\mathfrak{X}_{\mathcal{O}}{(\mathbb{C}^{n})}$ be a complete globally asymptotically stable vector field such that $F(0)=0$ and $\Omega$ be a spirallike domain with respect to $F$ containing the origin. Then, for every compact subset $K\subset\mathbb{C}^{n}$, there exists a real number $M_{K}>0$ such that $X(t,z)\in\Omega$ for all $t>M_{K}$ and for all $z\in K$.
Proof.
Let $r>0$ such that $B(0,r)\subset\Omega$. Since the origin is the globally asymptotic stable equilibrium point of $F$, there exists $\delta>0$ such that $X(t,z)\in B(0,r)$ $\forall t\geq 0$ and $\forall z\in B(0,\delta)$. We also have $\lim_{t\to\infty}X(t,z)\to 0$ for every $z\in\mathbb{C}^{n}$. Therefore, for every $z\in K$, there exist $t_{z}>0$, such that $X(t_{z},z)\in B(0,\frac{\delta}{3})$. Hence, from choice of $\delta>0$, we have
$X(t,z)\in B(0,r)\subseteq\Omega,\;\;\forall t\geq t_{z}.$
Let $K\subset\mathbb{C}^{n}$ be a compact subset. For every $z\in K$ there exist $r_{z},t_{z}>0$ such that $X(t,w)\in B(0,\frac{\delta}{3})$, for all $w\in B(z,r_{z})$ and $t\geq t_{z}$. Since $K$ is compact, hence, there exists $m\in\mathbb{N}$ and $\{z_{1},z_{2},\ldots,z_{m}\}\subset K$ such that $K\subset\cup_{j=1}^{m}B(z_{j},r_{z_{j}})$. Let $M_{K}=\max\{t_{z_{1}},t_{z_{2}},\ldots,t_{z_{m}}\}$. Therefore, $X(t,z)\in B(0,r)$, for all $z\in K$ , for all $t>M_{K}$.
We now present the proof of the spirallike domain with respect to a globally asymptotic stable vector field is Runge.
Proof of Theorem 1.1.
Let $f\in\mathcal{O}(\Omega)$ and $K\subset\Omega$ be a given compact subset and $\mu$ is a complex Borel measure on $K$. It is enough to show that if, for every $p\in\mathbb{C}[z_{1},z_{2},\ldots,z_{n}]$,
$\int_{K}p(z_{1},z_{2},\ldots,z_{n})\;\mathrm{d}\mu(z)=0$
then
$\int_{K}f(z)\;\mathrm{d}\mu(z)=0.$
Consider
$V_{K}:=\{e^{-\tau}:\tau\in\mathbb{C},X(\tau,z)\in\Omega,\forall z\in K\}\cup\{0\}\subset\mathbb{C}.$
Since $F$ is a $\mathbb{R}$-complete vector field, hence, using [Forst1996, Corollary 2.2], we get that $F$ is also $\mathbb{C}$-complete. Therefore, $X(\tau,z)$ is well defined for every $\tau\in\mathbb{C}$. We now show that $V_{K}$ is open and starshaped with respect to the origin. If $w\in V_{K}$, then $w=e^{-\tau_{1}}$ for some $\tau_{1}\in\mathbb{C}$ with $X(\tau_{1},z)\in\Omega$ for all $z\in K$. For any $t\in(0,1)$ we have
$tw=e^{\ln{t}}\cdot e^{-\tau_{1}}=e^{-(-\ln{t}+\tau_{1})}.$
Clearly, $-\ln{t}>0$. Since $\Omega$ is a spirallike domain with respect to the vector field $F$, hence, for every $z\in K$, it follows that
$X(-\ln{t}+\tau_{1},z)=X_{-\ln{t}}((X(\tau_{1},z)))\in\Omega.$
Consequently, $tw\in V_{K}$ for all $t\in(0,1)$. Clearly, $tw\in\Omega$ for $t=0,1$. Therefore, $V_{K}$ is starshaped. Hence, it is simply connected. Now we show that $V_{K}$ is an open subset. For that, we first show that $S=\{\tau\in\mathbb{C}|~{}~{}\forall~{}z\in K,~{}X(\tau,z)\in\Omega\}$ is an open subset of $\mathbb{C}$. Let $\tau_{1}\in S$. By definition of $S$ we get that $X(\tau_{1},z)\in\Omega$ for every $z\in K$. Let $\epsilon_{z}>0$ such that $B(X(\tau_{1},z),\epsilon_{z})\subset\Omega$. Since $X$ is a continuous map, hence, there exist $\delta_{z,\tau_{1}}>0$ such that $X(\tau,w)\in B(X(\tau_{1},z),\epsilon_{z})$ whenever $|\tau-\tau_{1}|<\delta_{z,\tau_{1}}$ and $\|w-z\|<\delta_{z,\tau_{1}}$. Since $K$ is compact, hence, we get that $K\subset\cup_{i=1}^{m}B(z_{i},\delta_{z_{i},\tau_{1}})$ for some $z_{1},z_{2},\ldots,z_{m}\in K$. Therefore, choosing $0<\delta_{1}<\min\{\delta_{z_{1},\tau_{1}},\delta_{z_{2},\tau_{1}},\ldots,\delta_{z_{m},\tau_{1}}\}$ we obtain that $B(\tau_{1},\delta_{1})\subset S$. By open mapping theorem, we get that all the points of the form $e^{-\tau}$ are interior points of $V_{K}$. We now show that there exists $r>0$ such that $B(0,r)\subset V_{K}$. Now for every non zero complex number $\tau\in\mathbb{C}$, we have $\tau=e^{-(\ln{\frac{1}{|\tau|}}-i\arg{\tau})},$ where $-\pi<\arg{\tau}\leq\pi$. Hence, we get that $X(\ln(|\frac{1}{\tau}|-i\arg{\tau})),z)=X_{\ln(|\frac{1}{\tau}|)}(X(-i\arg{\tau},z))$. Clearly, the set $K^{\prime}=\{X(-it,z):-\pi\leq t\leq\pi,~{}z\in K\}$ is compact. Since $\ln{|\dfrac{1}{\tau}|}\to\infty$ as $\tau\to 0$, hence, we infer from Proposition 3.1, that there exists $r>0$ such that $X(\ln{\frac{1}{|\tau|}},z^{\prime})\in\Omega$, whenever $|\tau|<r$ and $z^{\prime}\in K^{\prime}$. We also obtain that $X(-i\arg{\tau},z)\in K^{\prime}$ for all $z\in K$. Hence, for any $\tau\in B(0,r)$ and $z\in K$, we have $X(\ln{\frac{1}{|\tau|}},X(-i\arg{\tau},z))\in\Omega$. Therefore, $V_{K}$ is an open subset. Let $\widetilde{V}_{K}=V_{K}\setminus(-\infty,0]$. For each $\tau\in\widetilde{V}_{K}$, define $f_{\tau}:K\to\mathbb{C}$ by $f_{\tau}(z)=f(X(-\ln{\tau},z))$. Here $f(X(-\ln{\tau},z))$ is continuously differentiable with respect to the variable $\tau$ for each $z\in K$. Hence, the function $\phi:\widetilde{V}_{K}\to\mathbb{C}$ defined by
$\phi(\tau)=\int_{K}f_{\tau}(z)\;\mathrm{d}\mu(z)$
is holomorphic on $\widetilde{{V}}_{K}$. Since $K$ is compact in $\mathbb{C}^{n}$, there exist $R>0$ such that $K\subset\overline{B(0,R)}$. In view of Proposition 3.1, we get that $\exists~{}\delta>0$ such that $X(-\ln{\tau},y)\in\Omega$ for all $\tau\in\{\tau\in\mathbb{C}:|\tau|<\delta\}\cap\widetilde{{V}}_{K}$ and for all $y\in\overline{B(0,R)}$. Therefore, $f(X(-\ln{\tau},z))\in\mathcal{O}(B(0,R))$ for every $\tau\in\{|t|<\delta\}\cap\widetilde{{V}}_{K}$. Hence, there exists a sequence of polynomial $p_{n}$ such that $p_{n}$ uniformly converges to $f_{\tau}$ on every compact subset of $B(0,R)$, particularly on $K$. Therefore, for every $\tau$ in $V_{K}$ with $|\tau|<\delta$, it follows that
$\lim_{n\to\infty}\int_{K}p_{n}(z)\;\mathrm{d}\mu(z)=\int_{K}f((X(-\ln{\tau},z)))\;\mathrm{d}\mu(z).$
By assumption $\int_{K}p_{n}(X(-\ln{\tau},z))\;\mathrm{d}\mu(z)=0$. Hence, for all $\tau\in V_{K}$ with $|\tau|<\delta$, we obtain that
$\int_{K}f((X(-\ln{\tau},z)))\;\mathrm{d}\mu(z)=0.$
Hence, $\phi$ vanishes identically on an open ball in $\widetilde{V}_{K}$. Since $\phi$ is a holomorphic map hence, it vanishes identically on $\widetilde{V}_{K}$. Therefore, $\phi(1)\equiv 0$ and this implies that $\int_{K}f(z)\;\mathrm{d}\mu(z)=0$. Hence, $\Omega$ is Runge. ∎
4. Approximation of biholomorphisms
We begin this section with the following lemma which will be used in the proof of Theorem 1.5.
Lemma 4.1.
Let $\Omega$ be a domain in $\mathbb{C}^{n},n\geq 2,$ containing the origin. Assume that $\Omega$ is a spirallike domain with respect to a complete globally asymptotic stable vector field $V\in\mathfrak{X}_{\mathcal{O}}{(\mathbb{C}^{n})}$ with $V(0)=0$ and $DV(0)=A$. Suppose that there exists $\gamma,\alpha>0$ such that
  • (i)
    $\|X(t,z)\|\leq\gamma e^{-\alpha t}\|z\|$ for all $t\geq 0$ and for all $z\in\Omega$, where $X(t,z)$ is the flow of the vector field $V$.
  • (ii)
    $02\alpha+k_{-}(A)>$ 0 .
Then there exists a biholomorphic map $F$ on $\Omega$ into $\mathbb{C}^{n}$ such that $e^{-tA}\circ X_{t}(z)$ converges to $F(z)$ as $t\to\infty$ uniformly on compact subsets of $\Omega$.
Proof.
Fix a compact set $K\subset\Omega$. We will show that for every $\varepsilon>0$, there exists $N_{\varepsilon,K}\in\mathbb{N}$ such that $\|e^{-t_{1}A}X_{t_{1}}(z)-e^{-t_{2}A}X_{t_{2}}(z)\|<\varepsilon$ for every $t_{1},t_{2}>N_{\varepsilon,K}$ and for all $z\in K$. This ensures the uniform convergence of $\{e^{-tA}\circ X_{t}\}$ on $K$ as $t\to\infty$. Since $V$ is a holomorphic vector field on $\mathbb{C}^{n}$, we can write $V(z)=Az+R(z)$, where $R(z)=O(\|z\|^{2})$ as $z\to 0$ i.e., there exist $M,\delta>0$ such that $\|R(z)\|\leq M\|z\|^{2}$ for all $z\in B(0,\delta)$. Since $V$ is globally asymptotically stable vector field on $\mathbb{C}^{n}$, for every $z\in\mathbb{C}^{n}$ $\|X(t,z)\|\to 0$ as $t\to\infty$. We note that the convergence is uniform on every compact subset of $\mathbb{C}^{n}$. Hence, there exists $\widetilde{N_{K}}>0$ such that
$\|R(X(t,z))\|$ $<M\|X(t,z)\|^{2},\forall t>\widetilde{N_{K}},\forall z\in K.$ (4.1)
Define $f_{t}:\mathbb{C}^{n}\to\mathbb{C}^{n}$ by $f_{t}(z):=e^{-tA}X_{t}(z)$. Clearly, $f_{t}\in Aut(\mathbb{C}^{n})$ for all $t>0$. The family is also smooth in $t$-variable. Using the mean value theorem in the $t$ variable, we obtain that
$\|f_{t}(z)-f_{s}(z)\|$ $\leq\left\|\frac{df_{t}(z)}{\,dt}\bigg{|}_{t=s+a_{s}}\right\||t-s|,$ (4.2)
where $s\leq s+a_{s}\leq t$. We now find an estimate for $\left\|\frac{df_{t}(z)}{\,dt}\right\|$.
$\frac{df_{t}(z)}{\,dt}$ $=-Ae^{-tA}(X(t,z)+e^{-tA}(A(X(t,z)))+e^{-tA}(R(X(t,z)))$
$=e^{-tA}R(X(t,z)),$
Hence,
$\|\frac{df_{t}(z)}{\,dt}\bigg{|}_{s+a_{s}}\|$ $=\|e^{-(s+a_{s})A}R(X((s+a_{s}),z))\|.$ (4.4)
By assumption, choose $\sigma>0$ such that $02\alpha+k_{-}(A)-\sigma>$ 0. In view of Result 2.8, we deduce that
$\|\frac{df_{t}(z)}{\,dt}\bigg{|}_{s+a_{s}}\|$ $\leq\rho(\varepsilon_{1})e^{-(s+a_{s})(k_{-}(A)-\sigma)}\|R(X((s+a_{s}),z))\|.$ (4.5)
Using (4.1) and (4.5), we get that
$\|\frac{df_{t}(z)}{\,dt}\bigg{|}_{s+a_{s}}\|$ $\leq M\rho(\sigma)e^{-(s+a_{s})(k_{-}(A)-\sigma)}\|(X((s+a_{s}),z)\|^{2}\;\;\forall s>\widetilde{N_{K}}\;\;\forall z\in K.$ (4.6)
Using Assumption (i), equations (4.2), (4.1) and (4.6) we obtain that
$\|f_{t}(z)-f_{s}(z)\|\leq M^{\prime}e^{-Cs}|t-s|\;\;\forall t>s>\widetilde{N_{K}}\;\;\forall z\in K,$ (4.7)
where $C=2\alpha+k_{-}(A)-\sigma>0$ and $M^{\prime}:=\sup_{z\in K}M\gamma^{2}\rho(\varepsilon_{1})\|z\|^{2}$. Since there is an exponential decay on the right hand side of Equation 4.7, it follows that for every $\varepsilon>0$, there exists $N_{\varepsilon,K}\in\mathbb{N}$ such that $\|e^{-t_{1}A}X_{t_{1}}(z)-e^{-t_{2}A}X_{t_{2}}(z)\|<\varepsilon$ for every $t_{1},t_{2}>N_{\varepsilon,K}$ and for all $z\in K$. This proves that every subsequence of $\{f_{t}\}$ forms a Cauchy sequence on every compact subset of $\Omega$. Hence, there exists $F\in\mathcal{O}(\Omega,\mathbb{C}^{n})$ such that $\lim_{t\to\infty}e^{-tA}\circ X_{t}(z)=F(z)$ and the convergence is uniform on every compact subsets of $\Omega$.
Next, we show that $F$ is a biholomorphic map from $\Omega$ onto its range. We have $f_{t}=e^{-tA}\circ X_{t}\in\text{Aut}(\mathbb{C}^{n})$ and $D(e^{-tA}X_{t})(0)=e^{-tA}DX_{t}(0)=e^{-tA}e^{tA}=I_{n}$ for all $t\geq 0$. Since $Df_{t}(0)\to DF(0)$ as $t\to\infty$, hence, we obtain that $DF(0)=I_{n}$. Since $\text{det}(D(e^{-tA}X_{t})(z))\neq 0$ for all $z\in\mathbb{C}^{n}$, therefore, by Hurwitz theorem in several variables, we conclude that $\text{det}(DF(z))$ is either non-zero on $\Omega$ or identically zero on $\Omega$. But, $\text{det}(DF(0))=1$ implies that $\det(DF(z))\neq 0$ for all $z\in\Omega$. Therefore, by invoking inverse function theorem, we obtain that $F$ is a local biholomorphism. We need to show that $F$ is injective. Suppose not. Then, there exist two distinct points $z_{1},z_{2}$ in $\Omega$ such that $F(z_{1})=F(z_{2})$. Since $F$ is locally injective, hence, there exists $r>0$ such that $F$ is injective on $B(z_{1},r)$. Shrinking $r>0$, if needed, we assume that $z_{2}\notin\overline{B(z_{1},r)}$. Define $\widetilde{F}_{t}:=f_{t}-f_{t}(z_{2})$ and $\widetilde{F}:=F-F(z_{2})$. Clearly, $\widetilde{F}^{-1}(0)\cap\partial B(z_{1},r)=\emptyset$. Choose $\alpha>0$ such that $\inf_{\partial B(z_{1},r)}|\widetilde{F}|>\alpha$.
Now consider $g_{t}:=\widetilde{F}_{t}-\widetilde{F}$. We have
$\sup_{\partial B(z_{1},r)}|g_{t}(z)|\leq\sup_{\partial B(z_{1},r)}\left(|f_{t}(z)-F(z)|+|f_{t}(z_{2})-F(z_{2})|\right).$
Hence, there exists $N_{0}\in\mathbb{N}$ such that $\sup_{\partial B(z_{1},r)}|g_{t}(z)|\leq\alpha$ for all $t>N_{0}$. Hence, we obtain that
$\sup_{\partial B(z_{1},r)}|g_{t}(z)|\leq\alpha<|\widetilde{F}(z)|\;\;\forall z\in\partial B(z_{1},r).$
Therefore, we infer from Result 2.7, that $\widetilde{F}$ and $\widetilde{F_{t}}$ has same number of zeros on $B(z_{1},r)$ for all $t>N_{0}$. That contradicts the fact that $f_{t}$ are injective. Therefore, $F:\Omega\to\mathbb{C}^{n}$ is an injective holomorphic map. ∎
We now present the proof of the main theorem of this paper Theorem 1.5
Proof of Theorem 1.5.
Here $V$ is a complete holomorphic vector field, $X\colon\mathbb{R}\times\mathbb{C}^{n}\to\mathbb{C}^{n}$ is the flow of the vector field $V$. Let $\Phi:\Omega\to\Phi(\Omega)$ be a biholomorphism such that $\Phi(\Omega)$ is Runge. Define a map $H\colon[0,1]\times\Omega\to\mathbb{C}^{n}$ by
$H(t,z):=\begin{cases}e^{c\ln{t}A}(\Phi(X(-c\ln{t},z))),~{}\text{if}~{}t\in(0,1]\\ F(z),~{}\text{if}~{}t=0\\ \end{cases},$ (4.8)
where $F:\Omega\to\mathbb{C}^{n}$ is a biholomorphism onto its image is chosen in such way $e^{-c\ln{t}A}\circ X_{-c\ln{t}}(z)$ converges to the map $F(z)$ as $t\to 0^{+}$ uniformly over every compact subset of $\Omega$. In view of Lemma 4.1, we deduce the existence of such a biholomorphism $F\colon\Omega\to\mathbb{C}^{n}$. The biholomorphism $F$ is a limit of a sequence of automorphisms of $\mathbb{C}^{n}$ on the domain $\Omega$. According to the assumption, there exists $\epsilon_{1}>0$ such that $02\alpha+k_{-}(A)-\epsilon_{1}>$ 0. We choose a real number $c>\frac{1}{2\alpha+k_{-}(A)-\epsilon_{1}}>0$ and fix it. We will first show that $H$ is a $\mathcal{C}^{1}$-map and for each fixed $t\in[0,1]$, $H_{t}(\Omega)$ is Runge. Then, we will use the Result 1.3. It is enough to show that $H$ is $\mathcal{C}^{1}$ at the points of the form $(0,z)$. Since any translation map and an invertible linear map is particularly an automorphism of $\mathbb{C}^{n}$, hence, without loss of generality, we assume that $\Phi(0)=0$ and $D\Phi(0)=I_{n}$. Let $K\subset\Omega$ be a compact set. We now divide the proof into four steps.
Step I: Showing that $H$ is continuous at $(0,\zeta)$.
Expanding the power series expansion of the biholomorphism $\Phi:\Omega\rightarrow\mathbb{C}^{n}$ around the origin, we obtain that there exist constants $M_{1},\delta>0$ such that $\Phi$ can be expressed as:
$\Phi(z)$ $=z+R_{1}(z)\quad\text{for all}~{}z\in B(0,\delta),$ (4.9)
where
$\|R_{1}(z)\|\leq M_{1}\|z\|^{2}~{}\text{for all}~{}z\in B(0,\delta).$ (4.10)
Clearly, $-c\ln{t}\to\infty$ as $t\to 0^{+}$. Since by the assumption, the flow of the vector field $V$ has uniform exponential decay on the domain $\Omega$, hence, there exists $\delta_{K}>0$ such that $X(-c\ln{t},z)\in B(0,\delta)$, for all $z\in K$ and for all $t\in(0,\delta_{K})$. Hence, by (4.10), we obtain
$\|R_{1}(X(-c\ln{t},z))\|$ $\leq M_{1}\|X(-c\ln{t},z)\|^{2}~{}\forall z\in K,~{}\forall t\in(0,\delta_{K}).$ (4.11)
By assumption we obtain
$\|X(-c\ln{t},z)\|$ $\leq\gamma e^{\alpha c\ln{t}}\|z\|~{}\forall t>0.$ (4.12)
Hence, from (4.11), (4.12) we get that
$\|R_{1}(X(-c\ln{t},z))\|$ $\leq M_{1}\gamma^{2}t^{2c\alpha}\|z\|^{2}~{}\forall z\in K,~{}\forall t\in(0,\delta_{K}).$ (4.13)
By using Result 2.8 and (4.13), we deduce that for all $z\in K$ and for all $t\in(0,\delta_{K})$ the following holds:
$\|e^{c\ln{t}A}(R_{1}(X(-c\ln{t},z)))\|$ $\leq\rho(\varepsilon_{1})e^{c\ln{t}(k_{-}(A)-\varepsilon_{1})}\|R_{1}(X(-c\ln{t},z))\|$
$\leq M_{1}\rho(\varepsilon_{1})\gamma^{2}e^{c\ln{t}(k_{-}(A)-\varepsilon_{1})}t^{2c\alpha}\|z\|^{2}$
$=M_{1}\rho(\varepsilon_{1})\gamma^{2}t^{c(k_{-}(A)+2\alpha-\varepsilon_{1})}\|z\|^{2}$
$=M^{\prime}t^{c(k_{-}(A)+2\alpha-\varepsilon_{1})}.$ (4.14)
Here $M^{\prime}:=\sup_{z\in K}M_{1}\rho(\varepsilon_{1})\gamma^{2}\|z\|^{2}$. Since from (1.3) we have $k_{-}(A)+2\alpha-\varepsilon_{1}>0$, hence, from (4), we conclude that $e^{c\ln{t}A}(R_{1}(X(-c\ln{t},z)))\to 0$ as $t\to 0^{+}$, uniformly $K$. Now for any $(t,z)\in(0,\delta_{K})\times K$ we obtain
$H(t,z)$ $=e^{c\ln{t}A}(\Phi(X(-c\ln{t},z)))$
$=e^{c\ln{t}A}\big{(}X(-c\ln{t},z)+R_{1}(X(-c\ln{t},z))\big{)}$
$=e^{c\ln{t}A}(X(-c\ln{t},z))+e^{c\ln{t}A}\big{(}R_{1}(X(-c\ln{t},z))\big{)}.$ (4.15)
Therefore,
$\lim_{t\to 0^{+}}H(t,z)=\lim_{t\to 0^{+}}e^{c\ln{t}A}(X(-c\ln{t},z)).$
Since by Lemma 4.1, we have $\lim_{t\to 0^{+}}e^{c\ln{t}A}\circ(X(-c\ln{t},z))=F(z)$ and the convergence is uniform on every compact subset of $\Omega$, therefore, $\lim_{t\to 0^{+}}H(t,z)=F(z)=H(0,z)$. Hence, $H$ is continuous at $(0,z)$.
Step II: Showing that $H$ is differentiable at $(0,z)$.
We want to show that $DH(0,z)=DF(z)$ for every $z\in\Omega$. For every $t\in(0,\delta_{K})$ , $z\in K$ , and $h\neq 0$ inside a small enough neighborhood of origin we have
$\|H(t,z+h)-H(0,z)-DF(z)h\|$ $=\|e^{c\ln{t}A}(\Phi(X(-c\ln{t},z+h)))-F(z)-DF(z)h\|,$
using the expression of the map $\Phi$ from (4.9), we get from the above equation
$\|H(t,z+h)-H(0,z)-DF(z)h\|$ $=\|e^{c\ln{t}A}(X(-c\ln{t},z)+R_{1}(X(-c\ln{t},z)))-F(z)-DF(z)h\|$
$\leq\|e^{c\ln{t}A}(X(-c\ln{t},z+h))-F(z+h)\|$
$+\|e^{c\ln{t}A}\big{(}R_{1}(X(-c\ln{t},z+h)\big{)}+\|F(z+h)-F(z)-DF(z)h\|.$ (4.16)
Hence, from (4), we have the following for $(t,z)\in(0,\delta_{K})\times K$ and $h\neq 0$ inside a small enough neighborhood of origin
$\frac{\|H(t,z+h)-H(0,z)-DF(z)h\|}{\|(t,h)\|}$
$\leq\frac{\|e^{c\ln{t}A}(X(-c\ln{t},z+h))-F(z+h)\|}{\|(t,h)\|}+\frac{\|e^{c\ln{t}A}\big{(}R_{1}(X(-c\ln{t},z+h)\big{)}\|}{\|(t,h)\|}$
$+\frac{\|F(z+h)-F(z)-DF(z)h\|}{\|(t,h)\|}$
$\leq\frac{\|e^{c\ln{t}A}(X(-c\ln{t},z+h))-F(z+h)\|}{t}+\frac{\|e^{c\ln{t}A}\big{(}R_{1}(X(-c\ln{t},z+h)\big{)}\|}{t}$
$+\frac{\|F(z+h)-F(z)-DF(z)h\|}{\|h\|}.$ (4.17)
Next, we show that the three summands that appeared on the right-hand side of the (4), tend to zero as $\|(t,h)\|\to(0,0)$ with $t>0$.
Clearly,
$\frac{\|F(z+h)-F(z)-DF(z)h\|}{\|h\|}\to 0~{}~{}\text{as}~{}~{}\|h\|~{}~{}\to 0.$ (4.18)
By similar computation as (4), we conclude that there exists $M^{\prime\prime}>0$, such that for every $(t,h)\in(0,1)\times\mathbb{C}^{n}$ with $\|(t,h)\|$ is small enough, the following holds:
$\frac{\|e^{c\ln{t}A}\big{(}R_{1}(X(-c\ln{t},z+h)\big{)}\|}{t}$ $\leq M^{\prime\prime}t^{c(2\alpha+k_{-}(A)-\varepsilon_{1})-1},~{}~{}\forall z\in K.$ (4.19)
From the choice of $c>0$, it follows that the right-hand side of the (4.19) goes to 00 as $\|(t,h)\|\to 0$ with $t>0$. We now deal with the first summand of (4). We now view the vector field $V$ as an entire mapping such that $V(0)=0$, $DV(0)=A$. Hence, we can write
$V(z)$ $=Az+R(z).$ (4.20)
Define a map $S\colon[0,1]\times\Omega\to\mathbb{C}^{n}$ by
$S(t,z)=\begin{cases}e^{c\ln{t}A}\circ X_{-c\ln{t}}(z)~{}\quad\text{if}~{}t\in(0,1]\\ F(z)~{}\quad\text{if}~{}t=0\\ \end{cases}.$
The first summand in (4) is
$\frac{S(t,z+h)-S(0,z+h)}{t}.$ (4.21)
It follows from Lemma 4.1 that the map $t\to S(t,z+h)$ is continuous on $[0,1]$. It is also differentiable on $(0,1)$. Hence, by using the mean value theorem in (4.21), we obtain that
$\frac{\|S(t,z+h)-S(0,z+h)\|}{t}$ $\leq\bigg{\|}\frac{dS(t,z)}{dt}\bigg{|}_{t=\tau}\bigg{\|}.$ (4.22)
Here we do some computation. For $t\in(0,1)$, we have
$\frac{dS(t,z)}{dt}$ $=\frac{d}{dt}(e^{c\ln{t}A}\circ X_{-c\ln{t}}(z))$
$=\frac{c}{t}Ae^{c\ln{t}A}\big{(}X(-c\ln{t},z)\big{)}+e^{c\ln{t}A}\bigg{(}\frac{dX(-c\ln{t},z)}{dt}\bigg{)}$
$=\frac{c}{t}Ae^{c\ln{t}A}\big{(}X(-c\ln{t},z)\big{)}-\frac{c}{t}e^{c\ln{t}A}(V(X(-c\ln{t},z))).$
Since we have $V(z)=A(z)+R(z)$, hence, from (4) we get
$\frac{dS(t,z)}{dt}$ $=\frac{c}{t}Ae^{c\ln{t}A}\big{(}X(-c\ln{t},z)\big{)}-\frac{c}{t}e^{c\ln{t}A}(A(X(-c\ln{t},z)))-\frac{c}{t}Ae^{c\ln{t}A}\bigg{(}R(X(-c\ln{t},z))\bigg{)}$
$=-\frac{c}{t}Ae^{c\ln{t}A}\big{(}R(X(-c\ln{t},z))\big{)}.$ (4.24)
From (4.21) and (4) we deduce that
$\frac{\|S(t,z+h)-S(0,z+h)\|}{t}$ $\leq\bigg{\|}\frac{dS(t,z+h)}{dt}\bigg{|}_{t=\tau}\bigg{\|},$
where $0<\tau<t$. Therefore, we obtain
$\frac{\|S(t,z+h)-S(0,z+h)\|}{t}$ $\leq\frac{c}{\tau}\bigg{\|}Ae^{c\ln{\tau}A}\big{(}R(X(-c\ln{\tau},z+h))\big{)}\bigg{\|}.$ (4.25)
There is a neighborhood $U$ of the origin such that $\|R(z)\|\leq M\|z\|^{2}$ for all $z\in U$. Since $X(-c\ln{t},z)\to 0$ as $t\to 0^{+}$ uniformly on every compact subset of $\Omega$. Hence, from Result 2.8 and the equation (4.25), we obtain that
$\frac{\|S(t,z+h)-S(0,z+h)\|}{t}$ $\leq M\gamma^{2}\rho(\varepsilon_{1})t^{c(2\alpha+k_{-}(A)-\varepsilon_{1})-1}\|z+h\|^{2}.$ (4.26)
By the choice of $c>0$, it follows from (4.26) that
$\frac{\|e^{c\ln{t}A}(X(-c\ln{t},z+h))-F(z+h)\|}{t}\to 0~{}\text{as}~{}t\to 0^{+}~{}\text{uniformly on $K$.}$
Therefore, we infer from (4.26), (4.19) and (4.18) that left-hand side of (4) tends to 00 as $\|(t,h)\|\to 0$ uniformly on $K$. Consequently, it proves that the map $H$ is differentiable at $(0,z)$.
Step III: Showing that $H$ is continuously differentiable at $(0,z)$.
We show that $\frac{dH(t,z)}{dt}\to 0$ uniformly over $K$ as $t\to 0^{+}.$ We now compute the derivative.
$\frac{dH(t,z)}{dt}$ $=\frac{d}{dt}(e^{c\ln{t}A}(\Phi(X(-c\ln{t},z))))$
$=\frac{c}{t}e^{c\ln{t}A}\big{(}A\Phi(X(-c\ln{t},z))-D\Phi(X(-c\ln{t},z))V(X(-c\ln{t},z))\big{)}$
$=\frac{c}{t}e^{c\ln{t}A}\big{(}\Psi(X(-c\ln{t},z))\big{)}.$ (4.27)
Here $\Psi(z)=A(\Phi(z))-D\Phi(z)V(z)$. Clearly, $\Psi(0)=0$. We prove that $D\Psi(0)=0$. We deduce the following for all non-zero $h\in\mathbb{C}^{n}$
$\|\Psi(h)\|$ $=\|A(\Phi(h))-D\Phi(h)V(h)\|$
$=\|A(\Phi(h))-D\Phi(h)(Ah+R(h))\|$
$=\|A(\Phi(h))-Ah+A(h)-D\Phi(h)Ah-D\Phi(h)R(h))\|$
$\leq\|A(\Phi(h))-Ah\|+\|A(h)-D\Phi(h)Ah\|+\|D\Phi(h)R(h))\|$
$\leq\|Ah+A(R_{1}(h))-Ah\|+\|A(h)-D\Phi(h)Ah\|+\|D\Phi(h)R(h))\|$
$\leq\|A\|\|(R_{1}(h))\|+\|I-D\Phi(h)\|\|Ah\|+\|D\Phi(h)R(h))\|.$ (4.28)
We have $\|R_{1}(h)\|\leq M_{1}\|h\|^{2}$ and $\|R(h)\|\leq M\|h\|^{2}$ for all $z\in U\cap B(0,\delta^{{}^{\prime}})$. Also, $D\Phi(h)\to I$ as $h\to 0$. Hence, from (4) we get that $\frac{\|\Psi(h)\|}{\|h\|}\to 0$ as $\|h\|\to 0.$ This proves that $D\Psi(0)=0$. Therefore, we conclude that there exist $M_{2},\delta^{\prime}>0$, such that
$\|\Psi(z)\|\leq M_{2}\|z\|^{2},~{}~{}\forall z\in B(0,\delta^{\prime}).$
We note that $X(-c\ln{t},z)\to 0$ as $t\to 0^{+}$ uniformly on $K$. Hence, by invoking Result 2.8, we conclude that there exists $\delta^{\prime}_{K}>0$ such that
$\bigg{\|}\frac{dH(t,z)}{dt}\bigg{\|}$ $\leq c\rho(\varepsilon)M_{2}\gamma^{2}t^{c(2\alpha+k_{-}(A)-\varepsilon_{1})-1}\|z\|^{2}~{}~{}\forall t\in(0,\delta_{K}^{\prime}),~{}\forall z\in K$ (4.29)
It follows from the choice of $\varepsilon_{1},c>0$ that $\bigg{\|}\frac{dH(t,z)}{dt}\bigg{\|}\to 0$ as $t\to 0^{+}$ uniformly on $K$. Since $H_{t}$ is holomorphic on $\Omega$, hence $DH_{t}(z)\to DH_{0}(z)=DF(z)$ uniformly on $K$ as $t\to 0^{+}$. This proves that $H$ is $\mathcal{C}^{1}$-smooth map at $(0,z)$.
Step IV: Showing that $H_{t}(\Omega)$ is Runge for $t\in[0,1]$.
Since both $X_{-c\ln t}$ and $X_{c\ln t}$ are automorphisms of $\mathbb{C}^{n}$, it follows from Lemma 2.9, that $H_{t}(\Omega)$ is Runge. Hence, we are done. ∎
Proof of Corollary 1.6.
Define a vector field $V$ on $\mathbb{C}^{m+n}$ by
$V(z,w):=(V_{1}(z),V_{2}(w))\quad(z,w)\in\mathbb{C}^{n}\times\mathbb{C}^{m}.$
Let $X_{1}(t,z)$ and $X_{2}(t,w)$ be the flows of the vector field $V_{1}$ and $V_{2}$ respectively. Then $\theta(t,(z,w))=(X_{1}(t,z),X_{2}(t,w))$ is the flow of the vector field $V\in\mathfrak{X}_{\mathcal{O}}(\mathbb{C}^{n+m})$. Clearly, if $(z,w)\in\Omega_{1}\times\Omega_{2}$, then $\theta_{t}(z,w)\in\Omega_{1}\times\Omega_{2}$ for every $t\geq 0$. Hence, $\Omega_{1}\times\Omega_{2}$ is spirallike with respect to the vector field $V$. From the assumption we have that
$2\alpha_{1}+k_{-}(A)$ $>0$
$2\alpha_{2}+k_{-}(B)$ $>0.$
Here $A=DV_{1}(0)$ and $B=DV_{2}(0)$.
Therefore, we obtain
$\|X_{1}(t,z)\|^{2}+\|X_{2}(t,w)\|^{2}$ $\leq\gamma_{1}^{2}e^{-2\alpha_{1}t}\|z\|^{2}+\gamma_{2}^{2}e^{-2\alpha_{2}t}\|w\|^{2}$
$\|\theta(t,(z,w))\|^{2}$ $\leq\tilde{\gamma}^{2}e^{-2\alpha t}\|(z,w)\|^{2}$
$\|\theta(t,(z,w))\|$ $\leq\tilde{\gamma}e^{-\alpha t}\|(z,w)\|,$ (4.30)
where $\tilde{\gamma}:=\max\{\gamma_{1},\gamma_{2}\}$ and $\alpha:=\min\{\alpha_{1},\alpha_{2}\}$. From the assumption we have $\alpha+\min\{\frac{k_{-}(A)}{2},\frac{k_{-}(B)}{2}\}>0$.
Therefore, the rest of the proof follows from Theorem 1.5. ∎
Proof of Theorem 1.7.
(i) Suppose that $\Phi:\Omega\to\Phi(\Omega)$ is a biholomorphism with $J\Phi(z)\equiv 1$ and $\Phi(\Omega)$ is Runge domain. According to our assumption, $div_{\omega}(V)$ is a non-zero constant. Let $div_{\omega}(V)=a\neq 0$. By Lemma 2.10, we get that
$\frac{d}{dt}\det D_{z}(X(t,z))$ $=a\det~{}D_{z}(X(t,z)),~{}\text{det}D_{z}(X(0,z))=1,$
where $D_{z}(X(t,z))$ denotes the derivative matrix of the automorphism $X_{t}\in\text{Aut}(\mathbb{C}^{n})$ with respect to $z$. Solving the above differential equation, we obtain
$\text{det}D_{z}(X(t,z))$ $=e^{at},\;\;\forall z\in\mathbb{C}^{n},\;\;\forall t\geq 0.$ (4.31)
We know that $\det D_{z}X_{t}(0)=e^{tr(A)t}$, where $A=DV(0)$. Hence, it follows from (4.31), $a=tr(A)$. We get from (4.8)
$D_{z}H_{t}(z)=e^{c\ln{t}A}\circ D_{z}(\Phi(X(-\ln{t},z)))\circ D_{z}X_{-\ln{t}}(z).$
Since $\text{det}(D\Phi(X(-c\ln{t},z)))\equiv 1$, hence, from (4.31) we get that
$\det D_{z}H_{t}(z)=e^{tr(A)\ln{t}-tr(A)\ln{t}}=1,~{}~{}\forall t\in(0,1).$
Since $H_{t}(z)\to F(z)$ as $t\to 0^{+}$, hence we have $\text{det}H_{t}(z)\to\text{det}F(z)$ as $t\to 0^{+}.$ Therefore $\text{det}F(z)\equiv 1$ on $\Omega$.
We now show that identity map $\iota_{\Omega}:\Omega\to\Omega$ is homotopic equivalent to a constant map. Define $\widetilde{H}:[0,1]\times\Omega\to\mathbb{C}^{n}$ by
$\widetilde{H}(t,z)=\begin{cases}X(-\ln{t},z)~{}\text{if}~{}t\in(0,1]\\ 0~{}\text{if}~{}t=0\\ \end{cases}.$
The global exponential growth of the flow map provides the continuity of the map $\widetilde{H}(t,z)$ at the point $(0,z)$. This proves that the domain $\Omega$ is contractible to 00. Therefore, $H^{(n-1)}(\Omega,\mathbb{C})$ vanishes identically. Since $\Omega$ is also pseudoconvex, therefore, by Result 1.3, we conclude that $H_{1}(z)=\Phi(z)$ can be approximated by $Aut_{1}(\mathbb{C}^{n})$ uniformly over every compact subset of $\Omega$.
(ii) The flow of the symplectic vector field preserves the symplectic form. From our assumption $\Phi:\Omega\to\Phi(\Omega)$ is a symplectic holomorphic map. Hence, $Z_{t}^{*}(\omega)=\omega$ for all $t\in(0,1]$. We have seen in the proof of (ii) that $\Omega$ is contractible. Hence, $H^{(n-1)}(\Omega,\mathbb{C})=0$. We get that $Z_{t}(\Omega)$ is Runge for every $t\in[0,1]$, from Step IV of the proof of Theorem 1.5. Now using [symfor1995, Proposition 2.3 and Remark] we obtain our theorem.
5. Application in Loewner PDE
In this section first we state a generalization of [ABFW13, Theorem 3.4] and [Hamada15, Theorem 4.5].
Theorem 5.1.
Let $\Omega$ be a complex manifold of dimension $N\geq 2$. Let $\{\Omega_{j}\}_{j\in\mathbb{N}\cup\{0\}}$ be a sequence of open connected subset of $\Omega$ such that $\Omega_{j}\subset\Omega_{j+1}$ for every $j\in\mathbb{N}$ and $\Omega=\cup_{j\in\mathbb{N}\cup\{0\}}\Omega_{j}$. Assume that
  1. (i)
    each pair $(\Omega_{j},\Omega_{j+1})$ is a Runge pair.
  2. (ii)
    each $\Omega_{j}$ is biholomorphic to a Stein spiralshapelike domain with respect to some $V_{j}\in\mathfrak{X}_{\mathcal{O}}(\mathbb{C}^{n})$ satisfies the conditions of Theorem 1.5.
Then $\Omega$ is biholomorphic to a Runge and Stein domain in $\mathbb{C}^{n}$.
Since we have the approximation on certain spirallike domains satisfying conditions Theorem 1.5, Corollary 1.9, hence, the proof of the following theorems follow exactly the same way as [ABFW13, Theorem 3.4]. Therefore, we omit the proofs here.
We now give proof of Theorem 1.12.
Proof of Theorem 1.12.
Using Result 2.4 we get that there exists a Loewner chain $g_{t}:\Omega\to N$ of order $d$ that solves the Loewner PDE
$\frac{\partial f_{t}}{\partial t}(z)$ $=-df_{t}(G(z,t)),~{}~{}a.e.\quad t\geq 0,~{}~{}\forall z\in\Omega,$ (5.1)
where $N=\cup_{t\geq 0}g_{t}(\Omega)$ is a complex manifold of dimension $n$. We also obtain that any other solution to equation (5.1) with values in another $n$ dimensional complex manifold $Q$ is of the form $\Lambda\circ g_{t}$, where $\Lambda:N\to Q$ is holomorphic. By using Result 2.6, $\left(g_{j}(\Omega),g_{j+1}(\Omega)\right)$ is a Runge pair for every $j\in\mathbb{N}$. Since $g_{t}$ is a univalent holomorphic map, $g_{t}(\Omega)$ is an open connected subset of $N$. Clearly, $g_{t}(\Omega)$ is biholomorphic to $\Omega$. By assumption, $\Omega$ is spirallike with respect to $V$ such that $V$ satisfies the (1.3). Therefore, $g_{j}(\Omega)$ satisfies the Condition $(ii)$ of Theorem 5.1 for all $j\in\mathbb{N}$. Hence, thanks to Theorem 5.1, we get that $N$ is biholomorphic to some Runge Stein domain $G\subset\mathbb{C}^{n}$. Hence, it follows that there exists a biholomorphic map $F:N\to G$. Let $f_{t}=F\circ g_{t}$. We have the following
$\frac{\partial f_{t}}{\partial t}(z)$ $=DF(g_{t}(z))\frac{\partial g_{t}}{\partial t}(z)$
$=DF(g_{t}(z))(-dg_{t}(z)(G(z,t)))$
$=-D(F(g_{t}(z)))G((z,t))$
$=-D(f_{t}(z))(G(z,t)).$
This implies that $f_{t}$ satisfies (5.1). If $\tilde{g_{t}}(z)$ is another solution of the (5.1) with values in $\mathbb{C}^{n}$, then we deduce, from Theorem 2.4, that there is a biholomorphism $\Lambda:N\to\cup_{t\geq 0}\tilde{g_{t}}(\Omega)$ such that $\tilde{g_{t}}=\Lambda\circ g_{t}=\Lambda\circ F^{-1}(f_{t})$. Consider $\varphi=\Lambda\circ F^{-1}$. Therefore, any solution of (5.1) with values in $\mathbb{C}^{n}$ is of the form $\varphi(f_{t})$ for some suitable biholomorphism $\varphi\colon G\to\mathbb{C}^{n}$. ∎
6. Examples of spirallike domains
In this section, we provide a couple of examples of spirallike domains where every biholomorphism with Runge image can be approximated by elements Aut$(\mathbb{C}^{n})$.
Example 6.1.
Let $A=\lambda I_{n}+\mathcal{N}$, where $\mathcal{N}$ is a $n\times n$ matrix with ones in the first diagonal above the main diagonal and zeros elsewhere and $\lambda=\lambda_{1}+i\mu_{1}$, with $\lambda_{1}<0$. Let $\Omega:=\{(z_{1},z_{2},,\ldots,z_{n}):\bigg{|}z_{n-1}-\frac{z_{n}}{\lambda_{1}}\ln{|z_{n}|}\bigg{|}<1\}$. Then every biholomorphism from $\Omega$ onto a Runge domain can be approximated by elements of $\text{Aut}(\mathbb{C}^{n})$.
Explanation. Suppose that $A=-\lambda I_{n}+\mathcal{N}$. Let $w=e^{At}z$. We get that
$w=e^{At}=e^{\lambda_{1}t}\begin{bmatrix}1&t&\frac{t^{2}}{2}&\cdots&&\frac{t^{n}}{n!}\\ 0&1&t&\cdots&&\frac{t^{n-1}}{(n-1)!}\\ \vdots&\vdots&\vdots&\ddots&&\vdots\\ 0&0&0&\cdots&&t\\ 0&0&0&\cdots&&1\\ \end{bmatrix}.$
For all $z\in\Omega$ and for all $t\geq 0$, we obtain that
$\bigg{|}w_{n-1}-\frac{w_{n}}{\lambda_{1}}\ln{|w_{n}|}\bigg{|}$ $=\bigg{|}z_{n-1}e^{\lambda t}+tz_{n}e^{\lambda t}-\frac{z_{n}e^{\lambda t}}{\lambda_{1}}(\ln{|z_{n}|}+t\lambda_{1})\bigg{|}$
$=\bigg{|}z_{n-1}e^{\lambda t}-e^{\lambda t}\frac{z_{n}}{\lambda_{1}}\ln{|z_{n}|}\bigg{|}$
$=e^{\lambda_{1}t}\bigg{|}z_{n-1}-\frac{z_{n}}{\lambda_{1}}\ln{|z_{n}|}\bigg{|}$
$<1.$
Hence, $\Omega$ is a spirallike domain with respect to $A$. Clearly, the matrix $A$ satisfies $2k_{+}(A)=2\lambda_{1}<\lambda_{1}$. Therefore, from Corollary 1.9 it follows that any biholomorphism of $\Omega$, whose image is a Runge domain in $\mathbb{C}^{n}$, can be approximated by elements of $\text{Aut}(\mathbb{C}^{n})$.
Example 6.2.
Let $A=\text{Diag}(-\lambda_{1},-\lambda_{2},\cdots,-\lambda_{n})$, such that $\lambda_{i}\in\mathbb{N}$, with, $\lambda_{1}\leq\lambda_{2}\cdots\leq\lambda_{n}$ and $2\lambda_{1}>\lambda_{2}$. Let $i,j\in\{1,2,\cdots,n\}$ and $m_{i},m_{j}\in\mathbb{N}$, such that $m_{i}\lambda_{i}=m_{j}\lambda_{j}$. Let $\Omega:=\{z\in\mathbb{C}^{n}:\big{|}z_{i}^{m_{i}}-z_{j}^{m_{j}}\big{|}<1\}$. Then every biholomorphism from $\Omega$ onto a Runge domain can be approximated by elements of $\text{Aut}(\mathbb{C}^{n})$.
Explanation. Let $z\in\Omega$ and $w=e^{tA}z$. Then we obtain that
$|w_{i}-w_{j}|=|e^{m_{i}\lambda_{i}t}z_{i}-e^{m_{j}\lambda_{i}t}z_{j}|=e^{m_{i}\lambda_{i}}|z_{i}-z_{j}|<1.$
Hence, $\Omega$ is spirallike with respect to $A$. Since $\lambda_{1}>\lambda_{2}$, hence, it satisfies the condition of the Corollary 1.9. Therefore, the conclusion is true.
Example 6.3.
Let $D\subset\mathbb{C}^{n}$ be a spirallike domain with respect to $V\in\mathfrak{X}_{\mathcal{O}}(\mathbb{C}^{n})$. Assume that $V$ satisfies the condition of Theorem 1.5. Let $\Omega=\{(z,w)\in\mathbb{C}^{n+1}:z\in D,~{}\|w\|<1\}\subset\mathbb{C}^{n+1}$. Then $\Omega$ is spirallike with respect to vector field $\widetilde{V}=(V(z),-\beta w)$, where $\beta>0$ is chosen suitably.
Explanation. The flow of the vector field $\tilde{V}$ is $\tilde{X}(t,(z,w))=(X(t,z),e^{-\beta t}w)$, where $X(t,z)$ denotes the flow of the vector field $V$. Clearly, $\Omega$ is spirallike domain with respect to vector field $\widetilde{V}=(V(z),-\beta w)$, where $\beta>0$ will be chosen later. Let $DV(0)=A$. Since the flow of the vector field $V$ satisfies the condition of the Theorem 1.5, hence, for all $z=(z_{1},z_{2})\in D$, there exist $\gamma,\alpha>0$, such that
$\|X(t,z)\|$ $\leq\gamma e^{-\alpha t}\|z\|,\;\;\forall z\in D\;\;\forall t>0,$ (6.1)
and $\alpha>\frac{-k_{-}(A)}{2}$. Choose $\beta>0$ such that
$\min\{\alpha,-k_{-}(A)\}>\beta>\frac{-k_{-}(A)}{2}.$
For $(z,w)\in\Omega$ and $t>0$, we have the following estimate
$\|\tilde{X}(t,(z,w))\|^{2}$ $=\|X(t,z)\|^{2}+e^{-2\beta t}\|w\|^{2}.$
Hence, from (6.1) we get that
$\|\tilde{X}(t,(z,w))\|^{2}$ $\leq\tilde{\gamma}^{2}e^{-2\beta t}(\|z\|^{2}+\|w\|^{2}),$
where $\tilde{\gamma}^{2}:=\max\{1,\gamma^{2}\}.$ Therefore, for every $(z,w)\neq(0,0)$, we obtain that
$\|\tilde{X}(t,(z,w))\|$ $\leq\tilde{\gamma}e^{-\beta t}\|(z,w)\|.$
Clearly, from the choice of $\beta>0$ we have $k_{-}(D\tilde{V}(0))=\min\{k_{-}(A),-\beta\}=k_{-}(A)$. Again from the choice of $\beta>0$, we have $02\beta+k_{-}(D\widetilde{V}(0))>$ 0. Therefore, it satisfies Condition (1.3) of Theorem 1.5.
Next, we provide an example of a complete hyperbolic Hartogs domain which is spirallike with respect to a nonlinear vector field but not spirallike with respect to any invertible matrix. More precisely:
Theorem 6.4.
There is a complete hyperbolic domain $\Omega\subset\mathbb{C}^{2}$ containing the origin which is spirallike with respect to a nonlinear vector field satisfying the conditions of Theorem 1.5, but not spirallike with respect to any invertible matrix.
Let us consider the following domain.
$\Omega=\{(z_{1},z_{2})\in\mathbb{C}^{2}\;:\;{\sf Re}(z_{1})<3,~{}|z_{2}|<e^{-\frac{1}{2}{\sf Re}(z_{1})})\}.$
We will show that $\Omega$ is a candidate domain for Theorem 6.4. We will do this by proving a series of lemmas.
Lemma 6.5.
Suppose $\Omega$ is as above. Then $\Omega$ is a complete hyperbolic domain.
Proof.
We use Result 2.5 to prove this lemma. Since upper half space $H=\{z\in\mathbb{C}:\Im(z)>0\}$ is biholomorphic to a domain $G=\{z\in\mathbb{C}:{\sf Re}(z)<3\}\subset\mathbb{C}$. Hence, $G$ is a complete hyperbolic domain. We choose sequence $c_{j}=\frac{1}{2}$ for all $j\in\mathbb{N}$ and $h_{j}(z)=e^{z}$ for $j\in\mathbb{N}$ then $\{c_{j}\log|{h_{j}}(z)|\}$ converges to $\frac{1}{2}{\sf Re}(z)$. Therefore, using Result 2.5, we conclude that $\Omega$ is complete hyperbolic. ∎
Lemma 6.6.
$\Omega$ is spirallike with respect to the vector field $F(z_{1},z_{2})=(-2z_{1},-3z_{2}+z_{1}z_{2})$.
Proof.
The flow of $F$ is
$X(t,z)=\left(z_{1}e^{-2t},z_{2}e^{-3t}e^{\frac{z_{1}}{2}(1-e^{-2t})}\right),\;\;\forall t\in\mathbb{R}$
Let $(z_{1},z_{2})\in\Omega$ and $t\geq 0$. Let $w_{1}=z_{1}e^{-2t}$ and $w_{2}=z_{2}e^{-3t}e^{\frac{z_{1}}{2}(1-e^{-2t})}$. Then we have:
$|w_{2}|$ $=|z_{2}|e^{-3t}||e^{\frac{z_{1}}{2}(1-e^{-2t})}|$
$\leq|z_{2}||e^{\frac{z_{1}}{2}(1-e^{-2t})}|.$
Since $(z_{1},z_{2})\in\Omega$, we get that
$|w_{2}|$ $\leq e^{-\frac{1}{2}{\sf Re}(z_{1})}e^{\frac{1}{2}{\sf Re}(z_{1})(1-e^{-2t})}$
$\leq e^{-\frac{1}{2}{\sf Re}(z_{1}e^{-2t})}$
$<e^{-\frac{1}{2}{\sf Re}(w_{1})}.$
It is clear that ${\sf Re}(w_{1})<3$. Therefore, we get $(w_{1},w_{2})\in\Omega$ for all $t\geq 0$. Hence, $\Omega$ is spirallike with respect to $F$. ∎
We now show that the domain $\Omega$ satisfies our condition for approximation.
Lemma 6.7.
Any biholomorphism from $\Omega$ into $\mathbb{C}^{2}$ can be approximated uniformly on compact subsets of $\Omega$ by automorphisms of $\mathbb{C}^{2}$.
Proof.
We show that the flow of the vector field $F$ satisfies Condition (1.3). Let $\zeta=(\zeta_{1},\zeta_{2})\in\Omega$. We have the estimate of the flow on the domain $\Omega$
$\|X(t,\zeta)\|^{2}$ $=e^{-4t}\|\zeta_{1}\|^{2}+\|\zeta_{2}\|^{2}e^{-6t}e^{{\sf Re}(\zeta_{1})(1-e^{-2t})}.$
Since ${\sf Re}(\zeta_{1})<3$, there exist $M>0$ such that for all $t>M$ we have $-2t+(1-e^{-2t}){\sf Re}{\zeta_{1}}<1$. Let $s=\sup_{t\in[0,M]}-2t+3(1-e^{-2t})$ Therefore, we have
$\|X(t,\zeta)\|^{2}$ $\leq\gamma^{2}e^{-4t}\big{(}\|\zeta_{1}\|^{2}+\|\zeta_{2}\|^{2}\big{)}$
$\|X(t,\zeta)\|$ $\leq\gamma e^{-2t}\|\zeta\|,$
where $\gamma^{2}:=\max\{1,s\}$.
Here $\alpha=2$ and $k_{-}(DF(0))=-3$. Hence, $02\alpha+k_{-}(DF(0))>$ 0. Hence, $\Omega$ satisfies (1.3). Therefore. in view of Theorem 1.1, any biholomorphism of $\Omega$ can be approximated uniformly on compacts by $Aut(\mathbb{C}^{2})$. ∎
Next, we show that $\Omega$ is not spirallike with respect to any $A\in Gl_{2}(\mathbb{C})$. The way we follow is to construct a point lying in $\Omega$ and a positive time when the flow of a linear vector field leaves the domain. We construct it in different cases which are given in a series of lemmas that follows.
Lemma 6.8.
The domain $\Omega$ is not spirallike with respect to any diagonal matrix.
Proof.
Let $A=\begin{bmatrix}d_{1}&0\\ 0&d_{2}\\ \end{bmatrix}$ be a matrix with $d_{1}=a_{1}+ib_{1},\ d_{2}=a_{2}+ib_{2}$
Case 1.
$b_{1}\neq 0$.
Let $(z_{1},z_{2})=(-2p,1)$, where $p>0$ is chosen such that
$\frac{a_{2}\pi}{|b_{1}|}+pe^{\frac{a_{1}\pi}{|b_{1}|}}>0.$ (6.2)
Clearly, $(-2p,1)\in\Omega$. The flow of the linear vector field $A$ with initial point $(z_{1},z_{2})$ is
$\begin{bmatrix}w_{1}\\ w_{2}\end{bmatrix}=\begin{bmatrix}e^{d_{1}t}z_{1}\\ e^{d_{2}t}z_{2}\\ \end{bmatrix}.$
Choose $t_{1}=\frac{\pi}{|b_{1}|}$. Hence, at time $t_{1}$ we have $|w_{2}|=e^{\frac{a_{2}\pi}{|b_{1}|}}|z_{2}|=e^{\frac{a_{2}\pi}{|b_{1}|}}$ and $e^{-\frac{1}{2}{\sf Re}(w_{1})}=e^{-\frac{1}{2}(2p\cdot e^{\frac{a_{1}\pi}{|b_{1}|}})}=e^{-pe^{\frac{a_{2}\pi}{|b_{1}|}}}$. From (6.2) it follows that
$e^{\frac{a_{2}\pi}{|b_{1}|}+pe^{\frac{a_{1}\pi}{|b_{1}|}}}>1\Longrightarrow e^{\frac{a_{2}\pi}{|b_{1}|}}>e^{-pe^{\frac{a_{1}\pi}{|b_{1}|}}}\Longrightarrow|w_{2}|>e^{-\frac{1}{2}({\sf Re}(w_{1}))}.$
This implies that $(w_{1},w_{2})\notin\Omega$.
Case 2.
Suppose $b_{1}=0$. Since $A$ is invertible matrix, $a_{1}\neq 0.$
Subcase 1.
Assume that $a_{1}<0$. Hence, $d_{1}=a_{1}<0$. Choose $t_{1}=1$ and $z_{1}=-2p$, where $p=\frac{2|d_{2}|}{1-e^{a_{1}}}$. Hence, $p(1-e^{a_{1}})-|a_{2}|=2|d_{2}|-|a_{2}|>0\Longrightarrow p>pe^{a_{1}}+|a_{2}|.$ Choose $z_{2}\in\mathbb{R}$ such that
$e^{p}>z_{2}$ $>e^{pe^{a_{1}}+|a_{2}|}.$ (6.3)
The flow of the linear vector field $A$ with initial point $(z_{1},z_{2})$ is
$\begin{bmatrix}w_{1}\\ w_{2}\end{bmatrix}=e^{Dt}\begin{bmatrix}z_{1}\\ z_{2}\end{bmatrix}=\begin{bmatrix}e^{d_{1}t}z_{1}\\ e^{d_{2}t}z_{2}\\ \end{bmatrix}=\begin{bmatrix}e^{a_{1}t}z_{1}\\ e^{a_{2}t}(\cos{b_{2}t}+i\sin{b_{2}t})z_{2}\\ \end{bmatrix}.$
From (6.3) it follows that $z_{2}<e^{p}=e^{-\frac{1}{2}(-2p)}=e^{-\frac{1}{2}{\sf Re}(z_{1})}$, i.e., $(z_{1},z_{2})\in\Omega$. At $t=1$ we get that $|w_{2}|=e^{a_{2}}|z_{2}|$ and $e^{{\frac{-1}{2}{\sf Re}(w_{1})}=e^{-\frac{1}{2}e^{a_{1}}\cdot(-2p)}}=e^{pe^{a_{1}}}.$ Hence, from (6.3), we get that for any $a_{2}\in\mathbb{R}$, $|w_{2}|>e^{-\frac{1}{2}{\sf Re}(w_{1})}$. Therefore, $(w_{1},w_{2})\notin\Omega$.
Subcase 2.
Assume $a_{1}>0$. Note that $(2,\frac{1}{e^{2}})\in\Omega$. Choose $t=t_{1}^{\prime}>0$ such that $e^{a_{1}t_{1}^{\prime}}+a_{2}t_{1}^{\prime}-2>0$. At time $t^{\prime}_{1}>0$ with the initial point $(2,\frac{1}{e^{2}})$ we have $|w^{\prime}_{2}|=e^{a_{2}t^{\prime}_{1}}\frac{1}{e^{2}}=e^{a_{2}t^{\prime}_{1}-2},$ and $e^{\frac{-1}{2}{\sf Re}(w^{\prime}_{1})}=e^{-e^{a_{1}t_{1}^{\prime}}}.$ Since $t_{1}^{\prime}>0$, we obtain that
$e^{e^{a_{1}t_{1}^{\prime}}+a_{2}t_{1}^{\prime}-2}>1\Longrightarrow|w^{\prime}_{2}|>e^{-\frac{1}{2}{\sf Re}(w^{\prime}_{1})},$
which implies that $(w^{\prime}_{1},w^{\prime}_{2})\notin\Omega$. Therefore, $\Omega$ is not spirallike with respect to any diagonal matrix.
Lemma 6.9.
The domain $\Omega$ is not spirallike with respect to any upper triangular matrix.
Proof.
Suppose that $A=\begin{bmatrix}\lambda&1\\ 0&\lambda\end{bmatrix}$, where $\lambda=\lambda_{1}+i\lambda_{2}$
Case 1. $\lambda_{2}\neq 0$.
The flow of the upper triangular matrix with initial point $(z_{1},z_{2})$ is
$\begin{bmatrix}w_{1}\\ w_{2}\end{bmatrix}$ $=e^{At}\begin{bmatrix}z_{1}\\ z_{2}\end{bmatrix}=\begin{bmatrix}e^{\lambda t}&te^{\lambda t}\\ 0&e^{\lambda t}\end{bmatrix}\begin{bmatrix}z_{1}\\ z_{2}\end{bmatrix}=\begin{bmatrix}e^{\lambda t}(z_{1}+tz_{2})\\ e^{\lambda t}z_{2}\end{bmatrix}.$ (6.14)
Choose $p>0$ such that
$\frac{\lambda_{1}\pi}{|\lambda_{2}|}-\frac{\pi e^{\frac{\pi\lambda_{1}}{|\lambda_{2}|}}}{2|\lambda_{2}|}+pe^{\frac{\lambda_{1}\pi}{|\lambda_{2}|}}$ $>0,$ (6.15)
Since $|z_{2}|=1<e^{-\frac{1}{2}z_{1}}=e^{p}$, $(z_{1},z_{2})\in\Omega$. At $t=\frac{\pi}{|\lambda_{2}|}$ and with initial point $(z_{1},z_{2})=(2p,-1)$ such that (6.15) holds. We have that $|w_{2}|=|e^{\lambda t}||z_{2}|=e^{\frac{\pi\lambda_{1}}{|\lambda_{2}|}}$
$e^{-\frac{1}{2}{\sf Re}(w_{1})}$ $=e^{-\frac{1}{2}{\sf Re}(z_{1}+tz_{2})e^{\lambda t}}=e^{\frac{1}{2}\cdot(\frac{\pi}{|\lambda_{2}|}-2p)e^{\frac{\lambda_{1}\pi}{|\lambda_{2}|}}}.$ (6.16)
From (6.15)) we obtain that
$e^{\frac{\lambda_{1}\pi}{|\lambda_{2}|}}>e^{\frac{1}{2}\cdot(\frac{\pi}{|\lambda_{2}|}-2p)e^{\frac{\lambda_{1}\pi}{|\lambda_{2}|}}}.$ (6.17)
Using (6.16), (6.17) we get that $|w_{2}|>e^{-\frac{1}{2}{\sf Re}(w_{1})}$, which implies that $(w_{1},w_{2})\notin\Omega$ at $t=\frac{\pi}{|\lambda_{2}|}$. Case 2. $\lambda_{2}=0$. In this case $\lambda_{1}\neq 0$. From (6.14) it follows that, in this case, the flow is
$\begin{bmatrix}w_{1}\\ w_{2}\end{bmatrix}=e^{At}\begin{bmatrix}z_{1}\\ z_{2}\end{bmatrix}=\begin{bmatrix}e^{\lambda_{1}t}&te^{\lambda_{1}t}\\ 0&e^{\lambda_{1}t}\end{bmatrix}\begin{bmatrix}z_{1}\\ z_{2}\end{bmatrix}=\begin{bmatrix}e^{\lambda_{1}t}(z_{1}+tz_{2})\\ e^{\lambda_{1}t}z_{2}\end{bmatrix}$
Subcase 1. $\lambda_{1}>0$.
Clearly, $(0,\frac{1}{2})\in\Omega$. Choose $t=t_{1}>0$ such that $e^{\left(\lambda_{1}t_{1}+\frac{t_{1}}{4}e^{\lambda_{1}t_{1}}\right)}>2$. The flow at $t_{1}>0$ with initial point $(0,\frac{1}{2})$ satisfies $|w_{2}|=e^{\lambda_{1}t_{1}}\cdot\frac{1}{2}$ and $e^{-\frac{1}{2}{\sf Re}(w_{1})}=e^{-\frac{t_{1}e^{\lambda_{1}t_{1}}}{4}}.$ Hence, $|w_{2}|>e^{-\frac{1}{2}{\sf Re}(w_{1})}$, which implies that $(w_{1},w_{2})\notin\Omega$.
Subcase 2. $\lambda_{1}<0$.
Let $\lambda_{1}=-k$ for some $k>0$. The flow with initial point $(z_{1},z_{2})$ is
$\begin{bmatrix}w_{1}\\ w_{2}\end{bmatrix}=\begin{bmatrix}e^{-kt}(z_{1}+tz_{2})\\ e^{-kt}z_{2}\end{bmatrix}.$ (6.22)
Choose a fix $t^{\prime}_{1}>0$ such that $12e^{t_{1}(k-e^{kt_{1}})}<$ 1 and $(z_{1},z_{2})=(-2t_{1}e^{kt_{1}},2e^{kt_{1}})$. We obtain that $|z_{2}|\cdot e^{\frac{1}{2}{\sf Re}(z_{1})}=2e^{t_{1}(k-e^{kt_{1}})}<1$, hence, $(z_{1},z_{2})\in\Omega$. Hence, we get that $|w^{\prime}_{2}|=e^{-kt^{\prime}_{1}}\cdot 2e^{kt^{\prime}_{1}}=2$ and $e^{-\frac{1}{2}{\sf Re}(w^{\prime}_{1})}=e^{-\frac{1}{2}e^{-kt^{\prime}}(-2t^{\prime}_{1}e^{kt^{\prime}_{1}}+2t^{\prime}_{1}e^{kt^{\prime}_{1}})}=1.$ Hence, $|w^{\prime}_{2}|>e^{-\frac{1}{2}{\sf Re}(w^{\prime}_{1})}$, which implies that $(w^{\prime}_{1},w^{\prime}_{2})\notin\Omega$. Therefore, $\Omega$ is not spirallike with respect to any upper triangular matrix. ∎
Lemma 6.10.
$\Omega$ is not spirallike with respect to any diagonalizable matrix.
Proof.
Assume that $A$ is diagonalizable matrix. Hence, there exist $P\in Gl_{n}(\mathbb{C})$ such that
$PAP^{-1}=\begin{bmatrix}d_{1}&0\\ 0&d_{2}\\ \end{bmatrix}.$
In this case flow of the linear matrix $A$ with initial point $(z_{1},z_{2})\in\Omega$ is
$e^{At}\begin{bmatrix}z_{1}\\ z_{2}\end{bmatrix}=P^{-1}e^{Dt}P\begin{bmatrix}z_{1}\\ z_{2}\end{bmatrix}.$
If $P=\begin{bmatrix}a&b\\ c&d\\ \end{bmatrix}$ then
$\begin{bmatrix}w_{1}\\ w_{2}\end{bmatrix}$ $=e^{At}\begin{bmatrix}z_{1}\\ z_{2}\end{bmatrix}=\frac{1}{ad-bc}\begin{bmatrix}e^{d_{1}t}(adz_{1}+bdz_{2})+e^{d_{2}t}(-bcz_{1}-bdz_{2})\\ e^{d_{1}t}(-acz_{1}-bcz_{2})+e^{d_{2}t}(acz_{1}+adz_{2})\\ \end{bmatrix}.$ (6.29)
[…]
$\begin{bmatrix}w_{1}\\ w_{2}\end{bmatrix}=e^{At}\begin{bmatrix}z_{1}\\ z_{2}\end{bmatrix}=\begin{bmatrix}e^{dt}z_{1}\\ e^{dt}z_{2}\end{bmatrix},$
which is same as the flow of the diagonal matrix. Hence, from Lemma 6.8 we get that there exist $(z_{1},z_{2})\in\Omega$ and $t>0$, such that $(w_{1},w_{2})\notin\Omega$. Assume that $d_{1}\neq d_{2}$ and construct the initial and exit time of the flow in the following way.
Case 1. $ac\neq 0$. Subcase 1. $(e^{d_{1}}ad-e^{d_{2}}bc)\neq 0$. We choose the initial point
$(z_{1},z_{2})$ =$=\begin{cases}\left(\frac{3\mu\overline{(e^{d_{1}}ad-e^{d_{2}}bc)}}{|ac||e^{d_{1}}-e^{d_{2}}|(|e^{d_{1}}ad-e^{d_{2}}bc|)}i,0\right),~{}\text{if}~{}{\sf Re}(\mu\overline{(e^{d_{1}}ad-e^{d_{2}}bc)}\cdot i)<0\\ \left(\frac{-3\mu\overline{(e^{d_{1}}ad-e^{d_{2}}bc)}}{|ac||e^{d_{1}}-e^{d_{2}}|(|e^{d_{1}}ad-e^{d_{2}}bc|)}i,0\right),~{}\text{if}~{}{\sf Re}(\mu\overline{(e^{d_{1}}ad-e^{d_{2}}bc)}\cdot i)>0,\\ \end{cases}$ (6.30)
where $\mu=(ad-bc)$. Since $(z,0)\in\Omega$ for any $z$ with ${\sf Re}(z)<3$, we get that $(z_{1},z_{2})\in\Omega$. If the integral curves reaches at $(w_{1},w_{2})$, at $t=1$, then we get, from (6.29), that
$\begin{bmatrix}w_{1}\\ w_{2}\\ \end{bmatrix}$ $=\frac{1}{\mu}\begin{bmatrix}(e^{d_{1}}ad-e^{d_{2}}bc)z_{1}\\ (e^{d_{2}}-e^{d_{1}})acz_{1}\\ \end{bmatrix}.$
Hence, $|w_{2}|=\frac{1}{|\mu|}|ac||e^{d_{1}}-e^{d_{2}}||z_{1}|=3$. If $(z_{1},z_{2})=\left(\frac{3\mu\overline{(e^{d_{1}}ad-e^{d_{2}}bc)}}{|ac||e^{d_{1}}-e^{d_{2}}|(|e^{d_{1}}ad-e^{d_{2}}bc|)}i,0\right)$ then ${\sf Re}(w_{1})={\sf Re}\left(\frac{1}{\mu}(e^{d_{1}}ad-e^{d_{2}}bc)\cdot z_{1}\right)=0$. Also if
$(z_{1},z_{2})=\left(\frac{-3\mu\overline{(e^{d_{1}}ad-e^{d_{2}}bc)}}{|ac||e^{d_{1}}-e^{d_{2}}|(|e^{d_{1}}ad-e^{d_{2}}bc|)}i,0\right)$
then, by similar way, we get that ${\sf Re}(w_{1})=0$. Hence, $3=|w_{2}|>1=e^{-\frac{1}{2}{\sf Re}(w_{1})}$. Therefore, $(w_{1},w_{2})\notin\Omega$. Subcase 2. $e^{d_{1}}ad=e^{d_{2}}bc$. We choose $(z_{1},z_{2})=\left(\frac{-3|\mu|}{|ac||(e^{d_{2}}-e^{d_{1}})|},0\right)\in\Omega.$ If the integral curve starting at $(z^{\prime}_{1},z^{\prime}_{2})$ reaches at $(w^{\prime}_{1},w^{\prime}_{2})$ at time $t=1$ then, from (6.29), we get that
$\begin{bmatrix}w_{1}\\ w_{2}\\ \end{bmatrix}$ $=\frac{1}{\mu}\begin{bmatrix}0\\ acz_{1}(e^{d_{2}}-e^{d_{1}})\\ \end{bmatrix}$
Hence, $e^{-\frac{1}{2}{\sf Re}(w_{1})}=e^{0}=1$ and $|w_{2}|=|ac||z_{1}||(e^{d_{2}}-e^{d_{1}})|=\frac{|ac|}{|\mu|}\cdot\frac{3|\mu|}{|ac||(e^{d_{2}}-e^{d_{1}})|}|(e^{d_{2}}-e^{d_{1}})|=3.$ Therefore, $w_{2}>e^{-\frac{1}{2}{\sf Re}(w_{1})}$, which implies that $(w_{1},w_{2})\notin\Omega$.
Case 2. $ac=0$. Since $ad-bc\neq 0$, hence both $a$ and $c$ can not be zero. Either $a=0$ or $c=0$. we construct the initial point in $\Omega$ and exit time $t>0$, assuming $a=0,c\neq 0.$ Construction of exit time and the initial point is similar for the case $a\neq 0,~{}~{}c=0.$ If $d=0$, then it will be the case as in Lemma 6.8. We assume that $d\neq 0$. From (6.29), we get that the flow of the liner vector field $A$ with initial point $(z_{1},z_{2})$ is
$\begin{bmatrix}w_{1}\\ w_{2}\\ \end{bmatrix}$ $=e^{At}\begin{bmatrix}z_{1}\\ z_{2}\\ \end{bmatrix}=\begin{bmatrix}z_{1}e^{d_{2}t}+\frac{d}{c}\cdot z_{2}(e^{d_{2}t}-e^{d_{1}t})\\ e^{d_{1}t}z_{2}\end{bmatrix},$ (6.37)
where $d_{1}=a_{1}+ib_{1}$ and $d_{2}=a_{2}+ib_{2}$. Subcase 1. $a_{1}>0$.
Choose $t=t_{1}>0$ such that $\dfrac{e^{a_{1}t_{1}}}{2}>1$ and the initial point
$(z_{1},z_{2})=\left(0,\frac{c}{d}\cdot\frac{|d|\overline{(e^{d_{2}t_{1}}-e^{d_{1}t_{1}})}}{2|c||(e^{d_{2}t_{1}}-e^{d_{1}t_{1}})|}\cdot i\right).$
Since $|z_{2}|=\frac{1}{2}<e^{-\frac{1}{2}\cdot 0}=1$, $(z_{1},z_{2})\in\Omega$. Using (6.37) we get, at the time $t=t_{1}$, that $|w_{2}|=e^{a_{1}t_{1}}|z_{2}|=e^{a_{1}t_{1}}\frac{1}{2}>1$, and
$w_{1}$ $=z_{1}e^{d_{2}t_{1}}+\frac{d}{c}\cdot z_{2}(e^{d_{2}t_{1}}-e^{d_{1}t_{1}})=\frac{|d||(e^{d_{2}t_{1}}-e^{d_{1}t_{1}})|}{2|c|}i$
$\Longrightarrow$+⟹ + $e^{-\frac{1}{2}{\sf Re}(w_{1})}=1.$ (6.38)
Hence, we obtain that $w_{2}>e^{-\frac{1}{2}{\sf Re}(w_{1})}$. Therefore, $(w_{1},w_{2})\notin\Omega$.
[…]
$\begin{bmatrix}w_{1}\\ w_{2}\\ \end{bmatrix}$ $=\begin{bmatrix}z_{1}e^{a_{2}t}+\frac{d}{c}\cdot z_{2}(e^{a_{2}t}-e^{ib_{1}t})\\ e^{ib_{1}t}z_{2}\end{bmatrix}.$ (6.43)
Choose $t_{1}=\frac{(2m+1)\pi}{|b_{1}|}>0$, where $m\in\mathbb{N}$ will be chosen later. From (6.43) we get that $|w_{2}|=|e^{ib_{1}t_{1}}||z_{2}|=\frac{|d|}{|c|}.$ We have $(p-\frac{|d|^{2}}{2|c|^{2}})<0$. We again compute
$w^{\prime}_{1}$ $=z^{\prime}_{1}e^{a_{2}t^{\prime}_{1}}+\frac{d}{c}\cdot z^{\prime}_{2}(e^{a_{2}t^{\prime}_{1}}-e^{ib_{1}t^{\prime}_{1}})$
$=e^{a_{2}t^{\prime}_{1}}\left(\frac{|d|^{2}}{|c|^{2}}-2p\right)+\frac{|d|^{2}}{|c|^{2}}$
$-\frac{1}{2}{\sf Re}(w^{\prime}_{1})$ $=\left(p-\frac{|d|^{2}}{2|c|^{2}}\right)e^{a_{2}t^{\prime}_{1}}-\frac{|d|^{2}}{2|c|^{2}}.$ (6.44)
Since $a_{2}>0$, hence from (6), it follows that $e^{-\frac{1}{2}{\sf Re}(w^{\prime}_{1})}\to 0$, as $m\to\infty$. Hence, letting $m\to\infty$, we get that there exist $m_{1}\in\mathbb{N}$, such that we have $|w^{\prime}_{2}|>e^{-\frac{1}{2}{\sf Re}(w^{\prime}_{1})}$. Hence, $(w^{\prime}_{1},w^{\prime}_{2})\notin\Omega$.
Subcase 3. $a_{1}=0,~{}~{}b_{2}=0,~{}~{}a_{2}<0.$ In this subcase, we construct the initial point and the exist time in the following way. Let $a_{2}=-k_{2}$, where $k_{2}>0$. Choose $(\eta_{1},\eta_{2})=(-2p,\frac{\overline{d}}{\overline{c}})$, where $p>0$ is chosen in such a way that
$|\eta_{2}|=\frac{|d|}{|c|}<e^{-\frac{1}{2}{\sf Re}(\eta_{1})}=e^{p},$ (6.45)
which implies that $(\eta_{1},\eta_{2})\in\Omega$. Suppose that the integral curve of the vector field with initial point $(\eta_{1},\eta_{2})$ reaches at $(\zeta_{1},\zeta_{2})$ at time $t=T=\frac{(2m+1)\pi}{|b_{1}|}>0$. From (6.43), we get that $|\zeta_{2}|=|e^{ib_{1}T}||\eta_{2}|=\frac{|d|}{|c|}.$ By similar calculation as (6) we get that $-\frac{1}{2}{\sf Re}(\zeta_{1})=(p-\frac{|d|^{2}}{2|c|^{2}})e^{a_{2}T}-\frac{|d|^{2}}{2|c|^{2}}$ and choose $p=K^{\prime}\frac{|d|^{2}}{|c|^{2}}+\frac{|d|^{2}}{2|c|^{2}}$, where $K^{\prime}>0$ is chosen large enough such that 6.45 holds and we fix it. Then we get that
$-\frac{1}{2}{\sf Re}(\zeta_{1})$ $=K^{\prime}\frac{|d|^{2}}{|c|^{2}}e^{a_{2}T}-\frac{|d|^{2}}{2|c|^{2}}.$ (6.46)
Suppose that $\left(\frac{1}{2}-\frac{|c|^{2}}{|d|^{2}}\ln{\frac{|c|}{|d|}}\right)\neq 0$. Now choose $m\in\mathbb{N}$ such that
$e^{k_{2}T}$ $>\frac{K^{\prime}}{\left(\frac{1}{2}-\frac{|c|^{2}}{|d|^{2}}\ln{\frac{|c|}{|d|}}\right)}.$ (6.47)
Hence, we obtain that
$\left(\frac{1}{2}-\frac{|c|^{2}}{|d|^{2}}\ln{\frac{|c|}{|d|}}\right)>K^{\prime}e^{-k_{2}T}$
$\left(\frac{1}{2}-\frac{|c|^{2}}{|d|^{2}}\ln{\frac{|c|}{|d|}}\right)>K^{\prime}e^{a_{2}T}$
$e^{\left(\frac{|d|^{2}}{2|c|^{2}}-\frac{|d|^{2}K^{\prime}}{|c|^{2}}\cdot e^{a_{2}T}\right)}>\frac{|c|}{|d|}$
$\frac{|d|}{|c|}>e^{K^{\prime}\cdot\frac{|d|^{2}}{|c|^{2}}\cdot e^{a_{2}T}}\cdot e^{\frac{-|d|^{2}}{2|c|^{2}}}.$ (6.48)
Using (6.46) we get that $|\zeta_{2}|>e^{-\frac{1}{2}{\sf Re}(\zeta_{1})}.$ This implies that $(\zeta_{1},\zeta_{2})\notin\Omega$.
If $\frac{|c|^{2}}{|d|^{2}}\ln{\frac{|c|}{|d|}}=\dfrac{1}{2}$, then $\dfrac{|c|}{|d|}>1$. We construct the initial point and the exit time in the following way. Let $\alpha=\frac{|d|}{|c|}<1$. Choose $(\eta_{1},\eta_{2})=(-2p,\frac{\overline{d}}{\overline{c}}),$ where $p=\ln(\alpha)+e^{-k_{2}t}$ and $t=\frac{(2m+1)\pi}{|b_{1}|}>0$ with large enough $m\in\mathbb{N}$. Now for every $m\in\mathbb{N}$ we have
$1<e^{e^{-k_{2}t}}\Longrightarrow\alpha<e^{\ln(\alpha)}e^{e^{-k_{2}t}}\Longrightarrow|\eta_{2}|<e^{p}=e^{\frac{-1}{2}{\sf Re}(\eta_{1})}.$
Since $m\to\infty$ and $p\to\ln(\alpha)$ and $-2\ln{\alpha}<3$, hence, we get that for large enough $m\in\mathbb{N}$, ${\sf Re}(\eta)<3$, so $(\eta_{1},\eta_{2})\in\Omega$. We also have
$\frac{\alpha^{2}}{2}+\ln(\alpha)$ $=0.$ (6.49)
Suppose that the integral curve of the linear vector field reaches at $(\zeta_{1},\zeta_{2})$, at time $t=T=\frac{(2m+1)\pi}{|b_{1}|}>0$, where $m\in\mathbb{N}$ is large enough. Then, from (6.43), we get that $|\zeta_{2}|=|e^{ib_{1}T}||\eta_{2}|=\frac{|d|}{|c|}$ and
$-\frac{1}{2}{\sf Re}(\zeta_{1})$ $=(p-\frac{|d|^{2}}{2|c|^{2}})e^{a_{2}T}-\frac{|d|^{2}}{2|c|^{2}}$
$=\left(\ln(\alpha)+e^{-k_{2}T}-\frac{\alpha^{2}}{2}\right)e^{-k_{2}T}-\frac{\alpha^{2}}{2}.$ (6.50)
Since $\alpha<1$, we get that $\ln{\frac{1}{\alpha^{2}}}>0$. Choose $m\in\mathbb{N}$ such that $T>\frac{1}{k_{2}}\ln{\frac{1}{\alpha^{2}}},$. Then we obtain that
$0>e^{-k_{2}T}-\alpha^{2}$
$\Longrightarrow$ $0>e^{-k_{2}T}(e^{-k_{2}T}-\alpha^{2})$
$\Longrightarrow$ $\ln{\alpha}+\frac{\alpha^{2}}{2}>e^{-k_{2}T^{\prime}}(e^{-k_{2}T^{\prime}}-\alpha^{2}).$
From (6.49), we get that
$\Longrightarrow$ $\ln{\alpha}+\frac{\alpha^{2}}{2}>e^{-k_{2}T^{\prime}}(e^{-k_{2}T^{\prime}}-\frac{\alpha^{2}}{2}+\ln{\alpha})$
$\Longrightarrow$ $\alpha=e^{e^{-k_{2}T^{\prime}}\left(e^{-k_{2}T^{\prime}}-\frac{\alpha^{2}}{2}+\ln{\alpha}\right)-\frac{\alpha^{2}}{2}}.$
Using (6), we get that $|\zeta^{\prime}_{2}|>e^{\frac{-1}{2}{\sf Re}(\zeta^{\prime}_{1})},$ which implies that $(\zeta^{\prime}_{1},\zeta^{\prime}_{2})\notin\Omega$.
Subcase 4. $a_{1}=0,~{}~{}b_{2}\neq 0.$ Suppose that the integral curve of the matrix reaches at $(w_{1},w_{2})$ at $t>0$. From (6.37), we get that
$\begin{bmatrix}w_{1}\\ w_{2}\end{bmatrix}$ $=\begin{bmatrix}z_{1}e^{a_{2}t}(\cos{b_{2}t}+i\sin{b_{2}t})+\frac{d}{c}z_{2}\bigg{(}e^{a_{2}t}(\cos{b_{2}t}+i\sin{b_{2}t})-(\cos{b_{1}t}+i\sin{b_{1}t})\bigg{)}\\ e^{ib_{1}t}z_{2}\end{bmatrix}.$ (6.55)
Choose $(z_{1},z_{2})=(-2p,\frac{\overline{d}}{\overline{c}})$, where $p>0$ will be chosen sufficiently large and $t_{1}=\frac{(2m+1)\pi}{|b_{2}|}$. Clearly, for large $p>0$, $(z_{1},z_{2})\in\Omega$. At $t=t_{1}>0$ we get from (6.55),
$-\frac{1}{2}{\sf Re}(w_{1})$ $=-pe^{a_{2}t_{1}}+\frac{|d|^{2}}{2|c|^{2}}e^{a_{2}t_{1}}+\frac{1}{2}\frac{|d|^{2}}{|c|^{2}}\cos{\frac{b_{1}(2m_{1}+1)\pi}{|b_{2}|}}.$ (6.56)
From (6.56)) it follows that $\lim_{p\to+\infty}(e^{-\frac{1}{2}{\sf Re}(w_{1})})=0.$ Hence, there is a $p>0$ large enough such that $|w_{2}|>e^{-\frac{1}{2}{\sf Re}(w_{1})},$ which implies that $(w_{1},w_{2})\notin\Omega$. Therefore, $\Omega$ is not spirallike with respect to the matrices considered in this subcase.
Subcase 5. $a_{1}<0,~{}~{}b_{2}\neq 0.$ Suppose that $a_{1}=-L$, for some $L>0$. Choose $(z_{1},z_{2})=\left(-2p,\frac{c}{d}\frac{\overline{(e^{d_{2}t_{1}}-e^{d_{1}t_{1}}})}{|(e^{d_{2}t_{1}}-e^{d_{1}t_{1}})|}\cdot\frac{|d|}{|c|}i\right)$, where $t_{1}=\frac{\pi}{|b_{2}|}$ and $p>0$ will be chosen sufficiently large. Suppose that the integral curve reaches at $(w_{1},w_{2})$ at time $t_{1}>0$. We obtain, from (6.37), that $|w_{2}|=|e^{d_{1}t_{1}}||z_{2}|=e^{-Lt_{1}}$ and $w_{1}=-2p(\cos{b_{2}t_{1}}+i\sin{b_{2}t_{1}})e^{a_{2}t_{1}}+\frac{|d|}{|c|}\cdot|(e^{d_{2}t_{1}}-e^{d_{1}t_{1}})|i.$ This imples that $e^{-\frac{1}{2}{\sf Re}(w_{1})}=e^{-(pe^{a_{2}\frac{\pi}{|b_{2}|}})}.$ We choose $p>0$ such that $-Lt_{1}+pe^{a_{2}\frac{|\pi|}{|b_{2}|}}>0.$ Hence, we have $e^{-Lt_{1}}>e^{-\left(pe^{a_{2}\frac{|\pi|}{|b_{2}|}}\right)}$. This implies that $|w_{2}|>e^{\frac{-1}{2}{\sf Re}(w_{1})}.$ Therefore, it follows that $(w_{1},w_{2})\notin\Omega$.
Subcase 6 . $a_{1}<0,~{}~{}~{}~{}b_{2}=0,~{}~{}~{}a_{2}>0.$ Suppose the integral curve reaches at $(w_{1},w_{2})$ at time $t_{1}>0$. From (6.37), we get that
$\begin{bmatrix}w_{1}\\ w_{2}\end{bmatrix}$ $=\begin{bmatrix}z_{1}e^{a_{2}t_{1}}+z_{2}\cdot\frac{d}{c}\left(e^{a_{2}t_{1}}-e^{d_{1}t_{1}}\right)\\ e^{d_{1}t_{1}}z_{2}\end{bmatrix},$ (6.61)
where $d_{1}=a_{1}+ib_{1}$. We construct the initial point and exit time in the following way. Let $(z_{1},z_{2})=(0,\frac{\overline{d}}{M\overline{c}})$, where $M>0$ is so chosen that $|z_{2}|=\frac{|d|}{M|c|}<1=e^{-\frac{1}{2}{\sf Re}(z_{1})}$. Clearly, $(z_{1},z_{2})\in\Omega$. Assuming $b_{1}=0$. Using (6.61), we get that
$w_{1}=z_{1}e^{a_{2}t_{1}}+z_{2}\frac{d}{c}(e^{a_{2}t_{1}}-e^{d_{1}t_{1}})=\frac{|d|^{2}}{|c|^{2}M}(e^{a_{2}t_{1}}-e^{-Lt_{1}}).$
This implies that
$e^{-\frac{1}{2}{\sf Re}(w_{1})}$ $=e^{\left(-\frac{1}{2}\frac{e^{a_{2}t_{1}}|d|^{2}}{|c|^{2}M}+\frac{1}{2}\frac{|d|^{2}}{M|c|^{2}}\cdot e^{-Lt_{1}}\right)},\qquad|w_{2}|=e^{-Lt_{1}}\frac{|d|}{|c|M}$ (6.62)
Since $a_{2}>0$, hence we have
$\lim_{t_{1}\to\infty}Lt_{1}-\frac{1}{2}\frac{|d|^{2}}{|c|^{2}M}e^{a_{2}t_{1}}+\frac{1}{2}\frac{|d|^{2}}{|c|^{2}M}e^{-Lt_{1}}\to-\infty.$
Therefore, there exist a large enough $t_{1}>0$ such that
$e^{-Lt_{1}}\frac{|d|}{M|c|}$ $>e^{\frac{|d^{2}|}{2M|c|^{2}}e^{-Lt_{1}}-\frac{|d|^{2}}{2M|c|^{2}}e^{a_{2}t_{1}}}.$
Hence, it follows from (6.62), that, for large enough $t_{1}>0$, we have $|w_{2}|>e^{-\frac{1}{2}{\sf Re}(w_{1})}.$ If $b_{1}\neq 0$ then choosing $t_{1}=\frac{2m\pi}{|b_{1}|}$ and letting $m\to\infty$, we will have $|w_{2}|>e^{-\frac{1}{2}{\sf Re}(w_{1})}$. Therefore, $(w_{1},w_{2})\notin\Omega$.
Subcase 7. $a_{1}<0,~{}~{}~{}~{}b_{2}=0,~{}~{}~{}a_{2}<0.$
We construct the initial point in $\Omega$ and exit time in the following way. Let $a_{1}=-L,~{}~{}b_{1}=0$ and $a_{2}=-L_{1},~{}~{}b_{2}=0$, with $L>L_{1}>0$. We obtain from (6.37), that the flow of the linear vector field considered here is of the form
$\begin{bmatrix}w_{1}\\ w_{2}\\ \end{bmatrix}$ $=\begin{bmatrix}z_{1}e^{-L_{1}t}+z_{2}\frac{d}{c}(e^{-L_{1}t}-e^{-Lt})\\ e^{-Lt}z_{2}\\ \end{bmatrix}.$ (6.67)
Choose $(z_{1},z_{2})=\left(-2e^{\beta t_{1}},\frac{\overline{d}}{\overline{c}}e^{2Lt_{1}}\right)$, where $t_{1}>0$ will be chosen sufficiently large and $0<\beta<L_{1}$. We choose $M_{1}>0$ such that for $t>M_{1}$ we get $\frac{|d|}{|c|}e^{2Lt^{\prime}-e^{e^{\beta t^{\prime}}}}<1.$ This implies that $|z_{2}|<e^{-\frac{1}{2}{\sf Re}(z_{1})}.$ Suppose that the integral curves reaches at $(w_{1},w_{2})$ at time $t=t_{1}>0$, with initial point $(z_{1},z_{2})$. From (6.67), we obtain that $|w_{2}|=e^{-Lt_{1}}|z_{2}|=\frac{|d|}{|c|}e^{Lt_{1}}$ and $-\frac{1}{2}{\sf Re}(w_{1})=e^{\beta t_{1}}e^{-L_{1}t_{1}}-\frac{|d|^{2}}{2|c|^{2}}e^{Lt_{1}}(e^{(L-L_{1})t_{1}}-1).$ Since $(\beta-L_{1})<0$ and $(L-L_{1})>0$, hence $\lim_{t_{1}\to+\infty}(e^{-\frac{1}{2}{\sf Re}(w_{1})})=0,$ Therefore, there exist $M_{2}>0$ such that $|w_{2}|>e^{-\frac{1}{2}{\sf Re}(w_{1})}$ for all $t_{1}>M_{2}$. Taking $t_{1}>\max\{M_{1},M_{2}\}$, we get that $(w_{1},w_{2})\notin\Omega$. If $b_{1}\neq 0$, then we take $t_{1}=\frac{2m\pi}{|b_{1}|}$ and letting $m\to\infty$ we obtain that $(w^{\prime}_{1},w^{\prime}_{2})\notin\Omega$. If $L<L_{1}$, then we can also conjugate $A$ with the diagonal matrix $\begin{bmatrix}d_{2}&0\\ 0&d_{1}\end{bmatrix}$. Therefore, we obtain that $\Omega$ is not spirallike with respect to any diagonalizable matrix. ∎
Lemma 6.11.
The domain $\Omega$ is not spirallike with respect to any upper triangularizable matrix.
Proof.
Suppose that there exist $P=\begin{bmatrix}a&b\\ c&d\\ \end{bmatrix}\in Gl(n,\mathbb{C})$ such that $PAP^{-1}=\begin{bmatrix}\lambda&1\\ 0&\lambda\\ \end{bmatrix}$, where $\lambda=\lambda_{1}+i\lambda_{2}$ In this case flow of the matrix starting at $(z_{1},z_{2})$ is as follows:
${}\begin{bmatrix}w_{1}\\ w_{2}\\ \end{bmatrix}$ $=e^{At}\begin{bmatrix}z_{1}\\ z_{2}\\ \end{bmatrix}=\begin{bmatrix}e^{\lambda t}z_{1}+\frac{t}{\mu}e^{\lambda t}(cdz_{1}+d^{2}z_{2})\\ e^{\lambda t}z_{2}-\frac{t}{\mu}e^{\lambda t}(c^{2}z_{1}+dcz_{2})\end{bmatrix}.$ (6.74)
Case 1. $c\neq 0.$
Let $t_{1}>0$ such that $\frac{t_{1}}{|\mu|}|c|^{2}|1+\frac{t_{1}}{\mu}dc|>1$. Choose
$(z_{1},z_{2})=$$\begin{cases}\left(-e^{-\lambda t_{1}}\overline{(1+\frac{t_{1}}{\mu}\cdot dc)}i,0\right)~{}\text{if}~{}{\sf Re}\left(\left(e^{-\lambda t_{1}}\overline{(1+\frac{t_{1}}{\mu}\cdot dc)}i,0\right)\right)>0\\ \left(e^{\lambda t_{1}}\overline{(1+\frac{t_{1}}{\mu}\cdot dc)}i,0\right)~{}\text{if}~{}{\sf Re}\left(\left(e^{-\lambda t_{1}}\overline{(1+\frac{t_{1}}{\mu}\cdot dc)}i,0\right)\right)<0,\\ \end{cases}$ (6.75)
Where $\mu=ad-bc\neq 0$. Clearly, $(z_{1},z_{2})\in\Omega$. Suppose that the integral curve reaches at $(w_{1},w_{2})$ at $t=t_{1}$. Using (6.74), we get that $w_{1}=e^{\lambda t}z_{1}(1+\frac{t_{1}}{\mu}\cdot dc)=\pm|1+\frac{t_{1}}{\mu}\cdot dc|^{2}i.$ Hence, $e^{-\frac{1}{2}{\sf Re}(w_{1})}=1$. We also get that $|w_{2}|=\frac{t_{1}}{|\mu|}|c|^{2}|1+\frac{t_{1}}{\mu}dc|.$ From the choice of $t_{1}>0$ it follows that $|w_{2}|>1$. Hence, $(w_{1},w_{2})\notin\Omega.$
Case 2. $c=0$.
Since $c=0$, we get that $d\neq 0$. We construct the initial point and exit time in this case by following subcases.
Subcase 1. $\lambda_{2}\neq 0.$
We choose $(z_{1},z_{2})=(-2p,\frac{\mu}{d^{2}})$ and $p>0$ large enough such that $|\frac{\mu}{d^{2}}|<e^{-\frac{1}{2}(-2p)}=e^{p}.$ Hence, $(z_{1},z_{2})\in\Omega$. Now take $t_{1}=\frac{\pi}{|\lambda_{2}|}$. From (6.74) we have
$w_{1}=e^{\lambda t_{1}}\cdot(z_{1}+\frac{d^{2}t_{1}}{\mu}z_{2})=e^{\frac{\pi\lambda_{1}}{|\lambda_{2}|}}(2p-\frac{\pi}{|\lambda_{2}|}).$
Thus, we obtain that $e^{-\frac{1}{2}{\sf Re}(w_{1})}=e^{\left(-pe^{\frac{\pi\lambda_{1}}{|\lambda_{2}|}}+\frac{\pi}{2|\lambda_{2}|}e^{(\frac{\pi\lambda_{1}}{|\lambda_{2}|})}\right)}.$ Choosing $p>0$ large enough we get that
$|w_{2}|=e^{\lambda_{1}\frac{\pi}{|\lambda_{2}|}}|\frac{|\mu|}{|d^{2}|}|>e^{-\frac{1}{2}{\sf Re}(w_{1})}.$
Hence, $(w_{1},w_{2})\notin\Omega$.
Subcase 2. $\lambda_{2}=0$ and $\lambda_{1}<0.$
Let $\lambda_{1}=-k$ for some $k>0$ and $(z_{1},z_{2})=(-e^{kt_{1}},\frac{\mu}{d^{2}}e^{2kt_{1}})$, where $t_{1}>0$ is chosen such that $\frac{|\mu|}{|d|^{2}}e^{2kt_{1}-\frac{1}{2}e^{kt_{1}}}<1$. Therefore, $|z_{2}|<e^{-\frac{1}{2}{\sf Re}(z_{1})}$. Hence, $(z_{1},z_{2})\in\Omega$. From (6.74), we get that at time $t_{1}$
$w_{1}$ $=e^{-kt_{1}}z_{1}+\frac{td^{2}}{\mu}e^{-kt_{1}}z_{2}$
$=-1+t_{1}e^{kt_{1}}.$
Hence, we obtain that $e^{-\frac{1}{2}{\sf Re}(w_{1})}=e^{\frac{1}{2}(1-t_{1}e^{kt_{1}})}$ and $|w_{2}|=e^{kt_{1}}\frac{|\mu|}{|d^{2}|}$. Again choose $t_{1}>0$ large enough such that $\frac{|\mu|}{|d|^{2}}e^{kt_{1}+\frac{1}{2}(t_{1}e^{kt_{1}})-\frac{1}{2}}>1.$ Hence, we get that
$\frac{|\mu|}{|d|^{2}}e^{kt_{1}+\frac{1}{2}(t_{1}e^{kt_{1}})-\frac{1}{2}}>1\Longrightarrow|w_{2}|>e^{-\frac{1}{2}{\sf Re}(w_{1})}.$
Therefore, we get that $(w_{1},w_{2})\notin\Omega.$
Subcase 3. $\lambda_{2}=0,$ and $\lambda_{1}>0$.
Let $(z_{1},z_{2})=(-2p,\frac{\mu}{d^{2}}),$ where $p>0$ is chosen such that $(z_{1},z_{2})\in\Omega$. Since $\lambda t+e^{\lambda t}(\frac{t}{2}-p)\to\infty$ as $t\to\infty$. Therefore, we choose $t_{1}>0$ such that $\frac{|\mu|}{|d|^{2}}e^{\lambda_{1}t_{1}-e^{\lambda_{1}t_{1}}(p-\frac{t_{1}}{2})}>1.$ From (6.74) we get that, at $t=t_{1}$, $|w_{2}|=e^{\lambda_{1}t_{1}}\frac{|\mu|}{|d|^{2}}$ and $e^{\frac{1}{2}{\sf Re}(w_{1})}=e^{-pe^{\lambda_{1}t_{1}}+\frac{t_{1}}{2}e^{\lambda_{1}t_{1}}}$. Hence, from the assumption of $t_{1}$ we get that $|w_{2}|e^{\frac{1}{2}{\sf Re}(w_{1})}>1$. Therefore, $(w_{1},w_{2})\notin\Omega$. This shows that $\Omega$ is not spirallike with respect to any uppertriangulable matrix. ∎
Since any $A\in Gl(2,\mathbb{C})$ is either diagonalizable or upper triangulable, hence, from the above computations, we conclude that $\Omega$ is not spirallike with respect to any $A\in Gl(2,\mathbb{C})$. From Lemma 6.5 we get that $\Omega$ is complete hyperbolic. Therefore, Theorem 6.4 is proved.
Example 6.12.
Let $\Omega=\{(z_{1},\ldots,z_{n})\in\mathbb{C}^{n}:{\sf Re}(z_{1})<3,~{}|z_{j}|<e^{-\frac{1}{2}{\sf Re}(z_{1})},2\leq j\leq n\}$. Then $\Omega$ is spirallike domain with respect to the vector field $V(z_{1},z_{2},\ldots,z_{n})=(-2z_{1},-3z_{2}+z_{1}z_{2},-3z_{3}+z_{3}z_{1},\ldots,-3z_{n}+z_{n}z_{1})$.
Proof.
In this case, the flow of the vector field $V$ is
$X(t,(z_{1},z_{2},\cdots z_{n}))=(z_{1}e^{-2t},z_{2}e^{-3t}e^{\frac{z_{1}}{2}(1-e^{-2t})},z_{3}e^{-3t}e^{\frac{z_{1}}{2}(1-e^{-2t})},\\ \ldots,z_{n}e^{-3t}e^{\frac{z_{1}}{2}(1-e^{-2t})}).$
We show that $\Omega$ is spirallike with respect to $V$. Clearly, for all $t\geq 0$, we get that ${\sf Re}(z_{1}e^{-2t})<3$. Let $w_{j}=z_{j}e^{-3t}e^{\frac{z_{1}}{2}(1-e^{-2t})}$ for $j=2,3,\cdots,n$. We now have
$|w_{j}|$ $=|z_{j}|e^{-3t}||e^{\frac{z_{1}}{2}(1-e^{-2t})}|\leq|z_{j}||e^{\frac{z_{1}}{2}(1-e^{-2t})}|$
Since $(z_{1},z_{2})\in\Omega$, it follows that
$|w_{j}|$ $\leq e^{-\frac{1}{2}{\sf Re}(z_{1})}e^{\frac{1}{2}{\sf Re}(z_{1})(1-e^{-2t})}<e^{-\frac{1}{2}{\sf Re}(w_{1})}.$
Therefore, $\Omega$ is a spirallike domain with respect to $V$. By similar computation as in Theorem 6.4, we get that the flow satisfies the condition of Theorem 1.5. ∎
Acknowledgements. Sanjoy Chatterjee is supported by CSIR fellowship (File No-09/921(0283)/2019-EMR-I) and also would like to thank Golam Mostafa Mondal and Amar Deep Sarkar for several discussions. Sushil Gorai is partially supported by a Matrics Research Grant (MTR/2017/000974) of SERB, Dept. of Science and Technology, Govt. of India, for the beginning of this work and by a Core Research Grant (CRG/2022/003560) of SERB, Dept. of Science and Technology, Govt. of India, for the later part of the work.