### v5

# Approximations on certain domains of $\mathbb{C}^{n}$

Abstract.
In this paper, we study the domains in $\mathbb{C}^{n}$ that are invariant under the positive flows of some globally defined, complete holomorphic vector field with a globally attracting fixed point at the origin. Our first result says that such a domain $\Omega$ is always Runge. Next, with an additional assumption on the rate of convergence of the flow, we show that any biholomorphism $\Phi\colon\Omega\to\Phi(\Omega)$, with $\Phi(\Omega)$ is Runge, can be approximated by automorphisms of $\mathbb{C}^{n}$ uniformly on compacts. This generalizes all earlier known theorems in this direction substantially, even when the vector field is linear. As an application of our approximation results, on such domains that are also complete hyperbolic, we show that any Loewner PDE in a complete hyperbolic domain $\Omega$ admits an essentially unique univalent solution with values in $\mathbb{C}^{n}$. We also provide an approximation result for volume preserving biholomorphisms on above domains. We provide several examples of such domains.
Key words and phrases:
Approximation of biholomorphism, Automorphisms of $\mathbb{C}^{n}$, Runge domains, positive time flow invariant domains, Loewner chain
2020 Mathematics Subject Classification:
Primary: 32M17, 32E30, 30C45
1. Introduction
A domain $\Omega\subseteq\mathbb{C}^{n}$ is said to be a Runge domain if every holomorphic function $f:\Omega\to\mathbb{C}$ can be approximated by holomorphic polynomials uniformly on every compact subset of $\Omega$. Runge domains are fundamental in the study of function theory in several complex variables. In one variable, they are the simply connected domains in $\mathbb{C}$, hence, a topological property classifies them. In higher dimensions, no such characterization exists. Characterizing the Runge domain in higher dimensions is a very difficult problem. In this paper, we give a class of domains that are Runge. A holomorphic vector field $V$ on a domain $\Omega\subset\mathbb{C}^{n}$ is a real vector field on $\Omega$ such that
$V(z)=\sum_{j=1}^{n}a_{j}(z)\frac{\partial}{\partial x_{j}}+b_{j}(z)\frac{\partial}{\partial y_{j}}$
such that $(a_{j}+ib_{j})$ is holomorphic function on $\Omega$ for all $j\in\{1,2,\cdots,n\}$, $z_{j}=x_{j}+iy_{j}$ are coordinate of $\mathbb{C}^{n}$. We denote the set of all holomorphic vector fields on $\mathbb{C}^{n}$ by $\mathfrak{X}_{\mathcal{O}}(\mathbb{C}^{n})$. We now state our first theorem.
Theorem 1.1.
Let $F$ be a complete holomorphic vector field with a globally attracting fixed point at the origin. Assume that $\Omega$ is domain which is invariant under positive time flows of $F$ and contains the origin. Then $\Omega$ is a Runge domain.
It is known that any star shaped domain in $\mathbb{C}^{n}$ is Runge (see [Kasimi]). Any star shaped domain with respect to the origin is invariant under the positive time flows of the vector field $-I$. In [Hamada15, Theorem-3.1], Hamada proved the same when the domain is invariant under positive time flows of certain linear vector fields. Hamada calls those domains as spirallike domains. A detailed discussion about such domains are given in Section 2.
The main question that we study in this paper is: For what classes of domains $\Omega$ any biholomorphic map from $\Omega$ into $\mathbb{C}^{n}$ can be approximated by automorphisms of $\mathbb{C}^{n}$? Thanks to the Andersén-Lempert theory, a positive answer of this question will lead to approximating any biholomorphic map by simpler classes of maps, namely the finite composition of shears and overshears. In this paper, we give a large class of domains for which such approximation holds. To proceed further in our discussion, a brief description of the automorphism group of $\mathbb{C}^{n}$ will be useful. Let Aut$(\mathbb{C}^{n})$ be set of all holomorphic mappings $f:\mathbb{C}^{n}\to\mathbb{C}^{n}$ such that $f^{-1}$ exist and also holomorphic. For $n=1$, we know that $(\mathbb{C})$ consists only of affine mappings $z\mapsto az+b,(a,b\in\mathbb{C},\;a\neq 0)$. However, Aut$(\mathbb{C}^{n})$ is huge for $n\geq 2$. Writing the standard coordinates as $(z_{1},z_{2},\cdots z_{n})=(z^{\prime},z_{n})$ on $\mathbb{C}^{n}$ we see that Aut$(\mathbb{C}^{n})$ contains mappings of the form
$(z^{\prime},z_{n})\mapsto(z^{\prime},z_{n}+f(z^{\prime}))$ (1.1)
$(z^{\prime},z_{n})\mapsto(z^{\prime},e^{f(z^{\prime})}z_{n}),$ (1.2)
where $f:\mathbb{C}^{n-1}\to\mathbb{C}$ is any holomorphic mapping. The inverse of the above mapping is obtained by replacing $f$ with $-f$ in (1.1) and (1.2). The mappings of forms (1.1) and their $SL(n,\mathbb{C})$ conjugates are called shears or additive shears, and the mappings (1.2) and their $GL(n,\mathbb{C})$ conjugates are called overshears or multiplicative shears. The subgroup generated by elements of overshears and shears is denoted by $S(n)$, $S_{1}(n)$, respectively. The set of all volume preserving automorphisms of $\mathbb{C}^{n}$ is denoted by $\text{Aut}_{1}(\mathbb{C}^{n})$ and is defined by $\text{Aut}_{1}(\mathbb{C}^{n}):=\{F\in\text{Aut}(\mathbb{C}^{n}):\text{det}(DF(z))\equiv 1\}$. Our next result is about approximating biholomorphic maps on certain domains in $\mathbb{C}^{n}$ which are invariant under the positive time flow of certain holomorphic vector field by compositions of shears and overshears, special members of the $\text{Aut}(\mathbb{C}^{n})$. Before starting our main result let us survey a few things from the existing literature. The study of Aut$(\mathbb{C}^{n})$ was initiated by Rosay and Rudin [RosayRudin]. Questions about the density of the compositions of shears in the class of all volume preserving automorphisms of $\mathbb{C}^{n}$ and the density of compositions of overshears in Aut$(\mathbb{C}^{n})$ were asked in [RosayRudin], which were answered positively by Andersén [Andersen] and by Andersén and Lempert [AndersenLemp] respectively, giving birth of the now-famous Andersén-Lempert theory [Andersen, AndersenLemp] (see [Forstbook] for more details). It is a natural question to ask for characterization of the domains $\Omega\subset\mathbb{C}^{n}$ for which any biholomorphism from $\Omega$ into $\mathbb{C}^{n}$ can be approximated uniformly on compacts of $\Omega$ by automorphisms of $\mathbb{C}^{n}$, i.e. in view of Andersén-Lempert theory by composition of Overshears. Andersén and Lempert [AndersenLemp] considered this question and proved:
Result 1.2.
[AndersenLemp, Theorem 2.1] Let $\Omega$ be a starshaped domain $\mathbb{C}^{n}$ and $\Phi:\Omega\to\mathbb{C}^{n}$ be a biholomorphic map such that $\Phi(\Omega)$ is Runge. Then $\Phi$ can be approximated uniformly on compact subsets by compositions of overshears. If $\det D\Phi\equiv 1$, then $\Phi$ can be approximated uniformly on compact subsets by the composition of shears.
Note that, if a biholomorphism $\Phi:\Omega\to\Phi(\Omega)$, with $\Phi(\Omega)$ Runge, can be approximated uniformly on compacts by elements of Aut$({\mathbb{C}^{n}})$, then in view of the density theorem due to Andersén and Lempert [AndersenLemp, Theorem 1.3], we get that the biholomorphism $\Phi$ can be approximated by the composition of overshears.
In a seminal paper Forstnerič and Rosay [FR1993, Theorem-1.1], [Forst1994, Theorem-1.1] gave a sufficient condition for Runge domain under which any biholomorphic map can be approximated by automorphisms of $\mathbb{C}^{n}$.
Result 1.3 (Forstnerič-Rosay).
Let $\Omega\subset\mathbb{C}^{n}$ be a Runge domain. Assume that $H:[0,1]\times\Omega\to\mathbb{C}^{n}$ is a mapping of class $\mathcal{C}^{p}(0\leq p<\infty)$ such that for each $t\in[0,1]$, $H_{t}=H(t,\cdot)\colon\Omega\to\mathbb{C}^{n}$ is a biholomorphic mapping onto a Runge domain $\Omega_{t}:=H_{t}(\Omega)\subset\mathbb{C}^{n}$ and the map $H_{0}$ can be approximated uniformly on every compact subset of $\Omega$ by automorphisms of $\mathbb{C}^{n}$. Then
  • (i)
    for each compact subset $K\subset\Omega$ and each $\epsilon>0$ there exist a smooth map $\Psi:[0,1]\times\mathbb{C}^{n}\to\mathbb{C}^{n}$ such that $\Psi_{t}:=\Psi(t,\cdot)\in\text{Aut}(\mathbb{C}^{n})$, for every $t\in[0,1]$ and $\|H-\Psi\|_{\mathcal{C}^{p}([0,1]\times K)}<\epsilon$,
  • (ii)
    if we have $H_{0}\in Aut(\mathbb{C}^{n})$, then we can choose $\Psi$ such that $\Psi_{0}=H_{0}$,
  • (iii)
    if $\Omega$ is pseudoconvex, $H^{(n-1)}(\Omega,\mathbb{C})=0$, $\text{det}DH_{t}(z)\equiv 1$ on $\Omega$ and $H_{0}$ can be approximated on $\Omega$ by elements of $\text{Aut}_{1}(\mathbb{C}^{n})$, then every $H_{t}$ can be approximated on $\Omega$ by elements of $\text{Aut}_{1}(\mathbb{C}^{n})$.
This result by Forstnerič and Rosay is the most general result in this direction. The smoothness assumption of the family is lowered in [Forst1994] and [FKu]. But, for a given domain, constructing such an isotopy $\{\Phi_{t}\}$ in Result 1.3 is often difficult. Therefore, it is interesting to know the classes of domains to which the above result can be applied. Hamada [Hamada15] gave a class of domains in the following result.
Result 1.4.
[Hamada15, Theorem 4.2] Let $\Omega$ be a domain containing the origin that satisfies the following conditions: $e^{-tA}w\in\Omega$ for all $w\in\Omega$, for all $t\geq 0$ where $A\in M_{n}(\mathbb{C})$ such that $\inf_{||z||=1}{\sf Re}\langle Az,z\rangle>0$. If $\max_{\lambda\in\sigma(A)}{\sf Re}(\lambda)<2\inf_{||z||=1}{\sf Re}\langle Az,z\rangle$, where $\sigma(A)$ is the spectrum of $A$, then any biholomorphism $\phi:\Omega\to\mathbb{C}^{n}$ with $\phi(\Omega)$ Runge, can be approximated uniformly on compacts of $\Omega$ by automorphisms of $\mathbb{C}^{n}$.
In this paper, we focus on non-linear complete holomorphic vector fields with a globally attracting fixed point at the origin (see Section 2 and the domains that remains invariant under positive time flows of such vector fields. We show under a suitable growth condition on the flows the approximation result holds. To state it precisely, we need the following:
Let $\sigma(A)$ denote the spectrum of the matrix $A$. Let $k_{+}(A):=\max\{{\sf Re}\lambda:\lambda\in\sigma(A)\}$ and $k_{-}(A):=\min\{{\sf Re}\lambda:\lambda\in\sigma(A)\}$. We now present the main result of this article:
Theorem 1.5.
Let $V\in\mathfrak{X}_{\mathcal{O}}{(\mathbb{C}^{n})}$, $n\geq 2$, be a complete holomorphic vector field with a globally attracting fixed point at the origin and $DV(0)=A$. Assume that $\Omega\subseteq\mathbb{C}^{n}$ contains the origin and is invariant under the positive time flows of $V$. Suppose that there exist $\gamma,\alpha>0$ such that
  • (i)
    $\|X(t,z)\|\leq\gamma e^{-\alpha t}\|z\|$ for all $t\geq 0$ and for all $z\in\Omega$, where $X(t,z)$ is the flow of the vector field $V$.
  • (ii)
    $2\alpha+k_{-}(A)$ $>0.$ (1.3)
Then any biholomorphism $\Phi:\Omega\to\Phi(\Omega)\subseteq\mathbb{C}^{n}$ with $\Phi(\Omega)$ Runge can be approximated by automorphisms of $\mathbb{C}^{n}$ uniformly on every compact subset of $\Omega$.
Remark 1.6.
Conditions $(i)$ and $(ii)$ of Theorem 1.5 gives a uniform exponential decay which, in turn, provides the asymptotic behavior of the flows of $V$ as exponential type up to a biholomorphism of $\Omega$ (see Lemma 4.1). This plays a vital role in our proof.
We obtain the following corollary from Theorem 1.5, restricting the vector field to be linear which already gives a substantial generalization to Result 1.4.
Corollary 1.7.
Let $A\in GL(n,\mathbb{C})$ with $2k_{+}(A)<k_{-}(A)$ and $\Omega\subset\mathbb{C}^{n}$ be a domain containing the origin and $P^{-1}e^{tA}Pw\in\Omega$ for all $w\in\Omega$ for some $P\in GL(n,\mathbb{C})$. Then any biholomorphism $\Phi\colon\Omega\to\Phi(\Omega)$ with $\Phi(\Omega)$ Runge can be approximated by elements of $Aut(\mathbb{C}^{n})$ uniformly on every compact subset of $\Omega$.
We obtain the following corollary of the Theorem 1.5.
Corollary 1.8.
Let $\Omega_{1}\subset\mathbb{C}^{n}$, $\Omega_{2}\subset\mathbb{C}^{m}$ be two domains containing the origin and are invariant under positive time flow of complete vector fields $V_{1}\in\mathfrak{X}_{\mathcal{O}}(\mathbb{C}^{n})$ and $V_{2}\in\mathfrak{X}_{\mathcal{O}}(\mathbb{C}^{m})$ respectively. Both the vector fields has a globally attracting fixed point at the origin. Suppose that $X_{1}(t,z)$, $X_{2}(t,w)$ denotes the flow of the vector field $V_{1}$, $V_{2}$. Assume that there exists $\gamma_{1},\gamma_{2},\alpha_{1},\alpha_{2}>0$ such that $\|X_{j}(t,z_{j})\|\leq\gamma_{j}e^{-\alpha_{j}t}\|z_{j}\|$, for all $t\geq 0$, for all $z_{j}\in\Omega_{j}$ for $j\in\{1,2\}$. Suppose that
$\min\{\alpha_{1},\alpha_{2}\}>\max\{\frac{-k_{-}(A)}{2},\frac{-k_{-}(B)}{2}\}.$
Then, any biholomorphism $\Phi\colon\Omega_{1}\times\Omega_{2}\to\Phi(\Omega_{1}\times\Omega_{2})\subseteq\mathbb{C}^{n}\times\mathbb{C}^{m}$ with $\Phi(\Omega_{1}\times\Omega_{2})$ Runge can be approximated by elements of $\text{Aut}(\mathbb{C}^{m+n})$.
Here we pose a conjecture
Conjecture 1.9.
Conclusion of Theorem 1.5 holds true without condition (ii).
A holomorphic symplectic form $\omega$ on $\mathbb{C}^{2n}$ is closed holomorphic $2$-form on $\mathbb{C}^{2n}$ whose highest exterior power $\omega^{n}$ is nowhere vanishing. The standard holomorphic symplectic form on $\mathbb{C}^{2n}$ with coordinates $z_{1},\cdots,z_{n},~{}w_{1},\cdots,w_{n}$ is $\omega=\sum_{j=1}^{n}\,dz_{i}\wedge\,dw_{i}$ Let $\Omega\subset\mathbb{C}^{n}$ be a domain. A holomorphic map $f:\Omega\to\mathbb{C}^{2n}$ is called symplectic if $f^{*}(\omega)=\omega$. A holomorphic vector field $V$ on $\Omega$ is said to be symplectic if $\,d\iota_{V}(\omega)=0$. An automorphism $F:\mathbb{C}^{2n}\to\mathbb{C}^{2n}$ is symplectic if $F^{*}(\omega)=\omega$. We denote the set of all symplectic holomorphism by $Aut_{\omega}(\mathbb{C}^{n})$.
We now present an approximation result where a volume preserving biholomorphisms and symplectic biholomorphism that are defined on a domain that is invariant under positive time flows of some vector field, by volume preserving automorphism and symplectic automorphisms of $\mathbb{C}^{n}$ respectively.
Theorem 1.10.
Let $V\in\mathfrak{X}_{\mathcal{O}}{(\mathbb{C}^{n})}$, $n\geq 2$, be a complete vector field with a globally attracting fixed point at the origin and $DV(0)=A$. Assume that $\Omega\subseteq\mathbb{C}^{2n}$ is a pseudoconvex domain containing the origin and invariant under positive time flows of $V$. Suppose that there exist $\gamma,\alpha>0$ such that $\|X(t,z)\|\leq\gamma e^{-\alpha t}\|z\|$ for all $t\geq 0$ and for all $z\in\Omega$, where $X(t,z)$ is the flow of the vector field $V$. Assume that $02\alpha+k_{-}(A)>$ 0 . Let $\Phi:\Omega\to\Phi(\Omega)$ be a biholomorphism with $\Phi(\Omega)$ Runge. Then the following holds:
  • (i)
    If $div_{\omega^{n}}V(z)$ is non zero constant and $\text{det}(D\Phi(z))\equiv 1$, where $\omega^{n}$ is standard holomorphic $(n,0)$ form $\,dz_{1}\wedge\,dz_{2}\wedge\cdots\wedge\,dz_{n}$, then $\Phi$ can be approximated by elements of $Aut_{1}(\mathbb{C}^{n})$ uniformly on every compact subset of $\Omega$.
  • (ii)
    Let $n=2m$ and $(z_{1},z_{2},\ldots,z_{m},w_{1},w_{2},\ldots,w_{m})$ be a coordinate of $\mathbb{C}^{n}$ and $\omega=\sum_{j=1}^{m}\,dz_{i}\wedge\,dw_{i}$ is a symplectic form on $\mathbb{C}^{n}$. If $V$ is symplectic vector field and $\Phi^{*}(\omega)=\omega$, then $\Phi$ can be approximated by elements of $Aut_{\omega}(\mathbb{C}^{n})$ locally uniformly on $\Omega$.
Our next result is an application of Theorem 1.5 to the theory of Loewner PDE. The study was motivated by Arosio-Bracci-Wold [ABFW13] and Hamada [Hamada15]. We present a brief introduction to Loewner PDE before presenting our result. The Loewner ordinary differential equation in the unit disc was introduced by Loewner[Loew1923] in 1923 to study Bieberbach’s conjecture about the coefficients of a univalent map on the unit disc. The theory in one variable was developed later by Kufarev [Kuf] and Pommerenke [Pomm1965] (see [MBCD2010] for a brief survey of Loewner differential equations). The Loewner theory has been used to prove many deep results in geometric function theory [GKbook]. It was one of the main ingredients in de Brange’s proof of Bieberbach conjecture [dB1985]. Then there were several generalizations; Some of them are: chordal Loewner theory [KSS1968] and Loewner theory in several complex variables [Pfal1974, Pfal1975]. The theory of the Loewner equation in several complex variables was extended by Pfaltzgraff [Pfal1974, Pfal1975], and later studied by many others [Poreda1991, GHK2002, GHKK2008, HK2003, GKP2005]. A map $H:\Omega\times[0,\infty)\to\mathbb{C}^{n}$ is called a Herglotz vector field of order $d$ on $\Omega$ if
  • (i)
    For each $z\in\Omega$, $H(z,\cdot):[0,\infty)\to\Omega$ is measurable.
  • (ii)
    For each $s\in[0,\infty)$, $H(.,s)$ is holomorphic vector field and $H(\cdot,s)$ is a semicomplete holomorphic vector field for almost all $s\in[0,\infty)$, i.e. for each $z_{0}\in\Omega$, the ODE defined by
    $\dfrac{dw}{dt}=H(w(t),s),\;\;\;w(0)=z_{0}$
    has a solution for all $t\in[0,\infty)$ for almost all $s\in[0,\infty)$.
  • (iii)
    For each $t_{0}\in[0,\infty)$, and for each compact $K\subset\Omega$, there exists a function $c_{t_{0}}^{K}\in L^{d}([0,t_{0}],[0,\infty))$ such that $||H(z,t)||\leq c_{t_{0}}^{K}$ for all $z\in K$ almost all $t\in[0,t_{0}]$.
For a Herglotz vector field $H$ of order $d\in[1,\infty]$ on a complete hyperbolic domain $\Omega$ the Loewner PDE is the following equation:
$\dfrac{\partial f_{t}(z)}{\partial t}$ $=-df_{t}(z)H(z,t).$ (1.4)
A solution of the Loewner PDE is a family of holomorphic maps $\{f_{t}\}_{t\in[0,\infty)}$ from $\Omega$ to a complex manifold $M$ which is locally absolutely continuous on $[0,\infty)$ satisfies (1.4) and $t\mapsto f_{t}$ is continuous in the topology of holomorphic maps from $\Omega$ to $M$. Such a family of maps is closely related to evolution family $\{\varphi_{s,t}\}$ that satisfies Loewner ODE (1.5)
$\dfrac{\partial\varphi_{s,t}(z)}{\partial t}$ $=H(\varphi_{s,t}(z),t),~{}~{}a.e~{}~{}t\in[s,\infty)$ (1.5)
in the following sense:
[…]
The main open problem in the field is; To solve whether, on a complex complete hyperbolic domain, there is a solution to Loewner PDE with values in $\mathbb{C}^{n}$. Solution of such Loewner PDE with values in certain abstract complex manifold has already been found in [BCD2009]. Recently the connection between the approximation of biholomorphic maps and the solution of Loewner PDE has been found [ABFW13]. As an application of Theorem 1.5 we state the following.
Theorem 1.11.
Let $n\geq 2$. Let $\Omega\subset\mathbb{C}^{n}$ be a complete hyperbolic domain containing the origin and invariant under positive time flows of the vector field $V\in\mathfrak{X}_{\mathcal{O}}{(\mathbb{C}^{n})}$ with a globally attracting fixed point at the origin and $DV(0)=A$. Suppose that there exist $\gamma,\alpha>0$ such that $\|X(t,z)\|\leq\gamma e^{-\alpha t}\|z\|$, for all $t\geq 0$ and for all $z\in\Omega$, where $X(t,z)$ is the flow of the vector field $V$. Assume that
$2\alpha+k_{-}(A)$ $>0.$ (1.7)
Let $G:\Omega\times\mathbb{R}^{+}\to\mathbb{C}^{n}$ be a Harglotz vector field of order $d\in[1,\infty]$. Then there exist a Loewner chain $f_{t}:\Omega\to\mathbb{C}^{n}$ of order $d$ which solves the Loewner PDE
$\frac{\partial f_{t}}{\partial t}(z)$ $=-df_{t}(z)(G(z,t))~{}~{}a.e.~{}~{}t\geq 0~{}~{}\forall z\in\Omega.$ (1.8)
Moreover, $R=\cup_{t\geq 0}f_{t}(\Omega)$ is a Runge and Stein domain in $\mathbb{C}^{n}$ and any other solution to (1.8) with values in $\mathbb{C}^{n}$ is of the form $(\varphi\circ f_{t})$ for a suitable map $\varphi:R\to\mathbb{C}^{n}$.
Theorem 1.11 gives a partial answer to the open question whether any Loewner PDE in a complete hyperbolic domain in $\mathbb{C}^{n}$ admits a univalent solution in $\mathbb{C}^{n}$. This generalizes results of [ABFW13, Hamada15] Other partial positive answers are [ABFW13, Hamada15, Duren].
In the final section of the paper, we provide several examples of domains that satisfy the conditions of our theorems. Thus, those domains allow to approximate any biholomorphic maps onto a Runge domain by automorphisms of $\mathbb{C}^{n}$. we also provide an example of a complete hyperbolic Hartogs domain which is invariant with respect to positive time flow of a nonlinear holomorphic vector field but not with respect to any linear vector field.
2. Technical Preliminaries
In this section, we first discuss about the holomorphic vector fields and their flows. Then we make a comparison of domains that were considered by Hamada in [Hamada15] with ours when the vector field is linear. Let $\Omega\subset\mathbb{C}^{n}$ and $V$ be a holomorphic vector field on $\Omega$. To find the real flow of $V$ we consider the following equation
$\dot{x}(t)$ $=V(x(t)),\;\;\;x(0)=x_{0},$ (2.1)
We consider the holomorphic differential equation associated to (2.1) defined by
$\dot{x}(z)$ $=\widetilde{V}(x(z))$
$x(0)$ $=z_{0},$ (2.2)
[…]
$\frac{d\phi}{dt}=V(\phi(t));~{}~{}\frac{d\psi}{ds}=i\cdot V(\psi(s)),$
where $z=t+is$, $X(0)=z_{0}$. Thus, we have $X(t+is,z)=\phi_{t}(\psi_{s}(z_{0}))$.
[…]
In this paper we will consider the holomorphic vector fields with the following properties:
  • i)
    The vector field is defined on the whole of $\mathbb{C}^{n}$.
  • ii)
    The vector field is $\mathbb{R}$-complete.
  • iv)
    The vector field has a globally attracting fixed point.
Since the above properties play very important roles in our results, we just mention those here briefly before going further in our exposition. We consider the following system of differential equations: Let $E\subset\mathbb{R}^{n}$ be an open set containing the origin and $f:E\to\mathbb{R}^{n}$ is continuously differentiable mappings such that $f(0)=0$
$\frac{dX(t)}{dt}$ $=f(X(t)),~{}~{}X(0)=x_{0}.$ (2.3)
Assume that the solution of the system (2.3) exists for every $t\geq 0$, $\forall x_{0}\in E$.
  • Then origin is said to be a stable equilibrium point of the system (2.3) if for every $\epsilon>0$ there exist $\delta>0$ such that $X(t,x_{0})\in B(0,\epsilon)$ for every $x_{0}\in B(0,\delta)$ for every $t\geq 0$.
  • The origin is said to be globally attracting equilibrilum point or globally asymptotically stable equilibrium point if $E=\mathbb{R}^{n}$ , the origin is stable equilibrium point such that $\lim_{t\to\infty}X(t,x_{0})=0$ for all $x_{0}\in\mathbb{R}^{n}$. A vector field is said to be globally asymptotically stable if the origin is a globally asymptotically stable equilibrium point.
Any holomorphic vector field $F$ on $\mathbb{C}^{n}$ can also be seen as a holomorphic map from $F:\mathbb{C}^{n}\to\mathbb{C}^{n}$. For such a vector field $F$, we denote $DF(a)$ by the derivative matrix viewing it as a map.
Remark 2.1.
Let for $A\in M_{n}(\mathbb{C})$, $m(A):=\inf_{\|z\|=1}{\sf Re}\langle Az,z\rangle$. The condition in Result 1.4, becomes $k_{+}(A)<2m(A)$. As we had discussed earlier the differential equation that we consider in this case is:
$\frac{dX}{dt}$ $=AX(t)$ (2.4)
$X(0)$ $=z.$
The differential equation that Hamada [Hamada15] considered is the equation introduced by Gurganus [Gur] :
$\frac{dX}{dt}$ $=-AX(t)$ (2.5)
$X(0)$ $=z.$
The Runge property does not get affected by this change of definition. But, the approximation of the biholomorphic map of $\Omega$ by automorphisms of $\mathbb{C}^{n}$ get affected very much by this change. Hamada’s condition $k_{+}(A)<2m(A)$ (the condition in Result 1.4) does not remain invariant under the above change of definition. In general $m(A)$ is not a similarity invariant. Hence, the Corollary 1.7 really enlarges the class of domains for which every biholomorphism with Runge image can be approximated by elements of $\text{Aut}(\mathbb{C}^{n})$ locally uniformly on $\Omega$, even if the vector field is linear (see Example 6.2 for an explicit example).
Next, we state a result due to Arosio-Bracci-Hamada-Kohr [ABHK2013] which gives the existence of a solution of Loewner PDE with values in a complex manifold on a complete hyperbolic domain. A manifold is said to be completely hyperbolic if it forms a complete metric space with respect to the Kobayashi distance (see [Kobaya] for details.)
Result 2.2.
[ABHK2013, Theorem 5.2] Let $M$ be a complete hyperbolic complex manifold of dimension $n$. Let $G:M\times\mathbb{R}^{+}\to TM$ be a Herglotz vector field of order $d$. Then there exists a Loewner chain $f_{t}:M\to N$ of order $d$ which solves the Loewner PDE
$\frac{\partial f_{t}}{\partial t}(z)$ $=-df_{t}(z)(G(z,t))~{}~{}a.e.~{}t\geq 0~{}~{}\forall z\in M,$ (2.6)
where $N=\cup_{t\geq 0}f_{t}(M)$ is a complex manifold of dimension $n$ and any other solution to (2.6) with values in a complex manifold $Q$ is of the form $(\Lambda\circ f_{t})$ where $\Lambda:N\to Q$ is holomorphic.
We now state a result by Thai-Duc [Dudoc] which gives a characterization of the complete hyperbolic Hartogs domain. This plays a vital role in Example 6.4.
Result 2.3.
[Dudoc] Let $X$ be a complex space and $\phi$ a plurisubharmonic function on $X$. Assume that for every $x\in X$, there exists an open neighbourhood $U$ of $x$ in $X$ and a sequence $\{h_{j}\}$ of holomorphic functions on $U$ and a sequence $\{c_{j}\}$ of real numbers in the interval $(0,1)$ such that the sequence $\{c_{j}\log|h_{j}|\}$ converges uniformly on compact subsets of $U$ to the function $\phi$. Then the Hartogs domain $\Omega_{\phi}(X)=\{(x,z)\in X\times\mathbb{C}:|z|<e^{-\phi(x)}\}$ is complete hyperbolic if and only if $X$ is complete hyperbolic and $\phi$ is continuous on $X$.
Next, we state a result due to Hamada [Hamada15, Theorem 5.2] which describes a condition under which a pair of complex manifolds is a Runge pair.
Result 2.4.
[Hamada15] Let $M$ be a complete hyperbolic Stein Manifold and $Q$ be a complex manifold of the same dimension. Let $(f_{t}:M\to Q)$ be a Loewner chain of order $d\in[1,\infty]$. Then $(f_{s_{1}}(M),f_{s_{2}}(M))$ is a Runge pair for $0\leq s_{1}<s_{2}$.
The next result due to Lloyd [Lloyd, Theorem 3] will be used to prove Lemma 4.1.
Result 2.5.
[Lloyd, Theorem 3] Let $D$ be a bounded domain in $\mathbb{C}^{n}$. Suppose $f$ and $g$ are two continuous maps from $\overline{D}$ to $\mathbb{C}^{n}$ which are holomorphic on $D$ such that
$\|g(z)\|<\|f(z)\|~{}\forall z\in\partial D.$
[…]
We need the following lemma for the approximation result for the linear case.
Result 2.6.
[SimDav, Page-311] For every $\epsilon>0$ there exist $\rho(\epsilon),m(\epsilon)>0$, such that
$\|e^{-tA}z\|$ $\leq\rho(\epsilon)e^{-(k_{-}(A)-\epsilon)t}\|z\|$
$\|e^{tA}z\|$ $\leq m(\epsilon)e^{(k_{+}(A)+\epsilon)t}\|z\|$
for all $t\geq 0$, and for all $z\in\mathbb{C}^{n}$.
We now state and prove a lemma that will be useful for our proofs.
Lemma 2.7.
Let $\Omega$ be a domain in $\mathbb{C}^{n}$ and $\varphi:\Omega\to\mathbb{C}^{n}$ be a biholomorphic map with $\varphi(\Omega)$ Runge. Assume that $\mu:\Omega\to\mathbb{C}^{n}$ is a map such that $\mu(\Omega)$ is Runge and $\mu(\Omega)\subset\Omega$. Then $\lambda\circ\varphi\circ\mu(\Omega)$ is a Runge domain for every $\lambda\in Aut(\mathbb{C}^{n})$.
Proof.
It is enough to prove that $\varphi\circ\mu(\Omega)$ is Runge. Let $f\in\mathcal{O}(\varphi\circ\mu(\Omega))$ and $K$ be a compact subset of $\varphi\circ\mu(\Omega)$. Hence, $\varphi^{-1}(K)$ is a compact subset of $\mu(\Omega)$. Clearly, $f\circ\varphi\in\mathcal{O}(\mu(\Omega))$. Since $\mu(\Omega)$ is Runge, there exists a sequence of polynomials $\{p_{n}\}$ such that $p_{n}\to f\circ\varphi$ uniformly on $\varphi^{-1}(K)$. Hence, $p_{n}\circ\varphi^{-1}\to f$ uniformly on K. Since $\varphi^{-1}:\varphi(\Omega)\to\Omega$ is a holomorphic and $\varphi(\Omega)$ is Runge, there exists a sequence $\{q_{j}\}$ of polynomial maps from $\mathbb{C}^{n}\to\mathbb{C}^{n}$ such that $q_{j}\to\varphi^{-1}$ uniformly on $K$. By a diagonal-type process, we see that $f$ can be approximated by polynomials uniformly on $K$. Hence, $\varphi\circ\mu(\Omega)$ is Runge. ∎
We now state a result from [Forstbook, Page-110] that will be useful in the context of approximation by volume preserving automorphisms. Recall that, for holomorphic vector field $F=\sum_{j=1}^{n}F_{j}\frac{\partial}{\partial z_{j}}\in\mathfrak{X}_{\mathcal{O}}(\mathbb{C}^{n})$ and standard holomorphic $(n,0)$ form $\omega=\,dz_{1}\wedge\,dz_{2}\wedge\cdots\wedge\,dz_{n}$,
$\text{div}_{\omega}F=\frac{\partial F_{1}}{\partial z_{1}}+\frac{\partial F_{2}}{\partial z_{2}}+\cdots+\frac{\partial F_{n}}{\partial z_{n}}.$
Lemma 2.8.
Let $\Omega\subset\mathbb{C}^{n}$ be an open set and $F:\Omega\to\mathbb{C}^{n}$ be a complete $\mathcal{C}^{1}$- vector field with flow $X(t,z)$. Then
$\frac{d}{dt}\text{det}D_{z}(X(t,z))=\text{div}(F(X(t,z)))\cdot\text{det}~{}D_{z}(X(t,z)).$
3. Runge property
Before proving the main result of this section we state and prove a proposition that will be useful in our proof of Theorem 1.1.
Proposition 3.1.
Let $F\in\mathfrak{X}_{\mathcal{O}}{(\mathbb{C}^{n})}$ be a complete vector field such that the origin is a globally attracting fixed point and $\Omega$ be a positive time flow invariant domain containing the origin. Then, for every compact subset $K\subset\mathbb{C}^{n}$, there exists a real number $M_{K}>0$ such that $X(t,z)\in\Omega$ for all $t>M_{K}$ and for all $z\in K$.
Proof.
Let $r>0$ such that $B(0,r)\subset\Omega$. Since the origin is the globally attracting fixed point of $F$, there exists $\delta>0$ such that $X(t,z)\in B(0,r)$ $\forall t\geq 0$ and $\forall z\in B(0,\delta)$. We also have $\lim_{t\to\infty}X(t,z)\to 0$ for every $z\in\mathbb{C}^{n}$. Therefore, for every $z\in K$, there exist $t_{z}>0$, such that $X(t_{z},z)\in B(0,\frac{\delta}{3})$. Hence, from choice of $\delta>0$, we have
$X(t,z)\in B(0,r)\subseteq\Omega,\;\;\forall t\geq t_{z}.$
Let $K\subset\mathbb{C}^{n}$ be a compact subset. For every $z\in K$ there exist $r_{z},t_{z}>0$ such that $X(t,w)\in B(0,\frac{\delta}{3})$, for all $w\in B(z,r_{z})$ and $t\geq t_{z}$. Since $K$ is compact, hence, there exists $m\in\mathbb{N}$ and $\{z_{1},z_{2},\ldots,z_{m}\}\subset K$ such that $K\subset\cup_{j=1}^{m}B(z_{j},r_{z_{j}})$. Let $M_{K}=\max\{t_{z_{1}},t_{z_{2}},\ldots,t_{z_{m}}\}$. Therefore, $X(t,z)\in B(0,r)$, for all $z\in K$ , for all $t>M_{K}$.
We now present the proof of Theorem 1.1.
Proof of Theorem 1.1.
Let $f\in\mathcal{O}(\Omega)$ and $K\subset\Omega$ be a given compact subset and $\mu$ is a complex Borel measure on $K$. It is enough to show that if, for every $p\in\mathbb{C}[z_{1},z_{2},\ldots,z_{n}]$,
$\int_{K}p(z_{1},z_{2},\ldots,z_{n})\;\mathrm{d}\mu(z)=0$
then
$\int_{K}f(z)\;\mathrm{d}\mu(z)=0.$
Consider
$V_{K}:=\{e^{-\tau}:\tau\in\mathbb{C},X(\tau,z)\in\Omega,\forall z\in K\}\cup\{0\}\subset\mathbb{C}.$
Since $F$ is a $\mathbb{R}$-complete vector field, hence, using [Forst1996, Corollary 2.2], we get that $F$ is also $\mathbb{C}$-complete. Therefore, $X(\tau,z)$ is well defined for every $\tau\in\mathbb{C}$. We now show that $V_{K}$ is open and starshaped with respect to the origin. If $w\in V_{K}$, then $w=e^{-\tau_{1}}$ for some $\tau_{1}\in\mathbb{C}$ with $X(\tau_{1},z)\in\Omega$ for all $z\in K$. For any $t\in(0,1)$ we have
$tw=e^{\ln{t}}\cdot e^{-\tau_{1}}=e^{-(-\ln{t}+\tau_{1})}.$
Clearly, $-\ln{t}>0$. Since $\Omega$ is an invariant domain under the positive time flows of the vector field $F$, hence, for every $z\in K$, it follows that
$X(-\ln{t}+\tau_{1},z)=X_{-\ln{t}}((X(\tau_{1},z)))\in\Omega.$
Consequently, $tw\in V_{K}$ for all $t\in(0,1)$. Clearly, $tw\in\Omega$ for $t=0,1$. Therefore, $V_{K}$ is starshaped. Hence, it is simply connected. Now we show that $V_{K}$ is an open subset. For that, we first show that $S=\{\tau\in\mathbb{C}|~{}~{}\forall~{}z\in K,~{}X(\tau,z)\in\Omega\}$ is an open subset of $\mathbb{C}$. Let $\tau_{1}\in S$. By definition of $S$ we get that $X(\tau_{1},z)\in\Omega$ for every $z\in K$. Let $\epsilon_{z}>0$ such that $B(X(\tau_{1},z),\epsilon_{z})\subset\Omega$. Since $X$ is a continuous map, hence, there exists $\delta_{z,\tau_{1}}>0$ such that $X(\tau,w)\in B(X(\tau_{1},z),\epsilon_{z})$ whenever $|\tau-\tau_{1}|<\delta_{z,\tau_{1}}$ and $\|w-z\|<\delta_{z,\tau_{1}}$. Since $K$ is compact, hence, we get that $K\subset\cup_{i=1}^{m}B(z_{i},\delta_{z_{i},\tau_{1}})$ for some $z_{1},z_{2},\ldots,z_{m}\in K$. Therefore, choosing $0<\delta_{1}<\min\{\delta_{z_{1},\tau_{1}},\delta_{z_{2},\tau_{1}},\ldots,\delta_{z_{m},\tau_{1}}\}$ we obtain that $B(\tau_{1},\delta_{1})\subset S$. By open mapping theorem, we get that all the points of the form $e^{-\tau}$ are interior points of $V_{K}$. We now show that there exists $r>0$ such that $B(0,r)\subset V_{K}$. Now for every non zero complex number $\tau\in\mathbb{C}$, we have $\tau=e^{-(\ln{\frac{1}{|\tau|}}-i\arg{\tau})},$ where $-\pi<\arg{\tau}\leq\pi$. Hence, we get that $X(\ln(|\frac{1}{\tau}|-i\arg{\tau})),z)=X_{\ln(|\frac{1}{\tau}|)}(X(-i\arg{\tau},z))$. Clearly, the set $K^{\prime}=\{X(-it,z):-\pi\leq t\leq\pi,~{}z\in K\}$ is compact. Since $\ln{|\dfrac{1}{\tau}|}\to\infty$ as $\tau\to 0$, hence, we infer from Proposition 3.1, that there exists $r>0$ such that $X(\ln{\frac{1}{|\tau|}},z^{\prime})\in\Omega$, whenever $|\tau|<r$ and $z^{\prime}\in K^{\prime}$. We also obtain that $X(-i\arg{\tau},z)\in K^{\prime}$ for all $z\in K$. Hence, for any $\tau\in B(0,r)$ and $z\in K$, we have $X(\ln{\frac{1}{|\tau|}},X(-i\arg{\tau},z))\in\Omega$. Therefore, $V_{K}$ is an open subset. Let $\widetilde{V}_{K}=V_{K}\setminus(-\infty,0]$. For each $\tau\in\widetilde{V}_{K}$, define $f_{\tau}:K\to\mathbb{C}$ by $f_{\tau}(z)=f(X(-\ln{\tau},z))$. Here $f(X(-\ln{\tau},z))$ is continuously differentiable with respect to the variable $\tau$ for each $z\in K$. Hence, the function $\phi:\widetilde{V}_{K}\to\mathbb{C}$ defined by
$\phi(\tau)=\int_{K}f_{\tau}(z)\;\mathrm{d}\mu(z)$
is holomorphic on $\widetilde{{V}}_{K}$. Since $K$ is compact in $\mathbb{C}^{n}$, there exist $R>0$ such that $K\subset\overline{B(0,R)}$. In view of Proposition 3.1, we get that $\exists~{}\delta>0$ such that $X(-\ln{\tau},y)\in\Omega$ for all $\tau\in\{\tau\in\mathbb{C}:|\tau|<\delta\}\cap\widetilde{{V}}_{K}$ and for all $y\in\overline{B(0,R)}$. Therefore, $f(X(-\ln{\tau},z))\in\mathcal{O}(B(0,R))$ for every $\tau\in\{|t|<\delta\}\cap\widetilde{{V}}_{K}$. Hence, there exists a sequence of polynomial $p_{n}$ such that $p_{n}$ uniformly converges to $f_{\tau}$ on every compact subset of $B(0,R)$, particularly on $K$. Therefore, for every $\tau$ in $V_{K}$ with $|\tau|<\delta$, it follows that
$\lim_{n\to\infty}\int_{K}p_{n}(z)\;\mathrm{d}\mu(z)=\int_{K}f((X(-\ln{\tau},z)))\;\mathrm{d}\mu(z).$
By assumption $\int_{K}p_{n}(X(-\ln{\tau},z))\;\mathrm{d}\mu(z)=0$. Hence, for all $\tau\in V_{K}$ with $|\tau|<\delta$, we obtain that
$\int_{K}f((X(-\ln{\tau},z)))\;\mathrm{d}\mu(z)=0.$
Hence, $\phi$ vanishes identically on an open ball in $\widetilde{V}_{K}$. Since $\phi$ is a holomorphic map hence, it vanishes identically on $\widetilde{V}_{K}$. Therefore, $\phi(1)\equiv 0$ and this implies that $\int_{K}f(z)\;\mathrm{d}\mu(z)=0$. Hence, $\Omega$ is Runge. ∎
4. Approximation of biholomorphisms
We begin this section with the following lemma which will be used in the proof of Theorem 1.5. This lemma says under the conditions of Theorem 1.5 the asymptotic behavior of the flows of the vector field are like exponential up to a biholomorphism of the domain $\Omega$.
Lemma 4.1.
Let $V\in\mathfrak{X}_{\mathcal{O}}{(\mathbb{C}^{n})}$, $n\geq 2$, be a complete holomorphic vector field with $V(0)=0$ and $DV(0)=A$. Assume that $\Omega\subseteq\mathbb{C}^{n}$ contains the origin and is invariant under the positive time flows of $V$. Suppose that there exist real numbers $\gamma,\alpha>0$ such that
  • (i)
    $\|X(t,z)\|\leq\gamma e^{-\alpha t}\|z\|$ for all $t\geq 0$ and for all $z\in\Omega$, where $X(t,z)$ is the flow of the vector field $V$.
  • (ii)
    $02\alpha+k_{-}(A)>$ 0 .
Then there exists a biholomorphic map $F$ on $\Omega$ into $\mathbb{C}^{n}$ such that $e^{-tA}\circ X_{t}(z)$ converges to $F(z)$ as $t\to\infty$ uniformly on compact subsets of $\Omega$.
Proof.
Fix a compact set $K\subset\Omega$. We will show that for every $\varepsilon>0$, there exists $N_{\varepsilon,K}\in\mathbb{N}$ such that $\|e^{-t_{1}A}X_{t_{1}}(z)-e^{-t_{2}A}X_{t_{2}}(z)\|<\varepsilon$ for every $t_{1},t_{2}>N_{\varepsilon,K}$ and for all $z\in K$. This ensures the uniform convergence of $\{e^{-tA}\circ X_{t}\}$ on $K$ as $t\to\infty$. Since $V$ is a holomorphic vector field on $\mathbb{C}^{n}$, we can write $V(z)=Az+R(z)$, where $R(z)=O(\|z\|^{2})$ as $z\to 0$ i.e., there exist $M,\delta>0$ such that $\|R(z)\|\leq M\|z\|^{2}$ for all $z\in B(0,\delta)$. Since $V$ is globally asymptotically stable vector field on $\mathbb{C}^{n}$, for every $z\in\mathbb{C}^{n}$ $\|X(t,z)\|\to 0$ as $t\to\infty$. We note that the convergence is uniform on every compact subset of $\mathbb{C}^{n}$. Hence, there exists $\widetilde{N_{K}}>0$ such that
$\|R(X(t,z))\|$ $<M\|X(t,z)\|^{2},\forall t>\widetilde{N_{K}},\forall z\in K.$ (4.1)
Define $f_{t}:\mathbb{C}^{n}\to\mathbb{C}^{n}$ by $f_{t}(z):=e^{-tA}X_{t}(z)$. Clearly, $f_{t}\in Aut(\mathbb{C}^{n})$ for all $t>0$. The family is also smooth in $t$-variable. Using the mean value theorem in the $t$ variable, we obtain that
$\|f_{t}(z)-f_{s}(z)\|$ $\leq\left\|\frac{df_{t}(z)}{\,dt}\bigg{|}_{t=s+a_{s}}\right\||t-s|,$ (4.2)
where $s\leq s+a_{s}\leq t$. We now find an estimate for $\left\|\frac{df_{t}(z)}{\,dt}\right\|$.
$\frac{df_{t}(z)}{\,dt}$ $=-Ae^{-tA}(X(t,z)+e^{-tA}(A(X(t,z)))+e^{-tA}(R(X(t,z)))$
$=e^{-tA}R(X(t,z)),$
Hence,
$\|\frac{df_{t}(z)}{\,dt}\bigg{|}_{s+a_{s}}\|$ $=\|e^{-(s+a_{s})A}R(X((s+a_{s}),z))\|.$ (4.4)
By assumption, choose $\sigma>0$ such that $02\alpha+k_{-}(A)-\sigma>$ 0. In view of Result 2.6, we deduce that
$\|\frac{df_{t}(z)}{\,dt}\bigg{|}_{s+a_{s}}\|$ $\leq\rho(\varepsilon_{1})e^{-(s+a_{s})(k_{-}(A)-\sigma)}\|R(X((s+a_{s}),z))\|.$ (4.5)
Using (4.1) and (4.5), we get that
$\|\frac{df_{t}(z)}{\,dt}\bigg{|}_{s+a_{s}}\|$ $\leq M\rho(\sigma)e^{-(s+a_{s})(k_{-}(A)-\sigma)}\|(X((s+a_{s}),z)\|^{2}\;\;\forall s>\widetilde{N_{K}}\;\;\forall z\in K.$ (4.6)
Using Assumption (i), equations (4.2), (4.1) and (4.6) we obtain that
$\|f_{t}(z)-f_{s}(z)\|\leq M^{\prime}e^{-Cs}|t-s|\;\;\forall t>s>\widetilde{N_{K}}\;\;\forall z\in K,$ (4.7)
where $C=2\alpha+k_{-}(A)-\sigma>0$ and $M^{\prime}:=\sup_{z\in K}M\gamma^{2}\rho(\varepsilon_{1})\|z\|^{2}$. Since there is an exponential decay on the right hand side of (4.7), it follows that for every $\varepsilon>0$, there exists $N_{\varepsilon,K}\in\mathbb{N}$ such that $\|e^{-t_{1}A}X_{t_{1}}(z)-e^{-t_{2}A}X_{t_{2}}(z)\|<\varepsilon$ for every $t_{1},t_{2}>N_{\varepsilon,K}$ and for all $z\in K$. This proves that every subsequence of $\{f_{t}\}$ forms a Cauchy sequence on every compact subset of $\Omega$. Hence, there exists $F\in\mathcal{O}(\Omega,\mathbb{C}^{n})$ such that $\lim_{t\to\infty}e^{-tA}\circ X_{t}(z)=F(z)$ and the convergence is uniform on every compact subsets of $\Omega$.
Next, we show that $F$ is a biholomorphic map from $\Omega$ onto its range. We have $f_{t}=e^{-tA}\circ X_{t}\in\text{Aut}(\mathbb{C}^{n})$ and $D(e^{-tA}X_{t})(0)=e^{-tA}DX_{t}(0)=e^{-tA}e^{tA}=I_{n}$ for all $t\geq 0$. Since $Df_{t}(0)\to DF(0)$ as $t\to\infty$, hence, we obtain that $DF(0)=I_{n}$. Since $\text{det}(D(e^{-tA}X_{t})(z))\neq 0$ for all $z\in\mathbb{C}^{n}$, therefore, by Hurwitz theorem in several variables, we conclude that $\text{det}(DF(z))$ is either non-zero on $\Omega$ or identically zero on $\Omega$. But, $\text{det}(DF(0))=1$ implies that $\det(DF(z))\neq 0$ for all $z\in\Omega$. Therefore, by invoking inverse function theorem, we obtain that $F$ is a local biholomorphism. We need to show that $F$ is injective. Suppose not. Then, there exist two distinct points $z_{1},z_{2}$ in $\Omega$ such that $F(z_{1})=F(z_{2})$. Since $F$ is locally injective, hence, there exists $r>0$ such that $F$ is injective on $B(z_{1},r)$. Shrinking $r>0$, if needed, we assume that $z_{2}\notin\overline{B(z_{1},r)}$. Define $\widetilde{F}_{t}:=f_{t}-f_{t}(z_{2})$ and $\widetilde{F}:=F-F(z_{2})$. Clearly, $\widetilde{F}^{-1}(0)\cap\partial B(z_{1},r)=\emptyset$. Choose $\alpha>0$ such that $\inf_{\partial B(z_{1},r)}|\widetilde{F}|>\alpha$.
Now consider $g_{t}:=\widetilde{F}_{t}-\widetilde{F}$. We have
$\sup_{\partial B(z_{1},r)}|g_{t}(z)|\leq\sup_{\partial B(z_{1},r)}\left(|f_{t}(z)-F(z)|+|f_{t}(z_{2})-F(z_{2})|\right).$
Hence, there exists $N_{0}\in\mathbb{N}$ such that $\sup_{\partial B(z_{1},r)}|g_{t}(z)|\leq\alpha$ for all $t>N_{0}$. Hence, we obtain that
$\sup_{\partial B(z_{1},r)}|g_{t}(z)|\leq\alpha<|\widetilde{F}(z)|\;\;\forall z\in\partial B(z_{1},r).$
Therefore, we infer from Result 2.5, that $\widetilde{F}$ and $\widetilde{F_{t}}$ has same number of zeros on $B(z_{1},r)$ for all $t>N_{0}$. This contradicts the fact that $f_{t}$ are injective. Therefore, $F:\Omega\to\mathbb{C}^{n}$ is an injective holomorphic map. ∎
We now present the proof of the main theorem of this paper Theorem 1.5
Proof of Theorem 1.5.
Here $V$ is a complete holomorphic vector field, $X\colon\mathbb{R}\times\mathbb{C}^{n}\to\mathbb{C}^{n}$ is the flow of the vector field $V$. Let $\Phi:\Omega\to\Phi(\Omega)$ be a biholomorphism such that $\Phi(\Omega)$ is Runge. Define a map $H\colon[0,1]\times\Omega\to\mathbb{C}^{n}$ by
$H(t,z):=\begin{cases}e^{c\ln{t}A}(\Phi(X(-c\ln{t},z))),~{}\text{if}~{}t\in(0,1]\\ F(z),~{}\text{if}~{}t=0,\\ \end{cases}$ (4.8)
where $F:\Omega\to\mathbb{C}^{n}$ is a biholomorphism onto its image is chosen in such way $e^{c\ln{t}A}\circ X_{-c\ln{t}}(z)$ converges to the map $F(z)$ as $t\to 0^{+}$ uniformly over every compact subset of $\Omega$. In view of Lemma 4.1, we get the existence of such a biholomorphism $F\colon\Omega\to\mathbb{C}^{n}$. Clearly, the biholomorphism $F$ is a uniform limit of a sequence of automorphisms of $\mathbb{C}^{n}$ on the compact subsets of $\Omega$. By assumption, there exists $\epsilon_{1}>0$ such that $02\alpha+k_{-}(A)-\epsilon_{1}>$ 0. We choose a real number $c>\frac{1}{2\alpha+k_{-}(A)-\epsilon_{1}}>0$ and fix it. We will first show that $H$ is a $\mathcal{C}^{1}$-map and for each fixed $t\in[0,1]$, $H_{t}(\Omega)$ is Runge. Then, we will use the Result 1.3. It is enough to show that $H$ is $\mathcal{C}^{1}$ at the points of the form $(0,z)$. Since any translation map and an invertible linear map is particularly an automorphism of $\mathbb{C}^{n}$, hence, without loss of generality, we assume that $\Phi(0)=0$ and $D\Phi(0)=I_{n}$. Let $K\subset\Omega$ be a compact set. We now divide the proof into four steps.
Step I: Showing that $H$ is continuous at $(0,z)$.
Expanding the power series expansion of the biholomorphism $\Phi:\Omega\rightarrow\mathbb{C}^{n}$ around the origin, we obtain that there exist constants $M_{1},\delta>0$ such that $\Phi$ can be expressed as:
$\Phi(z)$ $=z+R_{1}(z)\quad\text{for all}~{}z\in B(0,\delta),$ (4.9)
where
$\|R_{1}(z)\|\leq M_{1}\|z\|^{2}~{}\text{for all}~{}z\in B(0,\delta).$ (4.10)
Clearly, $-c\ln{t}\to\infty$ as $t\to 0^{+}$. Since, by the assumptions $(i)$ and $(ii)$, the flow of the vector field $V$ has uniform exponential decay on the domain $\Omega$, hence, there exists $\delta_{K}>0$ such that $X(-c\ln{t},z)\in B(0,\delta)$, for all $z\in K$ and for all $t\in(0,\delta_{K})$. Hence, by (4.10), we obtain
$\|R_{1}(X(-c\ln{t},z))\|$ $\leq M_{1}\|X(-c\ln{t},z)\|^{2}~{}\forall z\in K,~{}\forall t\in(0,\delta_{K}).$ (4.11)
Again, using the uniform exponential decay in assumption $(ii)$, we obtain that
$\|X(-c\ln{t},z)\|$ $\leq\gamma e^{\alpha c\ln{t}}\|z\|~{}\forall t>0.$ (4.12)
Hence, from (4.11) and (4.12), we get that
$\|R_{1}(X(-c\ln{t},z))\|$ $\leq M_{1}\gamma^{2}t^{2c\alpha}\|z\|^{2}~{}\forall z\in K,~{}\forall t\in(0,\delta_{K}).$ (4.13)
By using Result 2.6 and (4.13), we deduce that for all $z\in K$ and for all $t\in(0,\delta_{K})$ the following holds:
$\|e^{c\ln{t}A}(R_{1}(X(-c\ln{t},z)))\|$ $\leq\rho(\varepsilon_{1})e^{c\ln{t}(k_{-}(A)-\varepsilon_{1})}\|R_{1}(X(-c\ln{t},z))\|$
$\leq M_{1}\rho(\varepsilon_{1})\gamma^{2}e^{c\ln{t}(k_{-}(A)-\varepsilon_{1})}t^{2c\alpha}\|z\|^{2}$
$=M^{\prime}t^{c(k_{-}(A)+2\alpha-\varepsilon_{1})}.$ (4.14)
Here $M^{\prime}:=\sup_{z\in K}M_{1}\rho(\varepsilon_{1})\gamma^{2}\|z\|^{2}$. Since from (1.3) we have $k_{-}(A)+2\alpha-\varepsilon_{1}>0$, hence, from (4), we conclude that $e^{c\ln{t}A}(R_{1}(X(-c\ln{t},z)))\to 0$ as $t\to 0^{+}$, uniformly $K$. Now for any $(t,z)\in(0,\delta_{K})\times K$ we obtain
$H(t,z)$ $=e^{c\ln{t}A}(\Phi(X(-c\ln{t},z)))$
$=e^{c\ln{t}A}(X(-c\ln{t},z))+e^{c\ln{t}A}\big{(}R_{1}(X(-c\ln{t},z))\big{)}.$ (4.15)
Therefore,
$\lim_{t\to 0^{+}}H(t,z)=\lim_{t\to 0^{+}}e^{c\ln{t}A}(X(-c\ln{t},z)).$
Since, by Lemma 4.1, we obtain that $\lim_{t\to 0^{+}}e^{c\ln{t}A}\circ(X(-c\ln{t},z))=F(z)$ and the convergence is uniform on every compact subset of $\Omega$, therefore, $\lim_{t\to 0^{+}}H(t,z)=F(z)=H(0,z)$. Hence, $H$ is continuous at $(0,z)$.
Step II: Showing that $H$ is differentiable at $(0,z)$.
We want to show that $DH(0,z)=DF(z)$ for every $z\in\Omega$. For every $t\in(0,\delta_{K})$ , $z\in K$ , and for $h$ in a small enough neighborhood $U_{0}$ of the origin we get that
$\|H(t,z+h)-H(0,z)-DF(z)h\|$
$\qquad\qquad\qquad\qquad\qquad=\|e^{c\ln{t}A}(\Phi(X(-c\ln{t},z+h)))-F(z)-DF(z)h\|.$
Using the expression of the map $\Phi$ from (4.9), we get that
$\|H(t,z+h)-H(0,z)-DF(z)h\|$
$=\|e^{c\ln{t}A}(X(-c\ln{t},z)+R_{1}(X(-c\ln{t},z)))-F(z)-DF(z)h\|$
$\leq\|e^{c\ln{t}A}(X(-c\ln{t},z+h))-F(z+h)\|$
$+\|e^{c\ln{t}A}\big{(}R_{1}(X(-c\ln{t},z+h)\big{)}+\|F(z+h)-F(z)-DF(z)h\|.$ (4.16)
Hence, from (4), we obtain, for $(t,z)\in(0,\delta_{K})\times K$ and $h\in U_{0},\;h\neq 0$, that
$\frac{\|H(t,z+h)-H(0,z)-DF(z)h\|}{\|(t,h)\|}$
$\leq\frac{\|e^{c\ln{t}A}(X(-c\ln{t},z+h))-F(z+h)\|}{\|(t,h)\|}+\frac{\|e^{c\ln{t}A}\big{(}R_{1}(X(-c\ln{t},z+h)\big{)}\|}{\|(t,h)\|}$
$+\frac{\|F(z+h)-F(z)-DF(z)h\|}{\|(t,h)\|}$
$\leq\frac{\|e^{c\ln{t}A}(X(-c\ln{t},z+h))-F(z+h)\|}{t}+\frac{\|e^{c\ln{t}A}\big{(}R_{1}(X(-c\ln{t},z+h)\big{)}\|}{t}$
$+\frac{\|F(z+h)-F(z)-DF(z)h\|}{\|h\|}.$ (4.17)
Next, we show that the three summands that appeared on the right-hand side of the (4), tend to zero as $\|(t,h)\|\to(0,0)$ with $t>0$. Clearly,
$\frac{\|F(z+h)-F(z)-DF(z)h\|}{\|h\|}\to 0~{}~{}\text{as}~{}~{}\|h\|~{}~{}\to 0.$ (4.18)
By similar computation as (4), we conclude that there exists $M^{\prime\prime}>0$, such that for every $(t,h)\in(0,1)\times\mathbb{C}^{n}$ with $\|(t,h)\|$ is small enough, the following holds:
$\frac{\|e^{c\ln{t}A}\big{(}R_{1}(X(-c\ln{t},z+h)\big{)}\|}{t}$ $\leq M^{\prime\prime}t^{c(2\alpha+k_{-}(A)-\varepsilon_{1})-1},~{}~{}\forall z\in K.$ (4.19)
From the choice of $c>0$, it follows that the right-hand side of the (4.19) tends to 00 as $\|(t,h)\|\to 0$ with $t>0$. We now deal with the first summand of (4). We now view the vector field $V$ as an entire mapping such that $V(0)=0$, $DV(0)=A$. Hence, we can write
$V(z)$ $=Az+R(z).$ (4.20)
Define a map $S\colon[0,1]\times\Omega\to\mathbb{C}^{n}$ by
$S(t,z)=\begin{cases}e^{c\ln{t}A}\circ X_{-c\ln{t}}(z)~{}\quad\text{if}~{}t\in(0,1]\\ F(z)~{}\quad\text{if}~{}t=0\\ \end{cases}.$
The first summand in (4) is
$\frac{S(t,z+h)-S(0,z+h)}{t}.$ (4.21)
It follows, from Lemma 4.1, that the map $t\to S(t,z+h)$ is continuous on $[0,1]$. It is also differentiable on $(0,1)$. Hence, by using the mean value theorem in (4.21), we obtain that
$\frac{\|S(t,z+h)-S(0,z+h)\|}{t}$ $\leq\bigg{\|}\frac{dS(t,z)}{dt}\bigg{|}_{t=\tau}\bigg{\|}.$ (4.22)
Here we do some computation. For $t\in(0,1)$, we have
$\frac{dS(t,z)}{dt}$ $=\frac{d}{dt}(e^{c\ln{t}A}\circ X_{-c\ln{t}}(z))$
$=\frac{c}{t}Ae^{c\ln{t}A}\big{(}X(-c\ln{t},z)\big{)}+e^{c\ln{t}A}\bigg{(}\frac{dX(-c\ln{t},z)}{dt}\bigg{)}$
$=\frac{c}{t}Ae^{c\ln{t}A}\big{(}X(-c\ln{t},z)\big{)}-\frac{c}{t}e^{c\ln{t}A}(V(X(-c\ln{t},z))).$
Since we have $V(z)=A(z)+R(z)$, hence, from (4) we get
$\frac{dS(t,z)}{dt}$ $=-\frac{c}{t}Ae^{c\ln{t}A}\big{(}R(X(-c\ln{t},z))\big{)}.$ (4.24)
From (4.21) and (4.24) we deduce that
$\frac{\|S(t,z+h)-S(0,z+h)\|}{t}$ $\leq\bigg{\|}\frac{dS(t,z+h)}{dt}\bigg{|}_{t=\tau}\bigg{\|},$
where $0<\tau<t$. Therefore, we obtain
$\frac{\|S(t,z+h)-S(0,z+h)\|}{t}$ $\leq\frac{c}{\tau}\bigg{\|}Ae^{c\ln{\tau}A}\big{(}R(X(-c\ln{\tau},z+h))\big{)}\bigg{\|}.$ (4.25)
There is a neighborhood $U$ of the origin such that $\|R(z)\|\leq M\|z\|^{2}$ for all $z\in U$. Since $X(-c\ln{t},z)\to 0$ as $t\to 0^{+}$ uniformly on every compact subset of $\Omega$. Hence, from Result 2.6 and the equation (4.25), we obtain that
$\frac{\|S(t,z+h)-S(0,z+h)\|}{t}$ $\leq M\gamma^{2}\rho(\varepsilon_{1})t^{c(2\alpha+k_{-}(A)-\varepsilon_{1})-1}\|z+h\|^{2}.$ (4.26)
By the choice of $c>0$, it follows from (4.26) that
$\frac{\|e^{c\ln{t}A}(X(-c\ln{t},z+h))-F(z+h)\|}{t}\to 0~{}\text{as}~{}t\to 0^{+}~{}\text{uniformly on $K$.}$
Therefore, we infer from (4.26), (4.19) and (4.18) that left-hand side of (4) tends to 00 as $\|(t,h)\|\to 0$ uniformly on $K$. Consequently, it proves that the map $H$ is differentiable at $(0,z)$.
Step III: Showing that $H$ is continuously differentiable at $(0,z)$.
We show that $\frac{dH(t,z)}{dt}\to 0$ uniformly over $K$ as $t\to 0^{+}.$ We now compute the derivative.
$\frac{dH(t,z)}{dt}$ $=\frac{d}{dt}(e^{c\ln{t}A}(\Phi(X(-c\ln{t},z))))$
$=\frac{c}{t}e^{c\ln{t}A}\big{(}A\Phi(X(-c\ln{t},z))-D\Phi(X(-c\ln{t},z))V(X(-c\ln{t},z))\big{)}$
$=\frac{c}{t}e^{c\ln{t}A}\big{(}\Psi(X(-c\ln{t},z))\big{)}.$ (4.27)
Here $\Psi(z)=A(\Phi(z))-D\Phi(z)V(z)$. Clearly, $\Psi(0)=0$. We prove that $D\Psi(0)=0$. We deduce the following for all non-zero $h\in\mathbb{C}^{n}$
$\|\Psi(h)\|$ $=\|A(\Phi(h))-D\Phi(h)V(h)\|$
$\leq\|A(\Phi(h))-Ah\|+\|A(h)-D\Phi(h)Ah\|+\|D\Phi(h)R(h))\|$
$\leq\|Ah+A(R_{1}(h))-Ah\|+\|A(h)-D\Phi(h)Ah\|+\|D\Phi(h)R(h))\|$
$\leq\|A\|\|(R_{1}(h))\|+\|I-D\Phi(h)\|\|Ah\|+\|D\Phi(h)R(h))\|.$ (4.28)
We have $\|R_{1}(h)\|\leq M_{1}\|h\|^{2}$ and $\|R(h)\|\leq M\|h\|^{2}$ for all $h\in U\cap B(0,\delta^{{}^{\prime}})$. Also, $D\Phi(h)\to I$ as $h\to 0$. Hence, from (4), we get that $\frac{\|\Psi(h)\|}{\|h\|}\to 0$ as $\|h\|\to 0.$ This proves that $D\Psi(0)=0$. Therefore, we conclude that there exist $M_{2},\delta^{\prime}>0$ such that
$\|\Psi(z)\|\leq M_{2}\|z\|^{2},~{}~{}\forall z\in B(0,\delta^{\prime}).$
We note that $X(-c\ln{t},z)\to 0$ as $t\to 0^{+}$ uniformly on $K$. Hence, by invoking Result 2.6, we conclude that there exists $\delta^{\prime}_{K}>0$ such that
$\bigg{\|}\frac{dH(t,z)}{dt}\bigg{\|}$ $\leq c\rho(\varepsilon)M_{2}\gamma^{2}t^{c(2\alpha+k_{-}(A)-\varepsilon_{1})-1}\|z\|^{2}~{}~{}\forall t\in(0,\delta_{K}^{\prime}),~{}\forall z\in K$ (4.29)
It follows from the choice of $\varepsilon_{1},c>0$ that $\bigg{\|}\frac{dH(t,z)}{dt}\bigg{\|}\to 0$ as $t\to 0^{+}$ uniformly on $K$. Since $H_{t}$ is holomorphic on $\Omega$, hence $DH_{t}(z)\to DH_{0}(z)=DF(z)$ uniformly on $K$ as $t\to 0^{+}$. This proves that $H$ is $\mathcal{C}^{1}$-smooth map at $(0,z)$.
Step IV: Showing that $H_{t}(\Omega)$ is Runge for $t\in[0,1]$.
Since both $X_{-c\ln t}$ and $X_{c\ln t}$ are automorphisms of $\mathbb{C}^{n}$, it follows from Lemma 2.7, that $H_{t}(\Omega)$ is Runge. Hence, we are done. ∎
Proof of Corollary 1.8.
Define a vector field $V$ on $\mathbb{C}^{m+n}$ by
$V(z,w):=(V_{1}(z),V_{2}(w))\quad(z,w)\in\mathbb{C}^{n}\times\mathbb{C}^{m}.$
Let $X_{1}(t,z)$ and $X_{2}(t,w)$ be the flows of the vector field $V_{1}$ and $V_{2}$ respectively. Then $\theta(t,(z,w))=(X_{1}(t,z),X_{2}(t,w))$ is the flow of the vector field $V\in\mathfrak{X}_{\mathcal{O}}(\mathbb{C}^{n+m})$. Clearly, if $(z,w)\in\Omega_{1}\times\Omega_{2}$, then $\theta_{t}(z,w)\in\Omega_{1}\times\Omega_{2}$ for every $t\geq 0$. Hence, $\Omega_{1}\times\Omega_{2}$ is invariant under the positive time flow of the vector field $V$.
From the assumption we have that
$2\alpha_{1}+k_{-}(A)$ $>0$
$2\alpha_{2}+k_{-}(B)$ $>0.$
Here $A=DV_{1}(0)$ and $B=DV_{2}(0)$.
Therefore, we obtain
$\|X_{1}(t,z)\|^{2}+\|X_{2}(t,w)\|^{2}$ $\leq\gamma_{1}^{2}e^{-2\alpha_{1}t}\|z\|^{2}+\gamma_{2}^{2}e^{-2\alpha_{2}t}\|w\|^{2}$
$\|\theta(t,(z,w))\|^{2}$ $\leq\tilde{\gamma}^{2}e^{-2\alpha t}\|(z,w)\|^{2}$
$\|\theta(t,(z,w))\|$ $\leq\tilde{\gamma}e^{-\alpha t}\|(z,w)\|,$ (4.30)
where $\tilde{\gamma}:=\max\{\gamma_{1},\gamma_{2}\}$ and $\alpha:=\min\{\alpha_{1},\alpha_{2}\}$. From the assumption we have $\alpha+\min\{\frac{k_{-}(A)}{2},\frac{k_{-}(B)}{2}\}>0$.
Therefore, the rest of the proof follows from Theorem 1.5. ∎
Proof of Theorem 1.10.
(i) Suppose that $\Phi:\Omega\to\Phi(\Omega)$ is a biholomorphism with $\text{det}D\Phi(z)\equiv 1$ and $\Phi(\Omega)$ is Runge domain. According to our assumption, $div_{\omega}(V)$ is a non-zero constant. Let $div_{\omega}(V)=a\neq 0$. By Lemma 2.8, we get that
$\frac{d}{dt}\det D_{z}(X(t,z))$ $=a\det~{}D_{z}(X(t,z)),~{}\text{det}D_{z}(X(0,z))=1,$
where $D_{z}(X(t,z))$ denotes the derivative matrix of the automorphism $X_{t}\in\text{Aut}(\mathbb{C}^{n})$ with respect to $z$. Solving the above differential equation, we obtain that
$\text{det}D_{z}(X(t,z))$ $=e^{at},\;\;\forall z\in\mathbb{C}^{n},\;\;\forall t\geq 0.$ (4.31)
Since $\det D_{z}X_{t}(0)=e^{tr(A)t}$, where $A=DV(0)$, hence, from (4.31), we get that $a=tr(A)$. We also get from (4.8)
$D_{z}H_{t}(z)=e^{c\ln{t}A}\circ D_{z}(\Phi(X(-\ln{t},z)))\circ D_{z}X_{-\ln{t}}(z).$
Since $\text{det}(D\Phi(X(-c\ln{t},z)))\equiv 1$, hence, from (4.31) we get that
$\det D_{z}H_{t}(z)=e^{tr(A)\ln{t}-tr(A)\ln{t}}=1,~{}~{}\forall t\in(0,1).$
Since $H_{t}(z)\to F(z)$ as $t\to 0^{+}$, hence, $\text{det}H_{t}(z)\to\text{det}F(z)$ as $t\to 0^{+}.$ Therefore, $\text{det}F(z)\equiv 1$ on $\Omega$.
We now show that identity map $\iota_{\Omega}:\Omega\to\Omega$ is homotopic equivalent to a constant map. Define $\widetilde{H}:[0,1]\times\Omega\to\mathbb{C}^{n}$ by
$\widetilde{H}(t,z)=\begin{cases}X(-\ln{t},z),~{}\text{if}~{}t\in(0,1]\\ 0,~{}\text{if}~{}t=0.\\ \end{cases}.$
The global exponential growth of the flow map provides the continuity of the map $\widetilde{H}(t,z)$ at the point $(0,z)$. This proves that the domain $\Omega$ is contractible to 00. Hence, $H^{(n-1)}(\Omega,\mathbb{C})$ vanishes identically. Since $\Omega$ is also pseudoconvex, therefore, by Result 1.3, we conclude that $H_{1}(z)=\Phi(z)$ can be approximated by $\text{Aut}_{1}(\mathbb{C}^{n})$ uniformly on every compact subset of $\Omega$.
(ii) The flow of the symplectic vector field preserves the symplectic form. From our assumption $\Phi:\Omega\to\Phi(\Omega)$ is a symplectic holomorphic map. Hence, $H_{t}^{*}(\omega)=\omega$ for all $t\in(0,1]$. We have seen in the proof of (i) that $\Omega$ is contractible. Hence, $H^{(n-1)}(\Omega,\mathbb{C})=0$. We get that $H_{t}(\Omega)$ is Runge for every $t\in[0,1]$, from Step IV of the proof of Theorem 1.5. Now using [symfor1995, Proposition 2.3] we obtain our theorem.
5. Application in Loewner PDE
In this section first we state a generalization of [ABFW13, Theorem 3.4] and [Hamada15, Theorem 4.5].
Theorem 5.1.
Let $\Omega$ be a complex manifold of dimension $N\geq 2$. Let $\{\Omega_{j}\}_{j\in\mathbb{N}\cup\{0\}}$ be a sequence of open connected subset of $\Omega$ such that $\Omega_{j}\subset\Omega_{j+1}$ for every $j\in\mathbb{N}$ and $\Omega=\cup_{j\in\mathbb{N}\cup\{0\}}\Omega_{j}$. Assume that
  1. (i)
    each pair $(\Omega_{j},\Omega_{j+1})$ is a Runge pair.
  2. (ii)
    each $\Omega_{j}$ is biholomorphic to a Stein domain that is invariant under positive time flows of a vector field $V_{j}\in\mathfrak{X}_{\mathcal{O}}(\mathbb{C}^{n})$ satisfying the conditions of Theorem 1.5.
Then $\Omega$ is biholomorphic to a Runge and Stein domain in $\mathbb{C}^{n}$.
Since we have the approximation on certain domains satisfying conditions Theorem 1.5, Corollary 1.7, hence, the proof of the above theorem follows exactly the same way as [ABFW13, Theorem 3.4]. Therefore, we omit the proof here.
We now give proof of Theorem 1.11.
Proof of Theorem 1.11.
Using Result 2.2 we get that there exists a Loewner chain $g_{t}:\Omega\to N$ of order $d$ that solves the Loewner PDE
$\frac{\partial f_{t}}{\partial t}(z)$ $=-df_{t}(z)(G(z,t)),~{}~{}a.e.\quad t\geq 0,~{}~{}\forall z\in\Omega,$ (5.1)
where $N=\bigcup_{t\geq 0}g_{t}(\Omega)$ is a complex manifold of dimension $n$. We also obtain that any other solution to equation (5.1) with values in another $n$ dimensional complex manifold $Q$ is of the form $\Lambda\circ g_{t}$, where $\Lambda:N\to Q$ is holomorphic. By using Result 2.4, $\left(g_{j}(\Omega),g_{j+1}(\Omega)\right)$ is a Runge pair for every $j\in\mathbb{N}$. Since $g_{t}$ is a univalent holomorphic map, $g_{t}(\Omega)$ is an open connected subset of $N$. Clearly, $g_{t}(\Omega)$ is biholomorphic to $\Omega$. By assumption, $\Omega$ is invariant under positive time flows of $V$ such that $V$ satisfies the (1.3). Therefore, $g_{j}(\Omega)$ satisfies the Condition $(ii)$ of Theorem 5.1 for all $j\in\mathbb{N}$. Hence, thanks to Theorem 5.1, we get that $N$ is biholomorphic to some Runge Stein domain $G\subset\mathbb{C}^{n}$. Hence, it follows that there exists a biholomorphic map $F:N\to G$. Let $f_{t}=F\circ g_{t}$. We have the following
$\frac{\partial f_{t}}{\partial t}(z)$ $=DF(g_{t}(z))\frac{\partial g_{t}}{\partial t}(z)$
$=DF(g_{t}(z))(-dg_{t}(z)(G(z,t)))$
$=-D(F(g_{t}(z)))G((z,t))$
$=-D(f_{t}(z))(G(z,t)).$
This implies that $f_{t}$ satisfies (5.1). If $\tilde{g_{t}}(z)$ is another solution of the (5.1) with values in $\mathbb{C}^{n}$, then we deduce, from Theorem 2.2, that there is a biholomorphism $\Lambda:N\to\cup_{t\geq 0}\tilde{g_{t}}(\Omega)$ such that $\tilde{g_{t}}=\Lambda\circ g_{t}=\Lambda\circ F^{-1}(f_{t})$. Consider $\varphi=\Lambda\circ F^{-1}$. Therefore, any solution of (5.1) with values in $\mathbb{C}^{n}$ is of the form $\varphi(f_{t})$ for some suitable biholomorphism $\varphi\colon G\to\mathbb{C}^{n}$. ∎
6. Examples
In this section, we provide a some examples of domains that satisfy the conditions of our theorems.
Example 6.1.
Let $A=\text{Diag}(-\lambda_{1},-\lambda_{2},\cdots,-\lambda_{n})$, such that $\lambda_{i}\in\mathbb{N}$, with, $\lambda_{1}\leq\lambda_{2}\cdots\leq\lambda_{n}$ and $2\lambda_{1}>\lambda_{n}$. Let $i,j\in\{1,2,\cdots,n\}$ and $m_{i},m_{j}\in\mathbb{N}$, such that $m_{i}\lambda_{i}=m_{j}\lambda_{j}$. Let $\Omega:=\{z\in\mathbb{C}^{n}:\big{|}z_{i}^{m_{i}}-z_{j}^{m_{j}}\big{|}<1\}$. Then every biholomorphism from $\Omega$ onto a Runge domain can be approximated by elements of $\text{Aut}(\mathbb{C}^{n})$.
Explanation. Let $z\in\Omega$ and $w=e^{tA}z$. Then we obtain that
$|w_{i}^{m_{i}}-w_{j}^{m_{j}}|=|e^{-m_{i}\lambda_{i}t}z_{i}-e^{-m_{j}\lambda_{j}t}z_{j}|=e^{-m_{i}\lambda_{i}}|z_{i}-z_{j}|<1.$
Since $2\lambda_{1}>\lambda_{n}$, hence, it satisfies the condition of the Corollary 1.7. Therefore, the conclusion is true.
The next example demonstrates a class of domains that satisfy the conditions of Corollary 1.7 but do not satisy the conditions of Result 1.4.
Example 6.2.
Let $A=\lambda I_{n}+\mathcal{N}$, where $\mathcal{N}$ is a $n\times n$ matrix with ones in the first diagonal above the main diagonal and zeros elsewhere and $\lambda=\lambda_{1}+i\mu_{1}$, with $\lambda_{1}<0$. Let $\Omega:=\{(z_{1},z_{2},,\ldots,z_{n}):\bigg{|}z_{n-1}-\frac{z_{n}}{\lambda_{1}}\ln{|z_{n}|}\bigg{|}<1\}$. Then every biholomorphism from $\Omega$ onto a Runge domain can be approximated by elements of $\text{Aut}(\mathbb{C}^{n})$.
Explanation. Suppose that $A=-\lambda I_{n}+\mathcal{N}$. Let $w=e^{At}z$. We get that
$w=e^{At}=e^{\lambda_{1}t}\begin{bmatrix}1&t&\frac{t^{2}}{2}&\cdots&&\frac{t^{n}}{n!}\\ 0&1&t&\cdots&&\frac{t^{n-1}}{(n-1)!}\\ \vdots&\vdots&\vdots&\ddots&&\vdots\\ 0&0&0&\cdots&&t\\ 0&0&0&\cdots&&1\\ \end{bmatrix}.$
For all $z\in\Omega$ and for all $t\geq 0$, we obtain that
$\bigg{|}w_{n-1}-\frac{w_{n}}{\lambda_{1}}\ln{|w_{n}|}\bigg{|}$ $=\bigg{|}z_{n-1}e^{\lambda t}+tz_{n}e^{\lambda t}-\frac{z_{n}e^{\lambda t}}{\lambda_{1}}(\ln{|z_{n}|}+t\lambda_{1})\bigg{|}$
$=\bigg{|}z_{n-1}e^{\lambda t}-e^{\lambda t}\frac{z_{n}}{\lambda_{1}}\ln{|z_{n}|}\bigg{|}$
$=e^{\lambda_{1}t}\bigg{|}z_{n-1}-\frac{z_{n}}{\lambda_{1}}\ln{|z_{n}|}\bigg{|}$
$<1.$
Clearly, the matrix $A$ satisfies $2k_{+}(A)=2\lambda_{1}<\lambda_{1}$. Therefore, from Corollary 1.7 it follows that any biholomorphism of $\Omega$, whose image is a Runge domain in $\mathbb{C}^{n}$, can be approximated by elements of $\text{Aut}(\mathbb{C}^{n})$. But the matrix $-A=-\lambda I_{n}-\mathcal{N}$ may not satisfy the condition of Result 1.4. Since for any $B\in M_{n}(\mathbb{C})$ we have that $m(B):=\inf\{{\sf Re}\langle Bz,z\rangle:\|z\|=1\}=\min\{{\sf Re}(\lambda):\lambda\in\sigma(\frac{1}{2}(B+B^{*}))\}$, hence for $n=3$ $m(A)=-\lambda_{1}-\sqrt{2}$ and $k_{+}(A)=-\lambda_{1}$. Therefore, for $n=3$ the matrix satisfies the condition given in Result 1.4 if and only if $\lambda_{1}<-2\sqrt{2}$.
Example 6.3.
Let $D\subset\mathbb{C}^{n}$ be a domain such that $D$ is invariant under the positive time flow of the vector field $V\in\mathfrak{X}_{\mathcal{O}}(\mathbb{C}^{n})$. Assume that $V$ satisfies the condition of Theorem 1.5. Let $\Omega=\{(z,w)\in\mathbb{C}^{n+1}:z\in D,~{}\|w\|<1\}\subset\mathbb{C}^{n+1}$. Then $\Omega$ is invariant under positive time flows of the vector field $\widetilde{V}=(V(z),-\beta w)$, where $\beta>0$ is chosen suitably so that the vector field $\widetilde{V}$ satisfies the condition of Theorem 1.5.
Explanation. The flow of the vector field $\tilde{V}$ is $\tilde{X}(t,(z,w))=(X(t,z),e^{-\beta t}w)$, where $X(t,z)$ denotes the flow of the vector field $V$. Clearly, $\tilde{X}_{t}(\Omega)\subset\Omega$ for all $t>0$. Here $\tilde{X}_{t}$ is the flow of the vector field $\widetilde{V}=(V(z),-\beta w)$ and $\beta>0$ will be chosen later. Let $DV(0)=A$. Since the flow of the vector field $V$ satisfies the condition of the Theorem 1.5, hence, for all $z=(z_{1},z_{2})\in D$, there exist $\gamma,\alpha>0$, such that
$\|X(t,z)\|$ $\leq\gamma e^{-\alpha t}\|z\|,\;\;\forall z\in D\;\;\forall t>0,$ (6.1)
and $\alpha>\frac{-k_{-}(A)}{2}$. Choose $\beta>0$ such that
$\min\{\alpha,-k_{-}(A)\}>\beta>\frac{-k_{-}(A)}{2}.$
For $(z,w)\in\Omega$ and $t>0$, we have the following estimate:
$\|\tilde{X}(t,(z,w))\|^{2}$ $=\|X(t,z)\|^{2}+e^{-2\beta t}\|w\|^{2}.$
Hence, from (6.1) we get that
$\|\tilde{X}(t,(z,w))\|^{2}$ $\leq\tilde{\gamma}^{2}e^{-2\beta t}(\|z\|^{2}+\|w\|^{2}),$
where $\tilde{\gamma}^{2}:=\max\{1,\gamma^{2}\}.$ Therefore, for every $(z,w)\neq(0,0)$, we obtain that
$\|\tilde{X}(t,(z,w))\|$ $\leq\tilde{\gamma}e^{-\beta t}\|(z,w)\|.$
Clearly, from the choice of $\beta>0$ we have $k_{-}(D\tilde{V}(0))=\min\{k_{-}(A),-\beta\}=k_{-}(A)$. Again from the choice of $\beta>0$, we have $02\beta+k_{-}(D\widetilde{V}(0))>$ 0. Therefore, it satisfies Condition (1.3) of Theorem 1.5.
Next, we provide an example of a complete hyperbolic Hartogs domain which is invariant in the positive time flow of a nonlinear vector field but not invariant under the flow of any matrix. More precisely:
Example 6.4.
Consider the domain $\Omega=\{(z_{1},z_{2})\in\mathbb{C}^{2}:~{}{\sf Re}(z_{1})<3,~{}|z_{2}|<e^{-\frac{1}{2}{\sf Re}(z_{1})})\}$. $\Omega$ is a complete hyperbolic Hartogs domain which is invariant under positive time flows of a non-linear holomorphic vector field but not invariant under positive time flows of any linear holomorphic vector field.
Explanation: We first show that $\Omega$ is invariant under positive time flows of the vector field $V(z_{1},z_{2})=(-2z_{1},-3z_{2}+z_{1}z_{2})$. The flow of the vector field $V$ is
$X(t,z)=\left(z_{1}e^{-2t},z_{2}e^{-3t}e^{\frac{z_{1}}{2}(1-e^{-2t})}\right),\;\;\forall t\in\mathbb{R}.$
Let $(z_{1},z_{2})\in\Omega$ and $t\geq 0$. Let $w_{1}=z_{1}e^{-2t}$ and $w_{2}=z_{2}e^{-3t}e^{\frac{z_{1}}{2}(1-e^{-2t})}$. Then we have:
$|w_{2}|$ $=|z_{2}|e^{-3t}||e^{\frac{z_{1}}{2}(1-e^{-2t})}|$
$\leq|z_{2}||e^{\frac{z_{1}}{2}(1-e^{-2t})}|.$
Since $(z_{1},z_{2})\in\Omega$, we get that
$|w_{2}|$ $\leq e^{-\frac{1}{2}{\sf Re}(z_{1})}e^{\frac{1}{2}{\sf Re}(z_{1})(1-e^{-2t})}$
$\leq e^{-\frac{1}{2}{\sf Re}(z_{1}e^{-2t})}$
$<e^{-\frac{1}{2}{\sf Re}(w_{1})}.$
It is clear that ${\sf Re}(w_{1})<3$. Therefore, we get $(w_{1},w_{2})\in\Omega$ for all $t\geq 0$.
Next, we claim that $\Omega$ is complete hyperbolic. We use Result 2.3 to prove this lemma. Consider the sequence $c_{j}=\frac{1}{2}$ and $h_{j}(z)=e^{z}$ for $j\in\mathbb{N}$. Then $\{c_{j}\log|{h_{j}}(z)|\}$ converges to $\frac{1}{2}{\sf Re}(z)$. Since upper half space $H:=\{z\in\mathbb{C}:\Im(z)>0\}$ is biholomorphic to a domain $G:=\{z\in\mathbb{C}:{\sf Re}(z)<3\}\subset\mathbb{C}$. Hence, $G$ is a complete hyperbolic domain. We now use Result 2.3 to conclude that $\Omega$ is a complete hyperbolic domain. Next, we show that $\Omega$ satisfies the conditions of Theorem 1.5. We show that the flow of the vector field $V$ satisfies Condition (1.3). Let $\zeta=(\zeta_{1},\zeta_{2})\in\Omega$. We have the estimate of the flow on the domain $\Omega$
$\|X(t,\zeta)\|^{2}$ $=e^{-4t}\|\zeta_{1}\|^{2}+\|\zeta_{2}\|^{2}e^{-6t}e^{{\sf Re}(\zeta_{1})(1-e^{-2t})}.$
Here we have ${\sf Re}(\zeta_{1})<3$. We also get that there exists $M>0$ such that $-2t+(1-e^{-2t}){\sf Re}{\zeta_{1}}<1$ for all $t>M$ Let $s=\sup_{t\in[0,M]}-2t+3(1-e^{-2t})$ Therefore, we have
$\|X(t,\zeta)\|^{2}$ $\leq\gamma^{2}e^{-4t}\big{(}\|\zeta_{1}\|^{2}+\|\zeta_{2}\|^{2}\big{)}$
$\|X(t,\zeta)\|$ $\leq\gamma e^{-2t}\|\zeta\|,$
where $\gamma^{2}:=\max\{1,s\}$.
Here $\alpha=2$ and $k_{-}(DF(0))=-3$. Hence, $02\alpha+k_{-}(DF(0))>$ 0. Hence, $\Omega$ satisfies (1.3). Therefore. in view of Theorem 1.1, any biholomorphism of $\Omega$ can be approximated uniformly on compacts by Aut$(\mathbb{C}^{2})$.
It can be shown that $\Omega$ is not invariant under positive time flows of any linear holomorphic vector field. It is clear that if all the eigenvalues of $A$ have positive real parts then a domain $\Omega$ containing the origin satisfies $e^{tA}\Omega\subset\Omega$ for all $t>0$ if and only if $\Omega=\mathbb{C}^{n}$. For given a matrix $A\in GL(n,\mathbb{C})$, all of whose eigenvalues have negative real part, we construct an initial point $z=(z_{1},z_{2})\in\Omega$ and $t>0$ (depending on $A$) such that $e^{tA}z\notin\Omega$. We provide the main steps here.
Step I: Showing that $\Omega$ is not invariant under positive time flows of the holomorphic linear vector fields corresponding to any diagonal matrix:
In this case, we may assume that $A=\begin{bmatrix}d_{1}&0\\ 0&d_{2}\\ \end{bmatrix}$ with $d_{1}=a_{1}+ib_{1}$ and $d_{2}=a_{2}+ib_{2}$ and $a_{j},b_{j}\in\mathbb{R}$ for $j=\{1,2\}$. If $b_{1}\neq 0$, then we choose $(-2p,1)\in\Omega$ with $\frac{a_{2}\pi}{|b_{1}|}+pe^{\frac{a_{1}\pi}{|b_{1}|}}>0$ and $t=\frac{\pi}{|b_{1}|}$. If $b_{1}=0$ then the choice of the initial point and the time depends on the sign of $a_{1}$ as follows: If $a_{1}<0$, then we choose the initial point $z=(-2p,z_{2})$ with $e^{p}>z_{2}>e^{pe^{a_{1}}+|a_{2}|}$ and $p=\frac{2|d_{2}|}{1-e^{a_{1}}}$ and $t_{1}=1$. If $a_{1}>0$ then we take $z=(2,\frac{1}{e^{2}})\in\Omega$ and $t>0$ with $e^{a_{1}t}+a_{2}t-2>0$.
Step II: Showing that $\Omega$ is not invariant under positive time flows of the holomorphic linear vector fields corresponding to any matrix of the form $\begin{bmatrix}\lambda&1\\ 0&\lambda\end{bmatrix}$: Here we consider the matrix of the form $\begin{bmatrix}\lambda&1\\ 0&\lambda\end{bmatrix}$, $\lambda=\lambda_{1}+i\lambda_{2}$. In case, $\lambda_{2}\neq 0$, we choose $z=(2p,-1)$ with $\frac{\lambda_{1}\pi}{|\lambda_{2}|}-\frac{\pi e^{\frac{\pi\lambda_{1}}{|\lambda_{2}|}}}{2|\lambda_{2}|}+pe^{\frac{\lambda_{1}\pi}{|\lambda_{2}|}}>0$ and $t=\frac{\pi}{|\lambda_{2}|}$. If $\lambda_{2}=0$ then the choice of the initial point and the time depends on the sign of $\lambda_{1}$. If $\lambda_{1}<0$ we take $(-2te^{-\lambda_{1}t},2e^{-\lambda_{1}t})\in\Omega$ such that $t>0$ and $12e^{t(k-e^{kt})}<$ 1.
If $A$ is similar to a diagonal matrix or a Jordan block then also we can compute the initial point inside the domain and the positive exit time of the integral curve. Since the computations are lengthy and of similar nature as Step I and II above, we omit the computation here.
Next, we provide an example of a domain in $\mathbb{C}^{n}$ similar to previous example.
Example 6.5.
Let $\Omega=\{(z_{1},\ldots,z_{n})\in\mathbb{C}^{n}:{\sf Re}(z_{1})<3,~{}|z_{j}|<e^{-\frac{1}{2}{\sf Re}(z_{1})},2\leq j\leq n\}$. Then $\Omega$ satisfies the condition of Theorem 1.5 with respect to the vector field $V(z_{1},z_{2},\ldots,z_{n})=(-2z_{1},-3z_{2}+z_{1}z_{2},-3z_{3}+z_{3}z_{1},\ldots,-3z_{n}+z_{n}z_{1})$. It can be proved, in similar lines to the proof of Example 6.4.
Acknowledgements. Sanjoy Chatterjee is supported by CSIR fellowship (File No-09/921(0283)/2019-EMR-I) and also would like to thank Golam Mostafa Mondal and Amar Deep Sarkar for several discussions. Sushil Gorai is partially supported by a Core Research Grant (CRG/2022/003560) of SERB, Dept. of Science and Technology, Govt. of India.